### v1

# Grassroots Consensus

Abstract
Grassroots platforms aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike. Within the grassroots architecture, consensus is needed to realize platforms that employ digital social contracts, which are like smart contracts except that they are among people not accounts and are executed by these people’s smartphones not by high-performance servers controlled by parties outside to the contract. Key envisioned grassroots platforms include sovereign democratic digital communities and federations, community banks and their grassroots cryptocurrencies, and digital cooperatives.
The grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during both low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots. The Grassroots Consensus protocol addresses all these requirements while having competitive performance in both low- and high-throughput scenarios and being one of the most concise and elegant consensus protocols for partial synchrony. It achieves that by building on two cutting-edge consensus protocols—the quiescent high-performance Morpheus and the blocklace-based Cordial Miners, improving the latter’s dissemination protocol and making it UDP-ready, and extending the protocol with a constitution and a constitutional amendment component, making it grassroots.
keywords:
Consensus Protocols, Grassroots Platforms, Blocklace, Digital Social Contracts, Democratic Digital Communities
1 Introduction
Grassroots. Grassroots platforms [31, 29, 27, 32, 33] aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike. A grassroots platform is different from a global platform in that it may have multiple instances that can operate independently of each other, without coordination or reliance on any third-party or global resource other than the network itself; furthermore, different instances of the same grassroots platform may interoperate and coalesce into ever-larger instances, possibly, but not necessarily, resulting in a single global instance.
Decentralized platforms for global cryptocurrencies [22, 37] and smart contracts [8, 5, 21] supporting Decentralized Finance (DeFi) [24, 13] and Decentralized Autonomous Organizations (DAOs) [11, 9, 35] often employ a consensus protocol as their foundation. Hence the surge in interest in high-performance and high-throughput consensus protocols [38, 15, 16, 19, 4, 14].
Grassroots platforms, on the other hand, are not consensus-based. Their foundational platform consists of a grassroots social graph implemented via a blocklace—a DAG-like generalization of the linear blockchain that is a universal Conflict-free Replicated Data Type (CRDT) [3]. The grassroots architecture [29, 31] facilitates the implementation of a grassroots social network [28] and grassroots cryptocurrencies [30, 19], which are also consensus-free, on top of the grassroots social graph. So where does consensus come into play?
Consensus. Within the grassroots architecture, consensus is needed to realize higher-level platforms that employ digital social contracts—the grassroots counterpart of smart contracts. Briefly, while a smart contract [8] is among anonymous accounts and runs on a consensus protocol executed by third-party servers, a social contract is among people known to each other and runs on a consensus protocol executed by the smartphones of these very same people [31, 33]. The two types of consensus-based contracts are compared in Table 1.
Table 1: Comparing Consensus-Based Smart Contracts and Social Contracts
Smart Contract [8] Digital Social Contract [7]
Among: Accounts People
Executed by: Third-party servers The people’s smartphones
In return for: Gas Love
Applications: Decentralized Finance (DeFi) [24, 13],
Decentralized Autonomous Organizations (DAOs) [11, 9],
Non-Fungible Tokens (NFTs) [36]
Grassroots democratic digital communities [23, 26] and their federation [31, 33],
community banks and digital cooperatives [30]
Amended via: Self-modifying code Constitutional amendments
Example: Grassroots Federation. A key application of digital social contracts mentioned in Table 1 is the grassroots federation of digital communities. Grassroots Federation [33] aims to address the egalitarian formation and the fair democratic governance of large-scale, decentralized, sovereign digital communities, the size of the EU, the US, existing social networks, and even humanity at large. A grassroots federation evolves via the grassroots formation of digital communities and their consensual federation. Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more. Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a small assembly elected by sortition [12, 33] among its members. It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community. Such a consensus protocol should satisfy all the requirements listed below.
Requirements of Grassroots Consensus. Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency. Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued one at a time—the low-throughput scenario—finalize each quickly and efficiently. As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless. At the same time, the protocol should also work efficiently and with low-latency on smartphones in a high-throughput scenario. Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput. In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by a conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [28], Cryptocurrencies [30, 19]) are blocklace-based, it would be very beneficial and integrable to also have a (v) blocklace-based consensus protocol. This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [16] and its extensions and improvements [4, 14].
We note that a blocklace [3] is not “just a DAG” the same way a blockchain is not “just a sequence”. In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$. An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer of $b^{\prime}$. Thus, similarly to a blockchain, the blocklace is:
  1. 1.
    Tamper-proof: A bit in a block cannot be changed without this being detected;
  2. 2.
    Non-repudiable: The creator of a block cannot deny having created it;
  3. 3.
    Acyclic: Even Byzantine agents cannot create blocks that form/close a cycle.
However, unlike the blockchain, which is extended competitively, the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other. In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution. Namely the intention is that the parties to a social contract will run the contract, as well as its underlying consensus protocol, on their smartphones. The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult. In particular, opening a direct TCP connection between two phones is either impossible, or impossible without the help of a proxy server, and even if established, it has to be re-established whenever one of the phones changes their IP address. Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations. Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (vi) UDP-ready.
The final requirement for a protocol for grassroots platforms is, naturally, that the protocol be (iv) grassroots. Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$. Any permissioned consensus protocol satisfies that. (ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$. For this to happen in a permissioned consensus protocol, the set of agents $P$ need to be “reconfigured” to include agents in $P^{\prime}\setminus P$. The question of reconfiguration has been studied extensively for decades [18, 10, 2, 34]. While the original informal exposition of the problem in the Paxos paper was with respect to a self-governed parliament [18], subsequent works in the context of permissioned consensus assume that the initial ‘permissioned’ agents are determined by an external authority, and so is their ‘reconfiguration’. Proof-of-Stake protocols such as Ethereum 2.0 offer a stake-based approach to reconfigure the participating agents on every epoch [17, 6].
A grassroots system cannot have external authorities and hence protocol reconfiguration must be entirely under the control of the very same agents that execute the protocol, as envisioned in the Paxos paper [18]. Furthermore, to be egalitarian rather than plutocratic, agents should have equal power in determining reconfiguration decisions. Specifically, in the grassroots context the challenge relates to the consensus protocol being executed by a self-governed community (the parties to the social contract running on top of the consensus protocol), allowing them to add or remove members at will, as well decide to change other aspects of its operation, for example the supermajority required to finalize a decision. To address that, we extend the protocol with a constitution that specifies the agents that execute the protocol as well as two key protocol parameters—the supermajority $\sigma$ by which decisions, including constitutional amendment decisions, are binding, and a time $\Delta$ estimating the unknown least upper bound on network delay during synchrony, $\delta$.
To summarize, the grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots. The Grassroots Consensus protocol was conceived to address all these requirements.
Paper structure. Section 2 presents the Grassroots Consensus protocol assuming a reliable network. The protocol is presented assuming a prevailing constitution that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony. Section 3 proves the safety and liveness of Grassroots Consensus. Section 4 extends the protocol for eventually-reliable networks, making it UDP-ready. Section 5 proves that the Grassroots Consensus protocol is indeed grassroots. Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present a distinct democratic constitutional amendment process for each parameter. Section 7 concludes.
2 Grassroots Consensus
2.1 Blocklace Preliminaries
We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$. We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$. While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it.
A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$, with $b$ called initial if $H=\emptyset$. An encoding of such a triple $b=(h,x,H)$ via a sequence of bits in an agreed-upon form is referred to as a well-formed block. We assume the hash function is collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$. Furthermore, the method of signature allows the public key $p$ to be recovered from a signature by $p$.
A block $b$ observes itself, any block $b^{\prime}$ pointed to by $b$, and any block observed by $b^{\prime}$. The ‘observes’ relation induces a partial order on any set of blocks. Two blocks that do not observe each other are conflicting, and if by the same agent they are equivocating. A block $b$ approves a block $b^{\prime}$ if $b$ observes $b^{\prime}$ and does not observe any block equivocating with $b^{\prime}$, and a set of blocks $B$ approves $b^{\prime}$ if every $b\in B$ approves $b^{\prime}$.
The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$. If a block variable $b$ is undefined then we deem its closure $[b]:=\emptyset$ to be empty rather than undefined. For a set of blocks $B$ the closure $[B]$ is defined by $[B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$. A set of blocks $B$ is closed if $[b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$. A blocklace is a closed set of blocks.
To help avoid confusion, we use $B$ below to denote a closed set of blocks (a blocklace) and $D$ to denote any set of blocks, not necessarily closed. Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$. Thus, a blocklace is a set of blocks with no dangling pointers. A block $b\in D$ is a tip of $D$ if no other block $b^{\prime}\neq b\in D$ observes $b$.
The depth $d(b)$ of a block $b$ is $1$ if $b$ is initial else $d(b)=d(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$. The depth $d(B)$ of a set of blocks $B$ is defined by $d(B):=\max_{b\in B}d(b)$. A round $r\geq 1$ in $B$ is the set of all blocks $b\in B$ of depth $d(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with depth $d(b)\leq r$, equivalently all rounds in $B$ up to and including $r$. Note that, by definition, all blocks of the same round do not observe each other.
2.2 Protocol Concepts
Models. We consider two models. The first is the standard model of eventual synchrony (aka partial synchrony) that assumes a reliable network in which every message sent among correct agents is eventually received. Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$. In particular, all messages sent any time before GST arrive by time GST+$\delta$.
In the second model the assumption of a reliable network is relaxed. It assumes an eventually-reliable network, in which any message sent infinitely often between correct agents eventually arrives, and a message sent after GST between correct agents arrives within $\delta$. This is the model assumed in the original Paxos paper [18] and many sequels. We refer to this model as the UDP-ready model.
Parameters. The protocol employs three key parameters $(P,\sigma,\Delta)$:
  1. 1.
    $P\subset\Pi$, a set of agents,
  2. 2.
    $\frac{1}{2}\leq\sigma<1$, a fraction, used to specify a supermajority among the agents
  3. 3.
    $\Delta>0$, a presumed upper-bound on message delay after GST, known to the agents.
We refer to the parameters $(P,\sigma,\Delta)$ employed by the protocol for a given block $b$ or blocklace $B$, with values as above, as the prevailing constitution in $b$, resp. $B$; when the value of $\Delta$ is immaterial we may refer to the prevailing constitution $(P,\sigma)$. We begin by assuming the prevailing constitution to consist of arbitrary constant values for all blocks during the protocol run, as is usually the case. Later, in Section 6, we augment the protocol with constitutional amendment blocks by which the prevailing constitution may be amended.
Supermajorities. Given agents $P\subset\Pi$, $|P|=n$, and $\frac{1}{2}\leq\sigma<1$, a $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$, or, equivalently, that $\frac{|Q|}{|P|}>\sigma$. A set of blocks $D$ is referred to as a $\sigma$-supermajority if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $D$ includes at least one $q$-block.
Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$. Given $f$, we say that a blocklace $B$ is $f$-safe if it includes equivocations by at most $f$ agents, a block $b$ is $f$-safe if $[b]$ is $f$-safe, and define $\sigma:=\frac{n+f}{2n}$. For example, if $f=0$ then $\sigma=\frac{1}{2}$; if $f=\frac{1}{2}n$ then $\sigma=\frac{3}{4}$; if $n=3f+1$ (the standard assumption), then $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
Under this assumption, a $\sigma$-supermajority includes a majority of the correct agents, and hence the intersection of any two $\sigma$-supermajorities includes at least one correct agent. Typically, permissioned consensus protocols refer to “blocks by at least $n-f$ agents among $P$”, which is equivalent to “a $\sigma$-supermajority among $P$” in the standard case of $n=3f+1$.
Waves. A blocklace is seen as a sequence of waves, where the $k^{th}$-wave in $B$ , $k\geq 1$, consists of three consecutive rounds $r$, $r+1$, $r+2$ in $B$, $r=3(k-1)+1$, referred to as the first, second, and third round of the wave.
For a given $P\subset\Pi$, the partial function leaderP assigns a leader $p\in P$ for the first round of every wave, where a first-round $p$-block $b$ with prevailing constitution $(P,\sigma)$ is a leader block if $p=\textit{leader}_{P}(d(b))$. We assume the leader function to be fair in that for every wave and any $p\in P$ there is a subsequent wave for which $p$ is the leader of its first round.
The consensus protocol progresses in waves. In each wave, a first round block is a candidate for finality in two cases—if it is a leader block or the only block in its round. These two possibilities typically arise in high- and low-throughput scenarios, respectively. During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with others, all agents produce second and third round blocks that finalize $b$. During high throughput, conflicting blocks may be produced and the leader function is used to break symmetry; leader blocks are finalized and induce the total ordering of all blocks.
A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows. Note that the next two definitions are mutually-recursive:
Definition 2.1 (Endorse).
Let $b$ be a first-round block with prevailing constitution $(P,\sigma)$ and $b^{\prime}$ a second-round block of the same wave in $B$. Then $b^{\prime}$ endorses $b$ if the previous wave is:
  1. 1.
    quiescent and $b$ is the only block in its round approved by $b^{\prime}$, or
  2. 2.
    non-quiescent and $b$ is a leader block approved by $b^{\prime}$.
A set of blocks $D$ endorses $b$ if every $b^{\prime}\in D$ endorses $b$.
A first-round block $b$ is final if the third round of the same wave has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally (text in parenthesis is explanatory, the definition is valid also without it):
Definition 2.2 (Ratified, Final, Ordered, Quiescent).
Given a blocklace $B$, a block $b\in B$ with prevailing constitution $(P,\sigma)$ is:
  1. 1.
    ratified by a block $b^{\prime}$ if $b^{\prime}$ approves a $\sigma$-supermajority that endorses $b$;
  2. 2.
    final in $B$ if there is a $\sigma$-supermajority in the third-round of $b$’s wave in $B$, in which each block ratifies $b$, referred to as a finalizing $\sigma$-supermajority.
A wave in such a blocklace $B$ is:
  1. 1.
    finalizing if it includes a (first-round) final block $b$; and
  2. 2.
    quiescent if in addition all the wave’s blocks except perhaps $b$ are empty and $b$ does not conflict with any block in $B$.
As an initial condition, we refer to the (non-existent) wave that ends in round 00 as quiescent. A block $b$ is slow if it is both finalized by a quiescent wave and follows a quiescent wave.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$. Note, however, that quiescence is not monotonic, both because conflicting blocks may be added to a round, and because more rounds can be added to the blocklace.
The next observation asserts that a wave’s final block is unique. {observation}[Uniqueness] A wave can have at most one ratified or final block.
Quiescence, low- and high-throughput scenarios. The protocol is initially quiescent. It begins operations and continues to do so as long as agents have payloads to send. If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode. It returns to being quiescent following a quiescent wave.
During low-throughput, in the good-case scenario, all waves are quiescent. A wave commences when some agent has a new transaction. This agent issues a (first-round) non-empty block, and all agents follow by issuing two rounds of blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final. The new wave is also quiescent provided all second- and third-round blocks are empty. The agent that issued a finalized, non-empty, first-round block in this scenario is referred to as a spontaneous leader, in contrast with the formal leader assigned by the leaderP function.
During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above, except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty (contain transactions to be ordered by the next-wave’s final leader block). The good-case scenarios are summarized in Figure 1.
Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own. Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block. And if during low-throughput there are multiple agents issuing first-round blocks, then it is possible that none of the first round blocks will become final. In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
To formalize these progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
Definition 2.3 (Advanced Round).
Given a blocklace $B$ with prevailing constitution $(P,\sigma)$, a round $r>0$ in $B$ is advanced if $r$ is a:
  1. 1.
    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a leader block or (iii) follows a quiescent wave and includes at least one block, or a
  2. 2.
    second or third round that includes a $\sigma$-supermajority of blocks among $P$.
Round 00 is trivially advanced in all blocklaces.
Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally, an ordinary block $b$ is valid if it is well-formed and round $d(b)-1$ is advanced in $[b]$.
Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block.
In the low-throughput scenario, the previous wave ending in round $r$ is quiescent, the spontaneous leader block (red) may be produced at any time after the previous round $r$ has advanced, and all other blocks are empty, thus the wave $w$ is also quiescent.
In the high-throughput scenario, the formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced, and all (or many) of the blocks are non-empty.
Figure 1: A good-case wave.
Dissemination. The protocol uses three types of dissemination-inducing blocks: ack, nack, and nudge, defined below; other blocks are referred to as ordinary.
Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
The Grassroots Consensus protocol we present in Section 2.3 assumes reliable links and uses nack and nudge blocks for dissemination. Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Grassroots Consensus for this model achieves $O(n)$ amortized communication complexity in the good case during high-throughput. The UDP-ready protocol (given in Section 4) uses also ack blocks for dissemination.
nack blocks are used to achieve reliable broadcast of ordinary blocks. They are needed in order to address partial dissemination by Byzantine agents. For example, consider a Byzantine agent that sends a block $b$ to correct agent $p$ but not to another correct agent $q$, and $p$ adds $b$ into its blocklace and sends a block $b^{\prime}$ that depends on $b$. Then when $q$ receives $b^{\prime}$, it cannot add it to its blocklace because its predecessor $b$ is missing. To this end, $q$ sends a nack block to $p$ specifying the blocks it is missing.
Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony, even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$. We therefore introduce a delay before sending a nack. Thus, in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no further communication is needed.
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $[b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. While obviously $p$ cannot know all the blocks it does not know, it can know the boundary of its knowledge—where $p$’s ignorance starts in relation to $b$. These are the tips of $[b]\setminus(B\cup\{b\})$. Note that if no block pointed to by $b$ is in $B$, then the pointers in $b$ point exactly to tips in $[b]\setminus(B\cup\{b\})$. Hence the following definition:
Definition 2.4 (nack-block).
Given a blocklace $B$, a nack-block for a block $b=(h,x,H)$ is the block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $[b]\setminus(B\cup\{b\})$.
Note that as defined above, $[b^{\prime}]\subseteq[b]$. Also note that, given a block $b=(h,x,H)$ and a blocklace $B$, the tips of $[b]\setminus(B\cup\{b\})$ are observed directly from $b$: These are the pointers in $H$ that are dangling in $B$, namely point to blocks not in $B$. These dangling pointers attest to the ‘knowledge gap’ of the recipient of $b$, namely, which blocks need to be added to $B$ so that $b$ will be in order following $B$. If there is no such gap, namely $[b]\subseteq B\cup\{b\}$, then the nack-block for $b$ and $B$ is undefined.
In the reliable network model, the reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block is in fact an ack-block, as it attests to all the blocks known to $p$ at the time of its creation; ack-blocks will be added in the UDP-ready protocol described in Section 4, when eventually-reliable networks are considered.
nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate progress using nudge-blocks.
Definition 2.5 (nudge-Block).
Given a blocklace $B$, a nudge-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{nudge}$ and pointers $H$ to the blocks of depth $r$ in $B$.
If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks, $p$ sends the leader a nudge-block pointing to the $r$ blocks in its blocklace. If the leader indeed misses any of these blocks, which after GST may happen due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
2.3 The Protocol
Here we present Grassroots Consensus for eventual synchrony with a reliable network. We prove the protocol’s safety and liveness in Section 3, and present its UDP-ready extension in Section 4. In this section, we present the protocol using arbitrarily-fixed values for the parameters $(P,\sigma,\Delta)$. These can be changed using constitutional amendments, discussed in Section 6.
In the Grassroots Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. The protocol is event-driven as described in Algorithm 1.
Overview. Each three consecutive rounds are grouped into a wave, which aims to finalize one block of its first round. Similarly to Morpheus [20], the protocol has two modes of operation, high-throughput and low-throughput, but here they are implicit— depending on whether the preceding wave is quiescent. In high-throughput mode, in the good-case scenario, only the leader sends a first round block. If the leader is faulty or does not communicate with other agents in a timely manner (bad-case scenario), other agents time-out on the leader and send blocks of their own. Agents wait either for a first-round leader block or for a supermajority of blocks in a round before producing a block in the next round.
Low-throughput is manifested by a quiescent wave, which finalizes its first-round spontaneous leader using two rounds of empty blocks. Following a quiescent wave, any agent may initiate a new wave by issuing a non-empty block. If two such blocks collide the protocol enters high throughput mode until, if ever, it quiesces again.
Initialize: $D\leftarrow\phi$, $B\leftarrow\phi$, $\emph{payload}\leftarrow\bot$.
Let $r$ always be the maximal advanced round in $B$, which is initially 0.
Receive: Upon receipt of $q$-block $b$, if $b$ a nack-block then judiciously send $[b]$ to $q$,
else if $b$ a nudge-block such that $p=\emph{leader}(d(b))$ and $[b]\not\subseteq B\cup\{b\}$ then send once a nack-block for $b$ to $q$,
else if $b$ is valid then add $b$ to $D$.
Accept & Nack: If $[b]\subseteq B\cup\{b\}$ for some $b\in D$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
else if a $q$-block $b^{\prime}$ was added to $D$ more than $\Delta$ time ago then send once a nack-block for $b^{\prime}$ to $q$.
Issue: Issue once a new block if $r+1$ is a second or third round or a first round that follows a wave that is: 1. quiescent (or $r=0$) and $\emph{payload}\neq\bot$, or 2. non-quiescent and (i) $p=\emph{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Nudge: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once to $\emph{leader}_{P}(r+1)$ a nudge-block for round $r+1$. Output: If $B$ contains a final block $b$ of depth greater than the previous final block (if any) then output every block in $\tau(b)$ unless it has already been output.
issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) a nack-block was received for $b^{\prime}$, or (iii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
Algorithm 1 Grassroots consensus protocol, code for agent $p\in P$, given $(P,\sigma,\Delta)$.
Delays. The protocol makes judicious use of delays, to ensure both liveness and efficiency. Here is the rationale for the various protocol delays, all expressed as multiples of $\Delta$:
$\Delta$ for issuing nack-blocks. In the good case we wish no nack-blocks to be sent at all. Agents wait $\Delta$ before complaining, so if a correct agent sends a block $b$ to $p$ and $q$ at the same time, $q$ may receive $b$ up to $\Delta$ sooner than $q$ and send a block $b^{\prime}$ that depends on $b$ to $p$; but $p$ will receive $b$ within $\Delta$ from receiving $b^{\prime}$.This delay is an optimization to avoid sending nack-blocks in the good case; eliminating it does not hamper safety or liveness.
$2\Delta$ for nudging the ensuing round’s leader. In the good case whence all agents are correct after GST, whenever an agent receives a block, all other agents receive it within $\Delta$ time. Thus, once a third round is advanced at the blocklace of a correct non-leader agent $p$, $p$ expects to receive a leader block of the ensuing (first) round within $2\Delta$. If this does not happen, this could be because (i) it is before GST, (ii) the leader is incorrect, or (iii) the leader has not seen yet that the previous round is advanced. This last third case can happen after GST even if the leader is correct, due to Byzantine partial dissemination of third-round blocks. Since $p$ knows that the previous third round is advanced, it sends a nudge-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing. This $2\Delta$ delay is an optimization to avoid spurious nudges in the good case; eliminating it does not hamper safety or liveness.
$9\Delta$ for an unresponsive leader. Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, t all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized. But because $p_{\ell}$ may be faulty, they need to eventually time-out and stop waiting. To determine the time-out, we consider the worst-case latencies as follows: Assume a third round $r$ is advanced at $p$ at time $t$. If $p$ does not receive a leader block by $t+2\Delta$, it nudges $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks. This exchange takes at most $3\Delta$, so by time $t+5\Delta$, $p_{\ell}$ issues a leader block $b$ for round $r+1$, which is received by $p$ by time $t+6\Delta$. Upon receipt of $b$, $p$ might be missing some blocks in $[b]$ and wait $\Delta$ before sending a nack and getting them from $q$ by time $t+9\Delta$. If $p$ does not accept any leader-block by $t+9\Delta$, it concludes that either $p_{\ell}$ is faulty or GST has not arrived yet, and sends a first-round block of its own in order to allow the protocol to proceed to the next round.
Protocol walk-through. Incoming blocks are inserted to $D$; if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $B$.
The variable $r$ always holds the highest advanced round in $B$. If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (initially and following a quiescent wave), $p$ issues a block only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the leader, or (ii) it times-out on the leader.
An agent $p$ always issues blocks that follow the latest advanced round $p$ saw. This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced. This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
After GST, whenever there is a wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block. Therefore, no agent can produce a valid second round block that does not endorse the leader block. So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized.
Ordering with $\tau$. Once the protocol finalizes a block $b$, it orders the predecessors of $b$ that have not been ordered yet using the function $\tau$. The definition of $\tau$ is similar to Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks.
The output of $\tau(b)$ is a sequence that includes all blocks observed by $b$, with equivocating blocks excluded. In Section 3 we prove that $\tau$ is monotonic and safe. The monotonicity of $\tau$ ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor. Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any. The safety of $\tau$ means that blocks are ordered the same way at all correct agents.
In anticipation of the protocol being constitutional, we already assume the notion of the prevailing constitution. For now, assume $(P,\sigma,\Delta)$ is fixed arbitrarily for all blocks. In Section 6, we define the constitution prevailing in a block, and have each call (initial and recursive) to $\tau$ with a block $b$ employ the parameters of the constitution prevailing in $b$.
Definition 2.6 ($\tau$).
We assume a fixed topological sort function $\textit{xsort}(b,B)$ (exclude and sort) that takes a block $b$ and a blocklace $B$, and returns a sequence that preserves the ‘observes’ partial order and includes all the blocks in $B$ that are approved by $b$. The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
$\tau(b):=\begin{cases}\tau(b^{\prime})\cdot\textit{xsort}(b,[b]\setminus[b^{\prime}])\text{\ \ if $b^{\prime}$ is the maximal-depth block ratified in $[b]$}\\ \ \ \ \ \ \ \ \ \ \textit{according to the constitution prevailing in $b$}\\ \textit{xsort}(b,[b])\text{\ \ \ \ \ if no such $b^{\prime}$ exists }\\ \end{cases}$
When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$. Note that $b^{\prime}$ is not necessarily final in $[b]$, nor final in the closure of the final block with which $\tau$ was called initially. In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
Latency and complexity. Regarding latency, we note that after GST and in the good case a (formal or spontaneous) leader block is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
Regarding communication complexity, during low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus [20], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$. A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, should carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.k
During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
3 Safety and Liveness of Grassroots Consensus
Here we prove the safety and liveness of the Grassroots Consensus protocol.
We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
Definition 3.1 (Safety and Liveness of an Ordering Consensus Protocol).
An ordering consensus protocol is:
Safe if output sequences of correct agents are always consistent.
Live if every non-empty block issued by a correct agent is eventually included in the output of every correct agent.
Next we prove safety and liveness of the basic Grassroots Consensus protocol.
3.1 Safety
Theorem 3.2.
Grassroots Consensus is safe.
To prove the theorem, we follow a condensed and simplified proof of the safety of Cordial Miners.
First, observe that Grassroots Consensus produces only $f$-safe blocks; if a block created by the protocol is not $f$-safe it means that its closure includes equivocations by more than $f$ agents, a contradiction. For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
Definition 3.3 ($\tau$ Safety).
A blocklace $B$ is $\tau$-safe if every final block in $B$ is ratified by every higher-depth block ratified by some block in $B$.
Proposition 3.4.
An f-safe blocklace is $\tau$-safe.
Proof 3.5.
Let $B$ be an $f$-safe blocklace, let $b$ be a block of wave $k$ final in $B$ and $b^{\prime}$ a block of a later wave $k^{\prime}>k$ ratified by some block in $B$, which means that $b^{\prime}$ is endorsed by a $\sigma$-supermajority in $B$. We have to show that $b$ is ratified by $b^{\prime}$.
Since $b^{\prime}$ is ratified by some block it means that it is valid, since correct agents do not include invalid blocks in their blocklace, let alone endorse them, and a $\sigma$-supermajority includes (a majority of) correct agents. Hence, being a first-round block of a wave, $b^{\prime}$ must observe a $\sigma$-supermajority of the last round of the previous wave, let’s call it the blue supermajority.
Then the blue supermajority and the finalizing $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $k^{\prime}$ happens to be $k+1$), which in turn must observe a supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
Proposition 3.6 ($\tau$ Monotonicity).
$\tau$ is monotonic wrt $\supset$ over $f$-safe blocklaces.
Proof 3.7.
We have to show that if $B\subset B^{\prime}$ for $f$-safe blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$. By Observation 2.2, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation 2.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their round number. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition 3.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
{observation}
If the union $B\cup B^{\prime}$ of two blocklaces is $f$-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
Proof 3.8.
By monotonicity of $\tau$ over safe blocklaces, $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
Proof 3.9 (Proof of Theorem 3.2).
Since in a run of Grassroots Consensus there are at most $f$ faulty agents in total, the union of the blocklaces of two correct agents is always $f$-safe. Hence, by Observation 3.1, the outputs of any two correct agents running Grassroots Consensus are always consistent.
3.2 Liveness
Next, we discuss liveness. We first prove that Grassroots Consensus disseminates all correct agents’ blocks and the blocks they depend on.
Proposition 3.10 (Dissemination).
If a block $b$ is issued by a correct agent in Grassroots Consensus then $[b]$ is eventually received by all correct agents.
Proof 3.11.
By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$. Since by assumption $b$ (and possibly other blocks in $[b_{p}]$) have not been received by $q$, then upon receipt of $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$. Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $[b_{q}]$, which includes $b$, to $q$. Since the network is reliable, $q$ will eventually received $b$, a contradiction.
The following observation follows directly from the definition of block validity:
{observation}
[Sequential Advance] If round $r$ is advanced in a blocklace $B$, then every round $0<r^{\prime}<r$ is also advanced in $B$.
We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress. Because Grassroots Consensus is quiescent, agents do not have to generate infinitely many blocks. But if some agent issues a block in a round, that round eventually advances.
Proposition 3.12 (Advance).
If a correct agent $p$ sends a block in some round $r$ in Grassroots Consensus, then $r$ eventually becomes advanced at every correct agent.
Proof 3.13.
Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, $r-1$ is advanced in $p_{1}$’s blocklace.
Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$. By Observation 3.11, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace. By Proposition 3.10, this means that no correct agent sends blocks in rounds $>r$.
Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition 3.10 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
By Proposition 3.10, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round $r-1$ is advanced in $p_{3}$’s blocklace.
In case $r$ is a first round following a quiescent wave at $p_{3}$, $r$ advances at $p_{3}$ upon receipt of $b$, and $p_{3}$ issues a round $r+1$ block, a contradiction.
Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction.
We next prove liveness for finite runs.
Proposition 3.14 (Finite implies quiescence).
Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
Proof 3.15.
Let $r$ be the last advanced round in $B$. (Note that round 00 is by definition advanced, so $r$ is always defined). Assume by way of contradiction that $B$ is not quiescent, then $p$ eventually issues a round $r+1$ block $b$. By Proposition 3.12, $r+1$ eventually becomes advanced at $p$. A contradiction.
Proposition 3.16 (Liveness in finite runs).
Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all non-empty ordinary blocks sent by correct agents in this run are ordered in $B$.
Proof 3.17.
By Proposition 3.10, $B$ includes all ordinary blocks sent by correct agents. By Proposition 3.14, $B$ is quiescent. This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty. Because $b$ is final, $p$ calls $\tau(b)$. Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$.
Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalizing. Because quiescence is not monotonic, we consider here only waves $k$ for which the preceding wave, $k-1$, is not quiescent at any correct agent at any time.
Proposition 3.18 (First round progress after GST).
Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round $r-1$ advances at time $t$>GST, and assume that the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time. Then $p$ sends a block in round $r$ if and only if it is the round’s leader. Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
Proof 3.19.
Let $p_{\ell}$ be the correct leader of round $r$. If $p=p_{\ell}$ then once $r-1$ is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ a nudge-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$. Because round $r-1$ is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$. Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block.
Below, we say that a round $r$ advances at time $t$ if there is some correct agent $p$ such that round $r$ is advanced in $p$’s blocklace at time $t$ and $r$ is not advanced at any correct agent before time $t$.
Proposition 3.20 (Progress after GST).
Let $r$ be a first round with a correct leader s.t. round $r-1$ advances after GST and the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time. Then a round $r$ leader block is finalized in its wave at all correct agents.
Proof 3.21.
By Proposition 3.18, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$. This means that it is impossible to create a valid round $r+1$ block that does not endorse the leader, as such a block would have to point to a super-majority of round $r$ blocks excluding the leader block. Therefore, every valid round $r+2$ block ratifies the leader block, and the block is finalized in round $r+2$ as soon as it advances at any correct agent.
Next, we consider runs with infinitely many quiescent waves.
Proposition 3.22 (Infinite quiescent waves).
Consider a run of Grassroots Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. Then all the blocks of correct agents are ordered by all agents in this run.
Proof 3.23.
Let $B$ be the (infinite) union of all blocklaces in the run. Let $b_{1}$ be a block of a correct agent in $B$. By Proposition 3.10, $b_{1}$ is eventually included in all correct agents’ blocklaces. Therefore, it is approved by all correct agents’ blocks from some wave onward. Because every quiescent wave is finalizing, and finalization is monotonic, there are infinitely many finalizing waves in $B$. Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$. Once $b_{2}$ is final at some agent’s blocklace, $b_{1}$ is ordered at that agent. We will show that $b_{2}$ is finalized at all agents.
Observe that since $b_{2}$ is finalized in wave $k$, it is approved by a super-majority in wave $k$. Because every valid first round block approves a super-majority in the preceding round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
Consider a finalizing wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$. Then $b_{3}$ is endorsed by a super-majority in the second round. Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 3.10). If $b_{3}$ is not a leader block, then to be validly endorsed, $[b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
By the same token, if any non-leader first round block $b$ is endorsed by a valid block of any agent in the second round of wave $k+1$, then $[b]$ is quiescent. This, in turn, means that every third round block in the wave either ratifies the leader block or observes a block whose closure is quiescent.
So when the third round of wave $k+1$ advances, every correct agent either observes a super-majority of blocks that ratify the leader block or observes a block whose closure is quiescent. In the former case the leader block is finalized, ordering all preceding ordinary blocks including $b_{2}$, and in the latter $b_{2}$ is already final in the block’s closure.
We are now ready to prove the liveness theorem.
Theorem 3.24.
Grassroots Consensus is live.
Proof 3.25.
Let $b$ be a non-empty ordinary $p$-block produced by a correct agent $p$ in a run of Grassroots Consensus. By Proposition 3.10, every correct agent receives and incorporates $b$ in its local blocklace.
We consider three cases: First, assume the run is finite. By Proposition 3.16, all non-empty ordinary blocks sent by correct agents in this run are ordered at all correct agents, and the theorem holds. Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time. By Proposition 3.22, $b$ is ordered.
Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time. Let $k$ be a wave whose last round advances after GST and after which no wave is quiescent at any correct agent’s blocklace. By fairness of leader selection, there are inifinitely many waves $k^{\prime}>k$ with a correct leader. By Proposition 3.20, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
4 UDP-Ready Grassroots Consensus for Eventually-Reliable Networks
Grassroots platforms and protocols are geared for smartphone-based serverless implementation, and hence depend on smartphone-to-smartphone communication. Smartphones often reside behind NATs and Firewalls, which in turn are programmed with policies that make smartphones finding each other difficult. To overcome this, servers with known IP addresses (or domain names), running the STUN protocol help devices behind NATs find each other. However, once they do, establishing a direct TCP connection is difficult, and hence direct phone-to-phone communication typically employs UDP.
Grassroots Consensus (Algorithm 1) was specified assuming a reliable network. Here, we relax the assumption to eventually-reliable networks, and extend the protocol to be UDP-ready, by adding ack-blocks and block resending. This protocol resends only ordinary blocks, as the reliable communication of dissemination-inducing blocks (nack, ack, and nudge) before GST is not required. Note that the protocol is very different from the trivial solution of “implementing TCP upon UDP”, as during correct execution the same block can arrive via multiple paths and blocks may arrive in many possible orders. The protocol does not need to to impose an ordering on agent-to-agent block communication.
Definition 4.1 (ack-block).
Given a block $b=(h,x,H)$, an ack-block for $b$ is the block $b^{\prime}=(h^{\prime},(\textsc{ack},h),\emptyset)$.
The changes to make Grassroots Consensus UDP-ready are adding ack to the Receive rule, adding resending to the nack-rule, a new Resend rule, and revising the notion of judiciously send, as detailed in Algorithm 2 below (additions to Algorithm 1 are highlighted). In Algorithm 2, $k\geq 2$ is a parameter chosen to taste. In principle, it could be added as the fourth parameter of the constitution and amended similarly to $\Delta$.
Receive: Upon receipt of $q$-block $b$, if $b$ a nack-block then judiciously send $[b]$ to $q$,
else if $b$ a nudge-block such that $p=\emph{leader}(d(b))$ and $[b]\not\subseteq B\cup\{b\}$ then send once a nack-block for $b$ to $q$,
else if $b$ is valid then add $b$ to $D$ and send an ack-block for $b$ to $q$.
Accept & Nack: If $[b]\subseteq B\cup\{b\}$ for some $b\in D$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$, else if a $q$-block $b$ was added to $D$ more than $\Delta$ time ago, and no nack-block for $b$ was sent during the last $k\Delta$, then send a nack-block for $b$ to $q$. Resend: If $k\Delta$ has passed since the most-recent ordinary $p$-block $b$ was sent to $q$, no ack-block or nack-block was received for $b$, and $b\notin[b_{q}]$ for any $q$-block $b_{q}\in B\cup D$, then resend $b$ to $q$.
judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) an ack-block or a nack-block was received for $b^{\prime}$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
Algorithm 2 UDP-Ready Extension to Grassroots Consensus.
Mobile devices. Smartphones are mobile, and when they move their IP address may change. Hence, even if a phone-to-phone connection is established, it has to be re-established every time one of the phones changes their IP address. Blocklace dissemination can support mobile agents recovering each other’s IP address via a joint stationary friend [28]. The problem is already present at protocols of lower-level than consensus such as the grassroots social graph [31] and the blocklace-based solution presented there [28] is applicable here as well.
Safety and liveness. The UDP-ready extension (Algorithm 2) to Grassroots Consensus only relates to dissemination-inducing blocks (an ack-block and its handing) and does not affect ordinary blocks. Hence the definitions, propositions, and proofs regarding the safety of Grassroots Consensus—all stated in terms of ordinary blocks only—remain valid verbatim. Safety of the UDP-ready extension is therefore a corollary of Theorem 3.2.
Regarding liveness, prior to GST ordinary blocks issued by correct agents are resent to correct agents until received. Thus, there is a time $t\geq\text{GST}$ by which every block sent by a correct agent priot to GST either has been received, or has been resent after GST but before $t$. Thus, any message sent before $t$ would be received by $t+\delta$. After GST, the two models under consideration are the same. Thus, the time $t$ in the UDP-ready protocol has the same properties as GST in the Grassroots Consensus protocol: Any message between correct agents, if sent before $t$ arrives by $t+\delta$, and if sent after $t$ arrives within $\delta$. Hence, the liveness proofs of Grassroots Consensus hold for the UDP-ready protocol, replacing GST by $t$ thus defined, with the liveness of the UDP-ready protocol being a corollary of Theorem 3.24.
5 Grassroots Consensus is Grassroots
We recall the formal definition of a grassroots protocol from reference [31], instantiate it informally (bypassing the formalities of the definition) to the constitutional Grassroots Consensus protocol, and then argue that the protocol is grassroots.
Definition 5.1 (Oblivious, Interactive, Grassroots).
A protocol $\mathcal{F}$ is: 1. oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, a run of the protocol over $P$ can proceed just the same in the presence of agents in $P^{\prime}$. 2. interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every run of the protocol over $P$, the protocol can proceed from any configuration $c$ in $r$ to interact with members in $P^{\prime}\setminus P$, namely produce a run that is not possible only over $P$. 3. grassroots if it is oblivious and interactive.
Note that the definition is stated in terms of any two populations $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$.
Being oblivious, in the context of Grassroots Consensus, means that if the prevailing populations of the protocol during a run are always included in $P$, then the presence of additional agents hanging around, those in $P^{\prime}\setminus P$, which do not participate in the protocol, should not interfere with the protocol or affect its behaviour. Clearly, this is the case with Grassroots Consensus, as its prevailing populations may always choose not to amend the constitution to include members outside $P$. Hence Grassroots Consensus satisfies the requirement of being oblivious.
Being interactive, in the context of Grassroots Consensus, means that if the prevailing populations of the protocol in some prefix of a run ending in a configuration $c$ are included in a set $P$, then the protocol has a computation starting from $c$ that interacts with members outside $P$ in a meaningful way, namely the protocol has a computation from $c$ that reaches a configuration $c^{\prime}$, in which members of $P^{\prime}\setminus P$ are included in the prevailing population, and thus $c^{\prime}$ cannot be reached if the prevailing populations are all restricted to $P$. Clearly, members of the prevailing population $P$, starting from configuration $c$, may amend the constitution to include in $P$ a new member $p^{\prime}\in(P^{\prime}\setminus P)$ and thus reach a configuration $c^{\prime}$ in which $p^{\prime}\notin P$ is a member. Hence Grassroots Consensus satisfies the requirement of being interactive. Together, these two properties imply that constitutional Grassroots Consensus is indeed grassroots. We thus conclude:
Theorem 5.2.
Grassroots Consensus is grassroots.
Note that the two requirements cannot even be expressed in the context of the basic (non-constitutional) Grassroots Consensus protocol, as the population $P$ is given and fixed. And that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
6 Constitutional Grassroots Consensus
The problem of reconfiguration of a consensus protocol—namely changing the set of its participants—has engaged the distributed computing research community for decades [34, 10, 2, 34], ever since it was first formulated in Lamport’s Paxos paper [18]. Here we explore the more general problem of constitutional amendment. Supporting constitutional amendments makes the protocol grassroots, as we formally prove in Section 5.
6.1 The Constitution
Generally, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself. The notion of a constitution and its amendment is different from the well-studied notion of reconfiguration in at least two respects:
  1. 1.
    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including of course the agents executing it but also the protocol itself and its parameters.
  2. 2.
    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to be coming “out of the blue”. Amending the constitution of a protocol, on the other hand, is done by the agents executing the protocol themselves, in accordance with the prevailing constitution’s constitutional amendment rule.
Here, we consider a restricted type of constitution $(P,\sigma,\Delta)$, with $\emptyset\subset P\subset\Pi$ being the population, $\frac{1}{2}\leq\sigma<1$ being a supermajority among the population the approval of which is needed to amend the constitution, which is also the supermajority used by the underlying consensus protocol, and $\Delta$ being the agreed-upon estimate of the prevailing population $P$ regarding the delay needed to ensure the liveness of the protocol.
The three components of the constitution affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony. Even more so, $P$, $\sigma$, and $\Delta$ are key components of Grassroots Consensus, given that its primary aim is to execute digital social contracts, as explained next. The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc. Clearly, they should be sovereign to accept new members and remove existing members, namely to amend $P$.
As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should $\sigma$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased. Adding new people to $P$, initially known to and trusted by only a subset of $P$, may require increasing $\sigma$, even if temporarily. The parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members are admitted with lower-capacity devices, $\Delta$ may need to be increased. In addition, technology advanced may allow decreasing $\Delta$. Thus, the prevailing population $P$ should be sovereign to amend the values of $\sigma$ and $\Delta$ as they see fit.
6.2 Democratic Constitutional Amendment
Here we describe meta-level democratic processes for deciding on amending each of the components of a constitution $(P,\sigma,\Delta)$.
(1) Amending $P$. Regarding amending $P$ to $P^{\prime}$, requiring $\sigma$-supermajorities among both $P$ and $P^{\prime}$ to approve the amendment has been proposed before [10], to avoid, among other problems, the ‘dead sailors problem’ (in which all of $P^{\prime}$ are in fact dead, resulting in a constitutional stalemate) presented in the original Paxos paper [18].
(2) Amending $\sigma$. Amending the supermajority by which a constitution can be amended has been studies by Abramowitz et al. [1], who showed that the $h$-rule (similar to the $h$-index of author citation metrics) is the only rule that satisfies certain intuitive and self-evident axioms. According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$. The value of $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which there is a $\sigma^{\prime}$-supermajority among the prevailing population $P$ who voted for a value greater or equal to $\sigma^{\prime}$. Else $\sigma$ is decreased to the minimal $\sigma^{\prime}$ for which there is a $\sigma$-supermajority among the prevailing population $P$ who voted for a value less than or equal to $\sigma^{\prime}$. Else $\sigma$ remains unchanged.
(3) Amending $\Delta$. The value of the parameter $\Delta$ should be increased or decreased only if the majority of the correct agents wish so. To this end, we use the Suppress Outer-f parameter update rule by Shahaf et al. [25] [^1] In reference [25] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion., as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified. The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good”.
If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median of the remaining votes $\Delta^{\prime}$ is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median of the remaining votes $\Delta^{\prime}$ is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$. Else $\Delta$ remains unchanged.
We note several desirable properties of the Suppress Outer-f rule: If all Byzantines wish to increase $\Delta$ to slow-down the protocol, and the correct agents do not, then $\Delta$ will not be increased. Similarly, if all Byzantines wish to decrease $\Delta$ to undermine the liveness of the protocol, and the correct agents do not, then $\Delta$ will not be decreased. If all correct agents unanimously wish to update $\Delta$ to $\Delta^{\prime}$ then $\Delta$ will change to $\Delta^{\prime}$ no matter what the Byzantine agents vote, provided $f<\frac{1}{3}n$. In particular, if the protocol is not making progress, it can be expected that all correct agents will vote to increase $\Delta$ and will succeed in doing so.
6.3 Grassroots Consensus with Constitutional Amendment
While we suggested specific meta-level democratic processes for amending the three components of the constitution, the safety and liveness of constitutional amendment does not depend on their specifics.
[…]
In a grassroots setup, the identifier of each instance of a grassroots consensus protocol should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents. This can be achieved, for example, by having the identifier Id of a protocol instance to consists of the founders $P_{1}$ together with the maximal serial number of any grassroots consensus protocol any founder has founded. This ensures that if $P_{1}$ has at least one correct agent, then Id would be unique.
Definition 6.1 (Valid Constitutional Amendment).
Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, $d_{i}$ is valid if $i=1$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid, and $Q_{i}$ is a $\max(\sigma_{i-1},\sigma_{i})$-supermajority in $P_{i-1}$ and a $\max(\sigma_{i-1},\sigma_{i})$-supermajority in $P_{i}$. The sequence is valid if all its members are valid. Two non-identical constitutional amendment decisions are conflicting if they have the same index.
Note that the two consecutive supermajorities are possibly overlapping, even identical. In the special case that only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed; if only $\sigma_{i}$ changes then a single $\max(\sigma_{i_{1}},\sigma_{i})$-supermajority among $P_{i}$ is needed; and if only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$ and a $\sigma_{i}$-supermajority among $P_{i}$ is needed.
This definition ensures that the $h$-rule is followed and that the ‘dead sailors’ scenario of Paxos (that the new $P^{\prime}$ does not exists) [18] has not happened.
We assume that correct agents do not sign invalid or two conflicting constitutional amendment decisions, and hence can make the following observation. {observation}[Safety of Constitutional Amendment] Given a valid sequence of constitutional amendment decisions $d_{1},\ldots d_{i}$, $i\geq 1$, if the number $f$ of faulty agents in $P_{i}$ among the total number of agents $n=|P_{i}|$ satisfies $\sigma_{i}\geq\frac{n+f}{2n}$, then the constitutional amendment process cannot extend this sequence with two conflicting valid constitutional amendment decisions.
A constitutional amendment decision is enacted by the consensus protocol via a block that includes it as payload. To ensure orderly transition among constitutions, we assume that a correct agent would not sign $d_{i+1}$ before observing that the amendment decision $d_{i}$ has been finalized by the underlying consensus protocol, namely that a block containing $d_{i}$ as payload has been finalized (as we explain below, only leader blocks carry constitutional amendments as payload). We exploit this in the consensus protocol by assuming that a $d_{i+1}$ amendment decision cannot occur as a new payload of a correct agent $p$ before a block with $d_{i}$ is final in the blocklace of $p$.
Definition 6.2 (Valid Constitutional Amendment Block).
A block $b$ is a valid constitutional amendment block if it is a valid initial block with an initial constitution $d_{1}$ as payload, in which case it is called a founding constitution block, or with a valid constitution $d_{i}$ as payload, $i>1$, and for which there is a valid constitutional block $b^{\prime}$ with $d_{i-1}$ as payload and $b^{\prime}$ is final in $[b]$.
Definition 6.3 (Constitutional, Founding, Prevailing, Pending).
A blocklace $B$ is constitutional if it has a sole founding constitution block, with its constitution referred to as the founding constitution of $B$. Given a constitutional blocklace $B$, the prevailing constitution of $B$ is the constitution of the most-recent final valid constitutional amendment block in $B$. The sole pending constitution of $B$, if any, is the constitution of the most-recent non-final valid constitutional amendment block in $B$, if there is one. When these terms apply to a block $b$ they refer to the blocklace $[b]$.
Importantly, we amend the definition of quiescence (Definition 2.2) to state that a blocklace with a pending constitution is not quiescent. Beyond that, Definitions 2.1, 2.2, and 2.3 are kept verbatim, with the prevailing constitution of each block interpreted as in Definition 6.3, and the $\sigma$-supermajority required for rounds to advance (in Definition 2.3) counted out of the prevailing constitution of the agent’s entire blockchain.
The Grassroots Consensus protocol remains the same except in the way it assigns payloads to blocks, which is changed to facilitate constitutional amendments and avoid conflicting decisions during the transition between constitutions. This change is effected without modifying the Grassroots Consensus protocol (Algorithm 1), but through constraining the payload variable while there are pending amendments, as described in Algorithm 3 – an agent that knows of a constitutional amendment $d_{i}$ that is not prevailing does not send the usual payload. If the agent is a round leader sending a leader block, it sets its payload to $d_{i}$, and otherwise it sets it to $\bot$.
Constitutional Amendment: If $p$ knows of a constitutional amendment decision $d$ not prevailing in $B$ then prior to issuing a block of round $r+1$, if $p$ is the leader of round $r+1$ then $\emph{payload}\leftarrow d$ else $\emph{payload}\leftarrow\bot$ (retaining the outstanding value of payload, if not $\bot$, for the future).
Algorithm 3 Grassroots Consensus with Constitutional Amendment.
The word knows in the Constitutional Amendment rule has a double meaning: Either $p$ knows that the meta-level process of constitutional amendment has produced a new valid constitutional amendment decision $d$ not yet prevailing in $B$, or $p$ knows that its blocklace $B$ includes a valid pending constitutional amendment block with payload $d$.
To ensure liveness, an correct agent $p$ that is removed by a constitution amendment $d_{i}$ (namely, $p$ is in $d_{i-1}$ and not in $d_{i}$) must continue to participate in the protocol until it observes a final leader block $b$ whose prevailing constitution is $d_{i}$. In particular, the constitution amendment $d_{i}$ is final in $[b]$.
Excluding faulty agents. One natural application of constitutional amendment is excluding exposed faulty agents, in particular equivocators or agents that produce invalid blocks. The protocol is amended so that a correct agent $p$ does not sign a constitutional amendment $i$ if $P_{i}$ includes an agent exposed as faulty in $p$’s blocklace $B$. As a block known to a correct agent is eventually known to every correct agent, every agent exposed as faulty by one correct agent will eventually be exposed as such by every correct agent, and all exposed faulty agents will eventually be excluded from the prevailing population.
7 Conclusions
The Grassroots Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts. Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve [33]. Even small community banks and cooperatives [30] may still be, at least initially, on the low-throughput side. As communities federate to form larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow. Until then, simulating the various scenarios would be productive future work.

### v2

@@
 Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more.
-Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a small assembly elected by sortition [12, 33] among its members.
+Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a similarly-small assembly elected by sortition [20, 46] among its members.
 It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community.
-Such a consensus protocol should satisfy all the requirements listed below.
+Such a consensus protocol should satisfy all the requirements explained next.
 Requirements of Grassroots Consensus.
 Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency.
-Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued one at a time—the low-throughput scenario—finalize each quickly and efficiently.
+Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued at low-throughput, finalize each quickly and efficiently.
 As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless.
@@
 Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput.
-In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by a conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
-A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [28], Cryptocurrencies [30, 19]) are blocklace-based, it would be very beneficial and integrable to also have a (v) blocklace-based consensus protocol.
+In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by an overly-conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
+A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [41], Cryptocurrencies [43, 27]) are blocklace-based, it would be very beneficial and integrable to also have a (iv) blocklace-based consensus protocol.
 This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [24] and its extensions and improvements [5, 22].
@@
 In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$.
-An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer of $b^{\prime}$.
+An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer to $b^{\prime}$.
 Thus, similarly to a blockchain, the blocklace is:
@@
     Acyclic: Even Byzantine agents cannot create blocks that form/close a cycle.
-However, unlike the blockchain, which is extended competitively, the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other.
+However, unlike the blockchain, which is extended competitively, with the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other.
 In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
 The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution.
-Namely the intention is that the parties to a social contract will run the contract, as well as its underlying consensus protocol, on their smartphones.
+Namely, the intention is that the parties to a social contract will run both the contract and its underlying consensus protocol on their smartphones.
 The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult.
@@
 Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations.
-Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (vi) UDP-ready.
-The final requirement for a protocol for grassroots platforms is, naturally, that the protocol be (iv) grassroots.
+Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (v) UDP-ready.
+The final sixth requirement for a protocol for grassroots platforms is, naturally, that the protocol be (vi) grassroots.
 Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$.
@@
 (ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$.
-For this to happen in a permissioned consensus protocol, the set of agents $P$ need to be “reconfigured” to include agents in $P^{\prime}\setminus P$.
+Any platform that operates on a shared global resource or employs a global replicated (Blockchain [31]), or distributed (IPFS [7], DHT [35]) shared data structure, or distributed pub/sub systems with a global directory [13, 14, 8], are all not grassroots.
+Server-based federated systems such as Mastodon [34] are also not grassroots, informally due to the control exerted by each server-operator on its members; formally, since a group of people of any size cannot function without first connecting to such a server.
+For a permissioned consensus protocol to be grassroots, the set of agents $P$ needs to be “reconfigured” to include agents in $P^{\prime}\setminus P$.
 The question of reconfiguration has been studied extensively for decades [26, 17, 2, 47].
@@
 Paper structure.
-Section 2 presents the Grassroots Consensus protocol assuming a reliable network.
-The protocol is presented assuming a prevailing constitution that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony.
+Section 2 presents the Grassroots Consensus protocol assuming a reliable network, as well as a prevailing constitution $(P,\sigma,\Delta)$, that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony.
 Section 3 proves the safety and liveness of Grassroots Consensus.
@@
 Section 5 proves that the Grassroots Consensus protocol is indeed grassroots.
-Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present a distinct democratic constitutional amendment process for each parameter.
+Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present democratic processes geared for amending each of the three parameters.
 Section 7 concludes.
@@
 issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$.
-judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) a nack-block was received for $b^{\prime}$, or (iii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
+judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
 Algorithm 1 Grassroots consensus protocol, code for agent $p\in P$, given $(P,\sigma,\Delta)$.
@@
 Note that the two requirements cannot even be expressed in the context of the basic (non-constitutional) Grassroots Consensus protocol, as the population $P$ is given and fixed.
-And that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
+Also note that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
 6 Constitutional Grassroots Consensus
@@
 Regarding amending $P$ to $P^{\prime}$, requiring $\sigma$-supermajorities among both $P$ and $P^{\prime}$ to approve the amendment has been proposed before [17], to avoid, among other problems, the ‘dead sailors problem’ (in which all of $P^{\prime}$ are in fact dead, resulting in a constitutional stalemate) presented in the original Paxos paper [26].
+Once $P$ can be amended, the question of sybils (fake and duplicate identities) entering $P$, in addition to Byzantines, comes to the fore.
+First, since one Byzantine agent, having penetrated $P$ with multiple identities, can cause as much harm as multiple Byzantine agents.
+In fact, the principal goal of permissionless protocols, both Proof-of-Work [32] and Proof-of-Stake [25, 10], is to augment consensus with mechanisms that make it sybil-resilient.
+Second, since a person controlling multiple digital identities in $P$—even if all correct at the consensus level—will have undue influence on the meta-level democratic processes, violating one person – one vote.
+Existing large-scale digital communities suffer from extensive penetration of sybils.
+For example, Facebook removes around 1Bn fake accounts every quarter [48].
+Most work on sybil detection and elimination employs top-down approaches suitable for global platforms [50, 18, 37, 12, 4, 6], but by their nature do not provide support for the democratic formation and conduct of digital communities.
+For grassroots consensus, egalitarian sybil-repelling protocols for the grassroots formation of digital communities [33, 39] may be more appropriate.
+They are based on mutual trust among the people forming a community and their willingness to backup this trust with mutual sureties.
+These protocols ensure, under certain conditions, that a digital community can grow indefinitely while retaining a bounded fraction of sybils.
+A complementary line of work [38, 29], relevant to the meta-level democratic processes, shows how a community can employ sybil-resilient democratic governance: How the safety and liveness of digital democratic processes can be ensured despite a bounded penetration of sybils, even in the face of partial participation by its non-sybil members.
+Together, the two lines of work offer a comprehensive solution for (i) sybil-repellence during the egalitarian formation of digital communities and (ii) sybil-resilience in the democratic governance of digital communities, both needed during by the meta-level processes of constitutional amendment of Grassroots Consensus, as discussed next.
 (2) Amending $\sigma$.

### v4

@@
-Grassroots Consensus
+Constitutional Consensus
 Abstract
-Grassroots platforms aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike.
-Within the grassroots architecture, consensus is needed to realize platforms that employ digital social contracts, which are like smart contracts except that they are among people not accounts and are executed by these people’s smartphones not by high-performance servers controlled by parties outside to the contract.
-Key envisioned grassroots platforms include sovereign democratic digital communities and federations, community banks and their grassroots cryptocurrencies, and digital cooperatives.
-The grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during both low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots.
-The Grassroots Consensus protocol addresses all these requirements while having competitive performance in both low- and high-throughput scenarios and being one of the most concise and elegant consensus protocols for partial synchrony.
-It achieves that by building on two cutting-edge consensus protocols—the quiescent high-performance Morpheus and the blocklace-based Cordial Miners, improving the latter’s dissemination protocol and making it UDP-ready, and extending the protocol with a constitution and a constitutional amendment component, making it grassroots.
+Consider people with smartphones operating without external authorities or global resources other than the network itself.
+In this setting, high-end applications supporting sovereign democratic digital communities, community banks, and digital cooperatives require consensus executed by community members, which must be reconfigurable to support community dynamics.
+The Constitutional Consensus protocol aims to address this need by introducing constitutional self-governance to consensus: participants dynamically amend the participant set, supermajority threshold, and timeout parameter through the consensus protocol itself.
+We achieve this by enhancing a DAG-based protocol (like Cordial Miners) with participant-controlled reconfiguration, while also supporting both high- and low-throughput operation (like Morpheus), remaining quiescent when idle.
+This three-way synthesis uniquely combines: (1) constitutional amendments for self-governance, (2) a cryptographic DAG structure for simplicity, parallelism, and throughput, and (3) both high- and low-throughput operation.
+The protocol achieves consensus in $3\delta$, maintains $O(n)$ amortized communication complexity during high throughput, and seamlessly transitions between modes.
+The basic protocol (without constitutional amendments) realizes these features in 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
 keywords:
@@
 1 Introduction
-Grassroots.
-Grassroots platforms [44, 42, 40, 45, 46] aim to offer an egalitarian alternative to global platforms—centralized/autocratic and decentralized/plutocratic alike.
-A grassroots platform is different from a global platform in that it may have multiple instances that can operate independently of each other, without coordination or reliance on any third-party or global resource other than the network itself; furthermore, different instances of the same grassroots platform may interoperate and coalesce into ever-larger instances, possibly, but not necessarily, resulting in a single global instance.
-Decentralized platforms for global cryptocurrencies [31, 52] and smart contracts [15, 9, 30] supporting Decentralized Finance (DeFi) [36, 21] and Decentralized Autonomous Organizations (DAOs) [19, 16, 49] often employ a consensus protocol as their foundation.
-Hence the surge in interest in high-performance and high-throughput consensus protocols [53, 23, 24, 27, 5, 22].
-Grassroots platforms, on the other hand, are not consensus-based.
-Their foundational platform consists of a grassroots social graph implemented via a blocklace—a DAG-like generalization of the linear blockchain that is a universal Conflict-free Replicated Data Type (CRDT) [3].
-The grassroots architecture [42, 44] facilitates the implementation of a grassroots social network [41] and grassroots cryptocurrencies [43, 27], which are also consensus-free, on top of the grassroots social graph.
-So where does consensus come into play?
-Consensus.
-Within the grassroots architecture, consensus is needed to realize higher-level platforms that employ digital social contracts—the grassroots counterpart of smart contracts.
-Briefly, while a smart contract [15] is among anonymous accounts and runs on a consensus protocol executed by third-party servers, a social contract is among people known to each other and runs on a consensus protocol executed by the smartphones of these very same people [44, 46].
-The two types of consensus-based contracts are compared in Table 1.
-Table 1: Comparing Consensus-Based Smart Contracts and Social Contracts
+Participant-Controlled Consensus.
+Consider people who wish to form self-governing digital communities without relying on servers or external authorities.
+These participants—whether members of community banks, digital cooperatives, or democratic assemblies—need consensus protocols they can execute on their smartphones while maintaining complete control over membership and governance rules.
+Unlike traditional reconfiguration where external authorities determine changes [40, 2], these communities require constitutional self-governance: participants must collectively control who joins, how decisions are made, and how the protocol operates, all through the consensus protocol itself.
+This paper presents Constitutional Consensus, a protocol that enables such self-governance by allowing participants to dynamically amend the participant set $P$, supermajority threshold $\sigma$, and timeout parameter $\Delta$ through democratic decisions implemented within the protocol.
+Digital Social Contracts.
+These self-governing communities implement digital social contracts [9]—deterministic state machines that encode agreements among people who know each other.
+Like smart contracts, they provide computational guarantees about contract execution.
+However, while smart contracts [10] operate among anonymous accounts and run on consensus protocols executed by third-party servers, social contracts are among identified people and run on consensus protocols executed by the participants’ own devices.
+Existing consensus-based smart contracts, notably in Decentralized Finance (DeFi) platforms [10, 7, 25], have driven development of high-performance consensus protocols [43, 17, 18, 4, 16], but these operate among pseudonymous accounts on global networks rather than identified people in local communities.
+This fundamental difference enables democratic rather than plutocratic governance, as shown in Table 1.
+Table 1: Comparing Smart Contracts and Consensus-Based Digital Social Contracts
 Smart Contract [10] Digital Social Contract [9]
@@
 Executed by: Third-party servers The people’s smartphones
-In return for: Gas Love
+In return for: Gas Naught
 Applications: Decentralized Finance (DeFi) [31, 15],
@@
 Amended via: Self-modifying code Constitutional amendments
-Example: Grassroots Federation.
-A key application of digital social contracts mentioned in Table 1 is the grassroots federation of digital communities.
-Grassroots Federation [46] aims to address the egalitarian formation and the fair democratic governance of large-scale, decentralized, sovereign digital communities, the size of the EU, the US, existing social networks, and even humanity at large.
-A grassroots federation evolves via the grassroots formation of digital communities and their consensual federation.
-Such digital communities may form according to geography, jurisdiction, affiliations, relations, interests, causes, and more.
-Small communities (say up to $100$ members) govern themselves; larger communities—no matter how large—are governed by a similarly-small assembly elected by sortition [20, 46] among its members.
-It is expected that the constitutional democratic governance of each community would be specified by a digital social contract running on top of a grassroots consensus protocol, which in turn would be executed by the members of said small community or its assembly, one protocol instance per community.
-Such a consensus protocol should satisfy all the requirements explained next.
-Requirements of Grassroots Consensus.
-Small communities engaged in democratic conduct, small community banks, small cooperatives, etc., would have a transaction rate much lower than network latency.
-Hence they would need a (i) quiescent protocol in which participants do nothing (in particular, send no messages) until there is a new transaction, and when subsequent transactions are issued at low-throughput, finalize each quickly and efficiently.
-As during low-throughput transactions are issued one-at-a-time, there is no symmetry to break or conflicts to resolve and hence—ideally—the agent issuing a transaction should be able to finalize it without the help of a leader, namely during low-throughput the protocol should be leaderless.
-At the same time, the protocol should also work efficiently and with low-latency on smartphones in a high-throughput scenario.
-Namely, it should be an (ii) efficient consensus protocol during both low- and high-throughput.
-In addition, a standard requirement, termed (iii) responsiveness, is that in the good case the protocol operates at network speed, with performance unaffected by an overly-conservative estimation $\Delta$ of the actual least upper bound on network delay $\delta$.
-A fourth requirement relates to the blocklace: As the lower-level grassroots platforms (Social Graph and Social Networks [41], Cryptocurrencies [43, 27]) are blocklace-based, it would be very beneficial and integrable to also have a (iv) blocklace-based consensus protocol.
-This is not a limitation as there are plenty of efficient and high-throughout blocklace-based protocols already, starting from Cordial Miners [24] and its extensions and improvements [5, 22].
-We note that a blocklace [3] is not “just a DAG” the same way a blockchain is not “just a sequence”.
-In a blocklace, directed edges are realized by signed cryptographic hash pointers: A $p$-block (block by agent $p$) $b$ includes a cryptographic hash of the rest of $b$, signed by $p$, which serves as its unique (whp) identifier, also referred to as a pointer to $b$.
-An edge from the $p$-block $b$ to the $q$-block $b^{\prime}$ is realized by $b$ including a pointer to $b^{\prime}$.
-Thus, similarly to a blockchain, the blocklace is:
+Governance: Token-weighted voting (plutocracy) One person, one vote (democracy)
+Amendment Authority: External governance tokens or DAOs Participants themselves via supermajority
+Parameter Control: Fixed in code or controlled by token holders Democratically adjustable by participants
+We note that while smart contracts are necessarily consensus-based, digital social contracts are not.
+Some grassroots platforms, notably grassroots social networks [35] and grassroots cryptocurrencies [37, 22], do not require consensus on the total ordering of the acts of their participants: Their safety depends only on the agents obeying the social contract [38].
+This can be ensured by mechanisms weaker then consensus, e.g.
+mutual attestations among agents operating within Trusted Execution Environments (TEEs) [30].
+However, applications requiring collective agreement—such as community banks with multiple signatories or democratic governance with voting—do require consensus to ensure all participants agree on the order of transactions and decisions.
+Technical Requirements.
+Communities implementing digital social contracts face unique technical challenges.
+Small communities may have transaction rates much lower than network latency, requiring a protocol that remains quiescent—sending no messages when idle.
+Yet the same protocol must efficiently handle high-throughput periods when the community is larger and/or more active.
+During low throughput, transactions arrive one-at-a-time with no conflicts to resolve, so ideally any participant should be able to finalize their transaction without designated leaders.
+During high throughput, the protocol should match the performance of modern consensus protocols.
+Specifically, an effective protocol for self-governing communities requires:
   1. 1.
-    Tamper-proof: A bit in a block cannot be changed without this being detected;
+    Quiescent operation: No activity when idle, spontaneous leader emergence for isolated transactions
   2. 2.
-    Non-repudiable: The creator of a block cannot deny having created it;
+    Dual-mode efficiency: Seamless transition between low and high throughput modes
   3. 3.
-    Acyclic: Even Byzantine agents cannot create blocks that form/close a cycle.
-However, unlike the blockchain, which is extended competitively, with the point-of-contention being adding the next block to the most-recent block, a blocklace is a naturally-cooperative data type, as agents may add blocks to the blocklace at will without conflicting with each other.
-In fact, the blocklace is a Universal Conflict-free Replicated Data Type (CRDT) [3].
-The grassroots architecture in general, and the consensus protocol presented here in particular, are geared for smartphone-based execution.
-Namely, the intention is that the parties to a social contract will run both the contract and its underlying consensus protocol on their smartphones.
-The vagaries of the Internet—NATs and firewalls—and the fact that mobile phones change their IP address as they roam about, make phone-to-phone communication difficult.
-In particular, opening a direct TCP connection between two phones is either impossible, or impossible without the help of a proxy server, and even if established, it has to be re-established whenever one of the phones changes their IP address.
-Hence, UDP is a much preferable protocol for phone-to-phone communication, and is the one typically used for audio and video phone-to-phone conversations.
-Thus, the fifth requirement is that the protocol work not only on reliable networks, as typically assumed by consensus protocols, but also on unreliable networks namely be (v) UDP-ready.
-The final sixth requirement for a protocol for grassroots platforms is, naturally, that the protocol be (vi) grassroots.
-Informally, this means two things: (i) That a group of agents $P$ running the protocol can do so unscathed if embedded within a larger group of agents $P^{\prime}\supset P$.
-Any permissioned consensus protocol satisfies that.
-(ii) That the protocol can do more if $P$ is embedded within a larger group of agents $P^{\prime}$.
-Any platform that operates on a shared global resource or employs a global replicated (Blockchain [31]), or distributed (IPFS [7], DHT [35]) shared data structure, or distributed pub/sub systems with a global directory [13, 14, 8], are all not grassroots.
-Server-based federated systems such as Mastodon [34] are also not grassroots, informally due to the control exerted by each server-operator on its members; formally, since a group of people of any size cannot function without first connecting to such a server.
-For a permissioned consensus protocol to be grassroots, the set of agents $P$ needs to be “reconfigured” to include agents in $P^{\prime}\setminus P$.
-The question of reconfiguration has been studied extensively for decades [26, 17, 2, 47].
-While the original informal exposition of the problem in the Paxos paper was with respect to a self-governed parliament [26], subsequent works in the context of permissioned consensus assume that the initial ‘permissioned’ agents are determined by an external authority, and so is their ‘reconfiguration’.
-Proof-of-Stake protocols such as Ethereum 2.0 offer a stake-based approach to reconfigure the participating agents on every epoch [25, 10].
-A grassroots system cannot have external authorities and hence protocol reconfiguration must be entirely under the control of the very same agents that execute the protocol, as envisioned in the Paxos paper [26].
-Furthermore, to be egalitarian rather than plutocratic, agents should have equal power in determining reconfiguration decisions.
-Specifically, in the grassroots context the challenge relates to the consensus protocol being executed by a self-governed community (the parties to the social contract running on top of the consensus protocol), allowing them to add or remove members at will, as well decide to change other aspects of its operation, for example the supermajority required to finalize a decision.
-To address that, we extend the protocol with a constitution that specifies the agents that execute the protocol as well as two key protocol parameters—the supermajority $\sigma$ by which decisions, including constitutional amendment decisions, are binding, and a time $\Delta$ estimating the unknown least upper bound on network delay during synchrony, $\delta$.
-To summarize, the grassroots architecture can benefit from a consensus protocol that is (i) quiescent, (ii) efficient during low- and high-throughput, (iii) responsive, (iv) blocklace-based, (v) UDP-ready, and (vi) grassroots.
-The Grassroots Consensus protocol was conceived to address all these requirements.
+    Responsiveness: Network-speed operation unaffected by conservative timeout estimates
+  4. 4.
+    DAG-based structure: Leveraging blocklace [3] parallelism and integration
+  5. 5.
+    Constitutional amendments: Participant-controlled governance of protocol parameters
+Technical Foundation.
+Constitutional Consensus achieves these requirements through a three-way synthesis.
+From Morpheus [23], it inherits dual-mode operation with spontaneous leaders during low throughput and formal leadership during high throughput.
+From Cordial Miners [18], it adopts the blocklace structure for parallelism and efficiency.
+To these, it adds constitutional amendment mechanisms enabling true self-governance.
+The protocol operates in waves of three rounds.
+During low throughput, any participant with a transaction can spontaneously lead a wave, achieving finality in $3\delta$.
+During high throughput, formal leaders are selected round-robin, maintaining $O(n)$ amortized communication complexity.
+The basic protocol (without amendments) requires only 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
+Constitutional Self-Governance.
+The key innovation lies in how participants control their own governance.
+Unlike smart contracts that rely on external governance tokens or DAOs [12], constitutional amendments are decided by the participants themselves through supermajority vote.
+While Lamport’s original Paxos envisioned a self-governing parliament [21], subsequent reconfiguration protocols [42, 2, 40, 20] assume external authorities determine membership changes.
+Constitutional Consensus returns to Paxos’s original vision of true self-governance.
+Unlike Proof-of-Stake systems where validators must stake capital [8], participation requires only the democratic approval of existing members.
+This creates a fundamental architectural difference: while blockchain systems typically support only a single global instance whose security depends on token value, Constitutional Consensus enables unlimited independent communities to self-govern simultaneously.
+Connection to Grassroots Architecture.
+This capability for multiple independent instances operating without central coordination or global resources makes Constitutional Consensus inherently grassroots [38].
+A community in Bangladesh, a cooperative in Brazil, and a federation in Berlin can each run their own instance, potentially with overlapping members, without any coordination or competition for resources.
+This architecture—multiple local instances rather than a single global instance—enables the broader grassroots vision of egalitarian digital platforms that operate without central authorities [36, 34].
+Within this grassroots architecture, Constitutional Consensus provides the consensus layer for applications requiring collective agreement, complementing consensus-free grassroots platforms like social networks [35] and cryptocurrencies [37, 22].
+Together, these protocols enable communities to build sovereign democratic digital infrastructure from the ground up, controlled entirely by their members rather than external authorities or token holders.
+This distinguishes Constitutional Consensus from global platforms—whether centralized like Facebook or decentralized like Bitcoin [26] and IPFS [5]—that can support only a single instance.
+Example: Grassroots Communities, Banks, and Federations.
+As a concrete scenario for grassroots platforms, consider a small town whose residents wish to form a digital community.
+They begin by deploying independent instances of consensus-free grassroots platforms on their smartphones—creating their own grassroots social network [35] and trading via self-minted grassroots cryptocurrencies [37, 22].
+Crucially, they do this without coordination with any global network or central authority.
+When their needs evolve to require collective agreement—community banks with multiple signatories, constitutional governance, or federation with neighbouring towns—they deploy their own Constitutional Consensus instance.
+This instance operates independently yet can interoperate with other communities’ instances, embodying the grassroots principle of multiple local instances rather than a single global system.
+Through Grassroots Federation [39], these autonomous instances can scale democratically from local to potentially global governance.
+Small communities ($\approx 100$ members) govern themselves directly; larger ones use sortition-selected assemblies [14, 39].
+Each community’s governance runs as a digital social contract on their own Constitutional Consensus instance, executed by members on their smartphones—no external servers or authorities required.
+Model.
+As deterministic fault-tolerant consensus is not solvable in asynchronous systems [13], we assume an eventual synchrony model, where communication delays are at most $\delta$ after some Global Stabilization Time (GST) that is unknown to the agents.
+We consider Byzantine failures, as some malicious actors may infiltrate the network, but we assume a bounded ratio of Byzantine agents (to be specified later), since, by and large, the agents trust each other.
 Paper structure.
-Section 2 presents the Grassroots Consensus protocol assuming a reliable network, as well as a prevailing constitution $(P,\sigma,\Delta)$, that specifies the set of agents $P$, the supermajority $\sigma$ among the agents by which the protocol operates, and a presumed upper-bound $\Delta$ on network delay during synchrony.
-Section 3 proves the safety and liveness of Grassroots Consensus.
-Section 4 extends the protocol for eventually-reliable networks, making it UDP-ready.
-Section 5 proves that the Grassroots Consensus protocol is indeed grassroots.
-Section 6 extends Grassroots Consensus with constitutional amendment, with which the agents may amend these three protocol parameters, and present democratic processes geared for amending each of the three parameters.
+Section 2 provides the model and problem definitions.
+Section 3 presents a single-epoch Constitutional Consensus protocol, with only an initial constitution and no constitutional amendments.
+Section 4 proves the safety and liveness of single-epoch Constitutional Consensus.
+Section 5 adds support for constitutional amendments – for each of the three parameters $P,\sigma$, and $\Delta$, we present a democratic process for amending it.
+We then present the protocol extension needed to realize constitutional amendments and prove its safety and liveness.
+Section 6 proves that Constitutional Consensus is grassroots.
 Section 7 concludes.
-2 Grassroots Consensus
+2 Model and Problem
+Model.
+We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$.
+While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it.
+A correct agent follows the protocol.
+An agent that does not follow the protocol is faulty or Byzantine.
+We assume an eventual synchrony (aka partial synchrony) model in which every message sent between correct agents is eventually received.
+Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
+Note that the abstraction of reliable messaging does not mean that agents must communicate via point-to-point TCP connections.
+Rather, it abstracts away a dissemination layer that can be implemented using acks/nacks, gossip, network coding, etc., and is orthogonal to the consensus protocol in this paper.
+A protocol proceeds through a sequence of epochs, each commences with a new constitution (defined next), referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution.
+The constitution defines the set of participants in the epoch, as well as key protocol parameters:
+Definition 2.1 (Constitution, $\sigma$-Supermajority).
+A constitution is a triple $(P,\sigma,\Delta)$ where:
+  1. 1.
+    $P\subset\Pi$ is a set of agents, with $n:=|P|$,
+  2. 2.
+    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the agents $P$,
+  3. 3.
+    $\Delta>0$ is a presumed upper-bound on message delay after GST among $P$.
+We may omit $\Delta$ when its value is immaterial (e.g., when discussing protocol safety) and refer to the constitution $(P,\sigma)$.
+A $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
+Supermajorities and faulty agents.
+Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$, and with this assumption we define $\sigma:=\frac{n+f}{2n}$.
+Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty agents in $P$.
+For example, for $f=0$, $\sigma=\frac{1}{2}$; for $f=\frac{1}{2}n$, $\sigma=\frac{3}{4}$; and if everyone but one agent is faulty then $\sigma=\frac{2n-1}{2n}=1-\frac{1}{2n}$, namely unanimity is required.
+For $n=3f+1$ (the standard assumption), $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
+Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”.
+Constitution safety and liveness.
+A constitution thus defines the protocol’s assumptions about the environment: $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$ and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent; $\Delta\geq\delta$ is an upper bound on the communication time after GST.
+Definition 2.2 (Safe and Live Constitution).
+A constitution $(P,\sigma,\Delta)$ is safe if $P$ includes at most $f=(2\sigma-1)n$ faulty agents, and live if after GST any message sent among correct agents in $P$ arrives within $\Delta$.
+Constitutional Consensus problem definition.
+A constitutional consensus protocol may have multiple instances.
+Each instance commences with an initial constitution $(P,\sigma,\Delta)$ that defines its participating agents $P$ and the key protocol parameters $\sigma$ and $\Delta$.
+As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
+In addition, constitutional consensus assumes an that participants engage in an unspecified democratic decision process that may result in a decision $d$ to amend the prevailing constitution.
+The discussion of such democratic processes is deferred to Section 5.
+For the purpose of the constitutional consensus protocol, all is needed is to assume: (i) the existence of such a process that may result in constitutional amendment decisions, which become known to the participants once made; (ii) that participants can verify a constitutional amendment decisions $d$ to be valid according to the rules specified in Section 5.
+A valid constitutional amendment decision $d$ defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$, in which the participants $P^{\prime}$ and/or the parameters $\sigma^{\prime}$ and $\Delta^{\prime}$ are the result of amending the prevailing constitution.
+In the following, constitution refers to the new constitution of a valid constitutional amendment decision, and an initial constitution is the new constitution of a valid amendment of a constitution with an empty population.
+Valid amendments are provided to the consensus protocol as inputs by participants that know of them.
+Thus, the output sequence of a constitutional consensus protocol includes both transactions and valid constitutional amendment decisions, as follows:
+Definition 2.3 (Constitutional sequence, epoch, epoch sequence, prevailing constitution, participants).
+A constitutional sequence is a sequence $T=t_{1},t_{2},\dots$ of transactions and valid constitutional amendment decisions beginning with an initial constitution.
+Each constitutional amendment decision $d$ in the sequence defines a new epoch, its epoch sequence is a maximal contiguous subsequence of $T$ that begins with $d$ and does not include another such decision, and its prevailing constitution is the new constitution of $d$.
+An agent participates in an epoch if it is a member of the population of its prevailing constitution.
+The constitutional output sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
+Definition 2.4 (Constitutional Consensus Correctness).
+In any run of an instance of constitutional consensus in which all valid constitutional amendment decisions made are safe and live, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following safety and liveness properties hold:
+  1. 1.
+    Safety:
+    Consistency
+    Every correct agent’s output sequence is a subsequence of $T$.
+    Validity
+    If a correct agent $p$ outputs a transaction $t$, then $p$ is a participant in $t$’s epoch in $T$.
+  2. 2.
+    Liveness:
+    Epoch Liveness
+    If a correct agent $p$ is a participant in an epoch $d$ in $T$, then $p$’s output sequence eventually includes all the transactions in $d$’s epoch sequence.
+Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in the epoch $d$ is included in $T$.
+    Constitution Amendment Liveness
+    If a valid constitutional amendment decision $d$ is provided as input to the protocol by a correct agent, then $T$ eventually includes $d$.
+Note that if there is no decision to amend the initial constitution then the protocol run has a single epoch with a fixed set of participants.
+In this case, the definition above coincides with the standard consensus correctness definition.
+Note also that progress is guaranteed (after GST) even if the constitution changes infinitely many times.
+However, normal transactions might starve due to frequent amendments.
+This definition allows the implementation to prioritize constitution changes over regular transactions, allowing the latter to starve.
+In practice, amendments are expected to be rare compared to transactions, so we believe this modus operandi to be reasonable.
+3 Single Epoch Constitutional Consensus
+Here we present the behaviour of the Constitutional Consensus protocol during a single epoch.
+The protocol is based on the blocklace data structure, which we define in Section 3.1.
+The protocol’s concepts are detailed in Section 3.2 and its pseudo-code in Section 3.3.
 3.1 Blocklace Preliminaries
 We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$.
-We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$.
-While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it.
-A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$, with $b$ called initial if $H=\emptyset$.
-An encoding of such a triple $b=(h,x,H)$ via a sequence of bits in an agreed-upon form is referred to as a well-formed block.
-We assume the hash function is collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$.
+A $p$-block $b$, with $p\in\Pi$, is a triple $b=(h,x,H)$ where $h$ is the hash of the pair $(x,H)$ signed by $p$, referred to as the identifier of $b$, $x$ is an arbitrary payload, which may be empty, $x=\bot$, in which case we refer to $b$ as an empty block, and $H$ a finite set of block identifiers, the blocks of which are pointed to by $b$.
+A block with no pointers, $H=\emptyset$, is called initial, in which case it can also be written as a pair $b=(h,x)$.
+An encoding of a block via a sequence of bits in an agreed-upon form is referred to as a well-formed block.
+We assume the hash function to be collision-free whp and cryptographic, so that blocks cannot form cycles and we may identify a block $b$ with its identifier $h$.
 Furthermore, the method of signature allows the public key $p$ to be recovered from a signature by $p$.
+A set of blocks $X$ is referred to as a $\sigma$-supermajority given a set of agents $P$ if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $X$ includes at least one $q$-block.
 A block $b$ observes itself, any block $b^{\prime}$ pointed to by $b$, and any block observed by $b^{\prime}$.
@@
 The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$.
-If a block variable $b$ is undefined then we deem its closure $[b]:=\emptyset$ to be empty rather than undefined.
-For a set of blocks $B$ the closure $[B]$ is defined by $[B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$.
-A set of blocks $B$ is closed if $[b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$.
+For a set of blocks $B$, its closure $[B]$ is defined by $B]:=\bigcup_{b\in B}[b]$, which is also the set of blocks observed by $B$.
+A set of blocks $B$ is closed if $b]\subseteq B$ for every $b\in B$, equivalently if $B=[B]$.
 A blocklace is a closed set of blocks.
-To help avoid confusion, we use $B$ below to denote a closed set of blocks (a blocklace) and $D$ to denote any set of blocks, not necessarily closed.
+To help avoid confusion, we use $B$ to denote a closed set of blocks (a blocklace) and $D$ or $X$ to denote any set of blocks, not necessarily closed.
+Also, we refer to a set of blocks by agents in $P\subset\Pi$ as $P$-blocks, the let the $P$-blocks in $B$ denote the set of all blocks in $B$ by agents in $P$.
 Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$.
@@
 A block $b\in D$ is a tip of $D$ if no other block $b^{\prime}\neq b\in D$ observes $b$.
-The depth $d(b)$ of a block $b$ is $1$ if $b$ is initial else $d(b)=d(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$.
-The depth $d(B)$ of a set of blocks $B$ is defined by $d(B):=\max_{b\in B}d(b)$.
-A round $r\geq 1$ in $B$ is the set of all blocks $b\in B$ of depth $d(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with depth $d(b)\leq r$, equivalently all rounds in $B$ up to and including $r$.
-Note that, by definition, all blocks of the same round do not observe each other.
+The depth of a block $b$, $\textit{depth}(b)$, is 0 if $b$ is initial else $\textit{depth}(b)=\textit{depth}(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$.
+The depth of a set of blocks $B$ is defined by $\textit{depth}(B):=\max_{b\in B}\textit{depth}(b)$.
+In the context of blocks, we use most-recent as synonym for deepest.
+A round $r\geq 0$ in $B$ is the set of all blocks $b\in B$ with $\textit{depth}(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with $\textit{depth}(b)\leq r$, equivalently all rounds in $B$ up to and including $r$.
+Note that, by definition, blocks of the same round do not observe each other.
 3.2 Protocol Concepts
-Models.
-We consider two models.
-The first is the standard model of eventual synchrony (aka partial synchrony) that assumes a reliable network in which every message sent among correct agents is eventually received.
-Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
-In particular, all messages sent any time before GST arrive by time GST+$\delta$.
-In the second model the assumption of a reliable network is relaxed.
-It assumes an eventually-reliable network, in which any message sent infinitely often between correct agents eventually arrives, and a message sent after GST between correct agents arrives within $\delta$.
-This is the model assumed in the original Paxos paper [26] and many sequels.
-We refer to this model as the UDP-ready model.
-Parameters. The protocol employs three key parameters $(P,\sigma,\Delta)$:
+The blocklace created during an epoch begins with a constitutional genesis block $(h,d)$ where $d$ specifies a new constitution $C=(P,\sigma,\Delta)$.
+In fact $d$ also records the constitutional decision to amend the old constitution (if any) to the new constitution and is signed by multiple agents; these are discussed in Section 5.
+Here, we focus on the behaviour of the protocol during a single epoch.
+For now we refer to a blocklace in which all initial blocks are constitutional genesis blocks as a constitutional blocklace.
+Every participant $p\in P$ commences an epoch $b$ with its blocklace including $b$ and its local epoch identifier $e$ set to $b$, $e\leftarrow b$.
+The protocol employs an evolving shared blocklace.
+Each participant $p\in P$ maintains a local copy of the constitutional blocklace.
+In the course of an epoch, agents create new blocks that observe the epoch’s genesis block and and disseminate them to other agents.
+Blocks may include transactions the agents wish to order, if they have any.
+While a blocklace is a partially ordered set of blocks, the protocol totally orders the blocks in an epoch in a consistent way, producing the same output sequence at all participants in an epoch.
+Moreover, epochs are also totally ordered, so that the totality of all blocks of all epochs are also totally ordered, even though not always by the same participants
+Because correct agents follow the protocol, they do not equivocate.
+Hence, for a blocklace $B$ created by the protocol with constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
+Waves.
+Given a constitutional blocklace $B$, a wave $W\subseteq B$ in a blocklace is either its constitutional genesis block, referred to as its $0^{th}$-wave, or the set of blocks of depths $r+1$, $r+2$, and $r+3$ in $B$, referred to as the the first, second, and third round of the wave, respectively, for any $r\geq 0$ such that $r\mod 3=0$.
+Such a wave is the $k^{th}$ wave of $B$, $k:=(r+3)/3$.
+For a given $P\subset\Pi$, the function leaderP maps integers to $P$.
+For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$.
+We assume the leader function is defined so that it is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
+Finalizing blocks in waves.
+The consensus protocol commences with a constitutional blocklace and progresses in waves.
+The protocol strives to finalize a new block in each wave.
+The first round of a wave includes candidates for finality, the second round includes endorsements of candidates, and the third round includes ratifications.
+A first round block is a candidate for finality in two cases—if it is a formal leader block or the only block in its round.
+These two possibilities typically arise in high- and low-throughput scenarios, respectively.
+During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with any others, all agents produce second and third round blocks that finalize $b$, in which case $b$ is referred to as spontaneous leader.
+During high throughput, multiple agents may have transactions to send, and the leader function is used to break symmetry, so that only the formal leader sends a leader block in the first round.
+Finalized leader blocks of both types induce the total ordering of all blocks, as in other DAG-based protocols [17, 18, 23].
+A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows.
+Note that the next two definitions are mutually-recursive.
+Definition 3.1 (Endorse).
+Let $b^{\prime}$ be a second round block that approves a first round block $b$ in a wave $W$ in the blocklace $B=[b^{\prime}]$. Let $(P,\sigma)$ be the prevailing constitution.
+Then $b^{\prime}$ endorses $b$ if the most-recent wave $W^{\prime}\subset B$ that precedes $b$ is:
   1. 1.
-    $P\subset\Pi$, a set of agents,
+    quiescent and $b$ is the only first-round block approved by $b^{\prime}$, or
   2. 2.
-    $\frac{1}{2}\leq\sigma<1$, a fraction, used to specify a supermajority among the agents
-  3. 3.
-    $\Delta>0$, a presumed upper-bound on message delay after GST, known to the agents.
-We refer to the parameters $(P,\sigma,\Delta)$ employed by the protocol for a given block $b$ or blocklace $B$, with values as above, as the prevailing constitution in $b$, resp.
-$B$; when the value of $\Delta$ is immaterial we may refer to the prevailing constitution $(P,\sigma)$.
-We begin by assuming the prevailing constitution to consist of arbitrary constant values for all blocks during the protocol run, as is usually the case.
-Later, in Section 6, we augment the protocol with constitutional amendment blocks by which the prevailing constitution may be amended.
-Supermajorities.
-Given agents $P\subset\Pi$, $|P|=n$, and $\frac{1}{2}\leq\sigma<1$, a $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$, or, equivalently, that $\frac{|Q|}{|P|}>\sigma$.
-A set of blocks $D$ is referred to as a $\sigma$-supermajority if there is a $\sigma$-supermajority $Q\subseteq P$ such that for every $q\in Q$, $D$ includes at least one $q$-block.
-Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$.
-Given $f$, we say that a blocklace $B$ is $f$-safe if it includes equivocations by at most $f$ agents, a block $b$ is $f$-safe if $[b]$ is $f$-safe, and define $\sigma:=\frac{n+f}{2n}$.
-For example, if $f=0$ then $\sigma=\frac{1}{2}$; if $f=\frac{1}{2}n$ then $\sigma=\frac{3}{4}$; if $n=3f+1$ (the standard assumption), then $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
-Under this assumption, a $\sigma$-supermajority includes a majority of the correct agents, and hence the intersection of any two $\sigma$-supermajorities includes at least one correct agent.
-Typically, permissioned consensus protocols refer to “blocks by at least $n-f$ agents among $P$”, which is equivalent to “a $\sigma$-supermajority among $P$” in the standard case of $n=3f+1$.
-Waves.
-A blocklace is seen as a sequence of waves, where the $k^{th}$-wave in $B$ , $k\geq 1$, consists of three consecutive rounds $r$, $r+1$, $r+2$ in $B$, $r=3(k-1)+1$, referred to as the first, second, and third round of the wave.
-For a given $P\subset\Pi$, the partial function leaderP assigns a leader $p\in P$ for the first round of every wave, where a first-round $p$-block $b$ with prevailing constitution $(P,\sigma)$ is a leader block if $p=\textit{leader}_{P}(d(b))$.
-We assume the leader function to be fair in that for every wave and any $p\in P$ there is a subsequent wave for which $p$ is the leader of its first round.
-The consensus protocol progresses in waves.
-In each wave, a first round block is a candidate for finality in two cases—if it is a leader block or the only block in its round.
-These two possibilities typically arise in high- and low-throughput scenarios, respectively.
-During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with others, all agents produce second and third round blocks that finalize $b$.
-During high throughput, conflicting blocks may be produced and the leader function is used to break symmetry; leader blocks are finalized and induce the total ordering of all blocks.
-A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows.
-Note that the next two definitions are mutually-recursive:
-Definition 2.1 (Endorse).
-Let $b$ be a first-round block with prevailing constitution $(P,\sigma)$ and $b^{\prime}$ a second-round block of the same wave in $B$. Then $b^{\prime}$ endorses $b$ if the previous wave is:
+    non-quiescent and $b$ is a $\textit{leader}_{P}$ block approved by $b^{\prime}$.
+A set of second-round blocks $X\subset W$ endorses $b$ if every $b^{\prime}\in X$ endorses $b$.
+Note that endorsement is an invariant block property, as it depends only on $b^{\prime}]$.
+Next, we specify the finality rule.
+A first-round block $b\in W$ is final if the third round of $W$ has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$.
+Formally:
+Definition 3.2 (Ratified, Final, Ordered, Quiescent).
+Assume a blocklace $B$ and a wave $W\subseteq B$. A first-round block $b\in W$ with prevailing constitution $(P,\sigma)$ is:
   1. 1.
-    quiescent and $b$ is the only block in its round approved by $b^{\prime}$, or
+    ratified by a third-round block $b^{\prime}\in W$ if $b^{\prime}$ approves a $\sigma$-supermajority in $P$ that endorses $b$, in which case $b$ is also ratified in $W$ and in $B$; and
   2. 2.
-    non-quiescent and $b$ is a leader block approved by $b^{\prime}$.
-A set of blocks $D$ endorses $b$ if every $b^{\prime}\in D$ endorses $b$.
-A first-round block $b$ is final if the third round of the same wave has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$.
-Formally (text in parenthesis is explanatory, the definition is valid also without it):
-Definition 2.2 (Ratified, Final, Ordered, Quiescent).
-Given a blocklace $B$, a block $b\in B$ with prevailing constitution $(P,\sigma)$ is:
+    final in $W$ and in $B$ if the third-round of $W$ includes a $\sigma$-supermajority in $P$, referred to as a finalising $\sigma$-supermajority, in which each block ratifies $b$.
+The wave $W$ is:
   1. 1.
-    ratified by a block $b^{\prime}$ if $b^{\prime}$ approves a $\sigma$-supermajority that endorses $b$;
+    finalising if it includes a block $b$ that is final in $W$; and
   2. 2.
-    final in $B$ if there is a $\sigma$-supermajority in the third-round of $b$’s wave in $B$, in which each block ratifies $b$, referred to as a finalizing $\sigma$-supermajority.
-A wave in such a blocklace $B$ is:
-  1. 1.
-    finalizing if it includes a (first-round) final block $b$; and
-  2. 2.
-    quiescent if in addition all the wave’s blocks except perhaps $b$ are empty and $b$ does not conflict with any block in $B$.
-As an initial condition, we refer to the (non-existent) wave that ends in round 00 as quiescent. A block $b$ is slow if it is both finalized by a quiescent wave and follows a quiescent wave.
+    quiescent if, in addition, all blocks in $W$ except perhaps $b$ are empty, and no block in $B$ conflicts with $b$.
+In addition, the $0^{th}$-wave is defined to be quiescent and its constitutional genesis block to be final.
 Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final.
+Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered.
 We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$.
-Note, however, that quiescence is not monotonic, both because conflicting blocks may be added to a round, and because more rounds can be added to the blocklace.
+Note, however, that quiescence is not monotonic, because conflicting and non-empty blocks may be added to a round.
 The next observation asserts that a wave’s final block is unique. {observation}[Uniqueness] A wave can have at most one ratified or final block.
+Proof 3.3.
+The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
 Quiescence, low- and high-throughput scenarios.
 The protocol is initially quiescent.
-It begins operations and continues to do so as long as agents have payloads to send.
+It begins operations and continues to operate as long as agents have payloads to send.
 If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode.
 It returns to being quiescent following a quiescent wave.
-During low-throughput, in the good-case scenario, all waves are quiescent.
+Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block, which is thus finalized by the wave.
+Low-throughput good case: The previous wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after the previous round $r$ has advanced, followed promptly by two rounds of empty blocks, resulting in a finalising and quiescent wave.
+High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced and all (or many) of the wave’s blocks are non-empty.
+Figure 1: A good-case wave $w$.
+We first consider the good-case scenarios summarized in Figure 1.
+During low-throughput all waves are quiescent.
 A wave commences when some agent has a new transaction.
-This agent issues a (first-round) non-empty block, and all agents follow by issuing two rounds of blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final.
-The new wave is also quiescent provided all second- and third-round blocks are empty.
-The agent that issued a finalized, non-empty, first-round block in this scenario is referred to as a spontaneous leader, in contrast with the formal leader assigned by the leaderP function.
-During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above, except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty (contain transactions to be ordered by the next-wave’s final leader block).
-The good-case scenarios are summarized in Figure 1.
+This agent issues a (first-round) non-empty block, and all agents follow promptly with two rounds of empty blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final, with the new wave also being quiescent.
+During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty, with transactions to be ordered by the the next final leader block.
 Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own.
-Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block.
-And if during low-throughput there are multiple agents issuing first-round blocks, then it is possible that none of the first round blocks will become final.
+Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block, with blocks by correct agents to be ordered by the leader of some subsequent finalising wave.
+Similarly, if during low-throughput there are multiple agents issuing first-round blocks, then typically none of the first-round blocks will become final.
 In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
@@
 Definition 3.4 (Advanced Round).
-Given a blocklace $B$ with prevailing constitution $(P,\sigma)$, a round $r>0$ in $B$ is advanced if $r$ is a:
+Assume a constitutional blocklace $B$ with prevailing constitution $(P,\sigma)$. A round $r$ in a wave $W\subseteq B$ is advanced if it is a:
   1. 1.
-    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a leader block or (iii) follows a quiescent wave and includes at least one block, or a
+    first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a formal leader block or (iii) is preceded by a quiescent wave in $B$ and includes at least one block, or a
   2. 2.
-    second or third round that includes a $\sigma$-supermajority of blocks among $P$.
-Round 00 is trivially advanced in all blocklaces.
-Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol.
-Formally, an ordinary block $b$ is valid if it is well-formed and round $d(b)-1$ is advanced in $[b]$.
-Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block.
-In the low-throughput scenario, the previous wave ending in round $r$ is quiescent, the spontaneous leader block (red) may be produced at any time after the previous round $r$ has advanced, and all other blocks are empty, thus the wave $w$ is also quiescent.
-In the high-throughput scenario, the formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced, and all (or many) of the blocks are non-empty.
-Figure 1: A good-case wave.
-Dissemination. The protocol uses three types of dissemination-inducing blocks: ack, nack, and nudge, defined below; other blocks are referred to as ordinary.
+    second or third round that includes blocks by a $\sigma$-supermajority among $P$.
+Round 0 is trivially advanced.
+Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally:
+Definition 3.5 (Valid Block).
+An ordinary block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
+Dissemination.
+The protocol uses two types of dissemination-inducing blocks: nack and inform; other blocks are referred to as ordinary.
 Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient.
 Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
-The Grassroots Consensus protocol we present in Section 2.3 assumes reliable links and uses nack and nudge blocks for dissemination.
-Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Grassroots Consensus for this model achieves $O(n)$ amortized communication complexity in the good case during high-throughput.
-The UDP-ready protocol (given in Section 4) uses also ack blocks for dissemination.
+Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
+Protocols such as TCP achieve the reliable transfer of a sequence of bytes from the sender to the recipient by employing ack and nack messages for transmission control.
+The purpose of an ack message is to acknowledge the receipt of a datagram, and of a nack message to report an out-of-order datagram to its sender, e.g., in the case of receiving the $(k+1)^{th}$-datagram before receiving the $k^{th}$-datagram.
+In the context of the blocklace, an ‘out-of-order’ block is a block received before some block it points to, and the purpose of a nack-block is indeed to report of such an out-of-order block to its creator (which is not necessarily its sender).
 nack blocks are used to achieve reliable broadcast of ordinary blocks.
-They are needed in order to address partial dissemination by Byzantine agents.
+They are needed in order to rectify partial dissemination by Byzantine agents.
 For example, consider a Byzantine agent that sends a block $b$ to correct agent $p$ but not to another correct agent $q$, and $p$ adds $b$ into its blocklace and sends a block $b^{\prime}$ that depends on $b$.
@@
 To this end, $q$ sends a nack block to $p$ specifying the blocks it is missing.
-Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony, even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$.
+Note that out-of-order blocks may naturally arise on a reliable network even during periods of synchrony and even when there are no Byzantine agents, because there is no bound on how quickly messages can arrive — in the example above, $b$’s sender might be correct and $b$ may still be under way to $q$.
 We therefore introduce a delay before sending a nack.
-Thus, in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no further communication is needed.
+The result is that in the good case, when the network is synchronous and there is no partial-dissemination by Byzantine agents, every correct agent sends its newly created blocks to all other agents and no nack-blocks or other communication is needed.
 The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$.
-If $[b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$.
-While obviously $p$ cannot know all the blocks it does not know, it can know the boundary of its knowledge—where $p$’s ignorance starts in relation to $b$.
-These are the tips of $[b]\setminus(B\cup\{b\})$.
-Note that if no block pointed to by $b$ is in $B$, then the pointers in $b$ point exactly to tips in $[b]\setminus(B\cup\{b\})$.
+If $b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$.
+These are the tips of $b]\setminus(B\cup\{b\})$.
 Hence the following definition:
 Definition 3.6 (nack-block).
-Given a blocklace $B$, a nack-block for a block $b=(h,x,H)$ is the block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $[b]\setminus(B\cup\{b\})$.
-Note that as defined above, $[b^{\prime}]\subseteq[b]$.
-Also note that, given a block $b=(h,x,H)$ and a blocklace $B$, the tips of $[b]\setminus(B\cup\{b\})$ are observed directly from $b$: These are the pointers in $H$ that are dangling in $B$, namely point to blocks not in $B$.
-These dangling pointers attest to the ‘knowledge gap’ of the recipient of $b$, namely, which blocks need to be added to $B$ so that $b$ will be in order following $B$.
-If there is no such gap, namely $[b]\subseteq B\cup\{b\}$, then the nack-block for $b$ and $B$ is undefined.
-In the reliable network model, the reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block is in fact an ack-block, as it attests to all the blocks known to $p$ at the time of its creation; ack-blocks will be added in the UDP-ready protocol described in Section 4, when eventually-reliable networks are considered.
-nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate progress using nudge-blocks.
-Definition 2.5 (nudge-Block).
-Given a blocklace $B$, a nudge-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{nudge}$ and pointers $H$ to the blocks of depth $r$ in $B$.
-If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks, $p$ sends the leader a nudge-block pointing to the $r$ blocks in its blocklace.
-If the leader indeed misses any of these blocks, which after GST may happen due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
-2.3 The Protocol
-Here we present Grassroots Consensus for eventual synchrony with a reliable network.
-We prove the protocol’s safety and liveness in Section 3, and present its UDP-ready extension in Section 4.
-In this section, we present the protocol using arbitrarily-fixed values for the parameters $(P,\sigma,\Delta)$.
-These can be changed using constitutional amendments, discussed in Section 6.
-In the Grassroots Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. The protocol is event-driven as described in Algorithm 1.
-Overview.
-Each three consecutive rounds are grouped into a wave, which aims to finalize one block of its first round.
-Similarly to Morpheus [28], the protocol has two modes of operation, high-throughput and low-throughput, but here they are implicit— depending on whether the preceding wave is quiescent.
-In high-throughput mode, in the good-case scenario, only the leader sends a first round block.
-If the leader is faulty or does not communicate with other agents in a timely manner (bad-case scenario), other agents time-out on the leader and send blocks of their own.
-Agents wait either for a first-round leader block or for a supermajority of blocks in a round before producing a block in the next round.
-Low-throughput is manifested by a quiescent wave, which finalizes its first-round spontaneous leader using two rounds of empty blocks.
-Following a quiescent wave, any agent may initiate a new wave by issuing a non-empty block.
-If two such blocks collide the protocol enters high throughput mode until, if ever, it quiesces again.
-Initialize: $D\leftarrow\phi$, $B\leftarrow\phi$, $\emph{payload}\leftarrow\bot$.
-Let $r$ always be the maximal advanced round in $B$, which is initially 0.
-Receive: Upon receipt of $q$-block $b$, if $b$ a nack-block then judiciously send $[b]$ to $q$,
-else if $b$ a nudge-block such that $p=\emph{leader}(d(b))$ and $[b]\not\subseteq B\cup\{b\}$ then send once a nack-block for $b$ to $q$,
-else if $b$ is valid then add $b$ to $D$.
-Accept & Nack: If $[b]\subseteq B\cup\{b\}$ for some $b\in D$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
-else if a $q$-block $b^{\prime}$ was added to $D$ more than $\Delta$ time ago then send once a nack-block for $b^{\prime}$ to $q$.
-Issue: Issue once a new block if $r+1$ is a second or third round or a first round that follows a wave that is: 1.
-quiescent (or $r=0$) and $\emph{payload}\neq\bot$, or 2.
-non-quiescent and (i) $p=\emph{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$.
-Nudge: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once to $\emph{leader}_{P}(r+1)$ a nudge-block for round $r+1$.
-Output: If $B$ contains a final block $b$ of depth greater than the previous final block (if any) then output every block in $\tau(b)$ unless it has already been output.
-issue a new block means create a block $b$ with payload and pointers to the tips of $B_{r}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$.
+Given a blocklace $B$, a nack-block by $p$ for a block $b=(h,x,H)$ such that $b]\not\subseteq B\cup\{b\}$ is the $p$-block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $b]\setminus(B\cup\{b\})$.
+Note that as defined above, $b^{\prime}]\subseteq[b]$, and if $b$ has no pointers dangling in $B$ then a nack-block for $b$ and $B$ is undefined.
+If the nack-block $b^{\prime}$ is defined, then its pointers $H^{\prime}$ are precisely the non-empty subset of $H$ that is dangling in $B$.
+The reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block functions as an ack-block, as it attests to all the blocks known to $p$ at the time of its creation.
+nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it.
+In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
+Definition 3.7 (inform-Block).
+Given a blocklace $B$, an inform-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{inform}$ and pointers $H$ to the blocks of depth $r$ in $B$.
+If an agent $p$ suspects that the leader of the next round $r+1$ does not proceed since it is missing some round $r$ blocks known to $p$, then $p$ sends the leader an inform-block pointing to the $r$ blocks in its blocklace.
+If the leader indeed misses any of these blocks, which may happen even after GST due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
+3.3 Single-Epoch Constitutional Consensus Pseudocode
+In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$.
+An epoch identifier $e$ points to the epoch’s constitutional genesis block.
+We use $B^{e}$ to denote the blocklace of epoch $e$ in $B$, which is the subset of $B$ that observes $e$.
+The protocol is event-driven.
+Algorithm 1 describes a single-epoch Constitutional Consensus, in which the prevailing constitution is fixed throughout the run.
+Initialisation.
+The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$.
+The initialisation of subsequent epochs is discussed in Algorithm 2.
+Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Receive: Upon receipt of a $q$-block $b$, $q\in P$
+if $b$ is a nack-block for a $p$-block then judiciously send $[b]$ to $q$,
+else if $b$ an inform-block and $b]\not\subseteq B\cup\{b\}$
+——- then send once a nack-block for $b$ to $q$,
+else if $b$ is a valid ordinary block then $D\leftarrow D\cup\{b\}$.
+Accept or Nack: For every $q$-block $b\in D$, $q\in P$,
+if $b]\subseteq B\cup\{b\}$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
+else if a $b$ was added to $D$ more than $\Delta$ time ago
+——- then send once a nack-block for $b^{\prime}$ to $q$.
+Issue: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: 1.
+quiescent and $\emph{payload}\neq\bot$, or 2.
+non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$.
+Issue Backlog: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block.
+Inform Leader: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$.
+Output: If $B$ contains a final block $b$ of depth greater than the previous final block then output every block in $\tau(b)$ (Definition 3.8) unless it has already been output.
+issue a round $k$ block means create a block $b$ with payload and pointers to the $P$-blocks that are tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$.
 judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
-Algorithm 1 Grassroots consensus protocol, code for agent $p\in P$, given $(P,\sigma,\Delta)$.
+Algorithm 1 Constitutional Consensus for a single epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$. Code for agent $p\in P$.
 Delays. The protocol makes judicious use of delays, to ensure both liveness and efficiency. Here is the rationale for the various protocol delays, all expressed as multiples of $\Delta$:
@@
 In the good case we wish no nack-blocks to be sent at all.
-Agents wait $\Delta$ before complaining, so if a correct agent sends a block $b$ to $p$ and $q$ at the same time, $q$ may receive $b$ up to $\Delta$ sooner than $q$ and send a block $b^{\prime}$ that depends on $b$ to $p$; but $p$ will receive $b$ within $\Delta$ from receiving $b^{\prime}$.This delay is an optimization to avoid sending nack-blocks in the good case; eliminating it does not hamper safety or liveness.
-$2\Delta$ for nudging the ensuing round’s leader.
+Agents wait $\Delta$ before complaining, so if a correct agent sends a block $b$ to $p$ and $q$ at the same time, $q$ may receive $b$ up to $\Delta$ sooner than $q$ and send a block $b^{\prime}$ that depends on $b$ to $p$; but $p$ will receive $b$ within $\Delta$ from receiving $b^{\prime}$.
+This delay is an optimization to avoid sending nack-blocks in the good case; eliminating it does not hamper safety or liveness.
+$2\Delta$ for informing the ensuing round’s leader.
 In the good case whence all agents are correct after GST, whenever an agent receives a block, all other agents receive it within $\Delta$ time.
@@
 This last third case can happen after GST even if the leader is correct, due to Byzantine partial dissemination of third-round blocks.
-Since $p$ knows that the previous third round is advanced, it sends a nudge-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing.
-This $2\Delta$ delay is an optimization to avoid spurious nudges in the good case; eliminating it does not hamper safety or liveness.
+Since $p$ knows that the previous third round is advanced, it sends an inform-block to the leader, which will cause the leader to send a nack-block requesting the blocks it is missing.
+This $2\Delta$ delay is an optimization to avoid spurious informs in the good case; eliminating it does not hamper safety or liveness.
 $9\Delta$ for an unresponsive leader.
-Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, t all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized.
+Following a non-quiescent wave after GST, if the leader $p_{\ell}$ is correct, all correct agents should wait for $p_{\ell}$’s first round block before proceeding to the second round in order to allow that block to be finalized.
 But because $p_{\ell}$ may be faulty, they need to eventually time-out and stop waiting.
 To determine the time-out, we consider the worst-case latencies as follows: Assume a third round $r$ is advanced at $p$ at time $t$.
-If $p$ does not receive a leader block by $t+2\Delta$, it nudges $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks.
+If $p$ does not receive a leader block by $t+2\Delta$, it informs $p_{\ell}$, which issues a nack-block, upon receipt of which $p$ sends $p_{\ell}$ any missing blocks.
 This exchange takes at most $3\Delta$, so by time $t+5\Delta$, $p_{\ell}$ issues a leader block $b$ for round $r+1$, which is received by $p$ by time $t+6\Delta$.
@@
 Protocol walk-through.
-Incoming blocks are inserted to $D$; if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $B$.
-The variable $r$ always holds the highest advanced round in $B$.
+Receive: valid incoming ordinary blocks are inserted to $D$.
+Accept or Nack: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
+Issue: The variable $r$ always holds the highest advanced round in $B$, and an agent typically issues blocks in round $r+1$.
+This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced.
+This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
+The only exception (to sending in round $r+1$) is the Issue Backlog rule: In case $p$ has payload to send, has not sent a round $r$ block, and is not supposed to send a round $r+1$ block, $p$ issues a round $r$ block to ensure its payload is disseminated.
 If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block.
 If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode.
-During low throughput (initially and following a quiescent wave), $p$ issues a block only if it has payload to send.
-During high-throughput, $p$ issues a block in case (i) it is the leader, or (ii) it times-out on the leader.
-An agent $p$ always issues blocks that follow the latest advanced round $p$ saw.
-This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced.
-This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
-After GST, whenever there is a wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block.
-Therefore, no agent can produce a valid second round block that does not endorse the leader block.
-So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized.
-Ordering with $\tau$.
-Once the protocol finalizes a block $b$, it orders the predecessors of $b$ that have not been ordered yet using the function $\tau$.
-The definition of $\tau$ is similar to Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks.
-The output of $\tau(b)$ is a sequence that includes all blocks observed by $b$, with equivocating blocks excluded.
-In Section 3 we prove that $\tau$ is monotonic and safe.
-The monotonicity of $\tau$ ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor.
-Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any.
-The safety of $\tau$ means that blocks are ordered the same way at all correct agents.
-In anticipation of the protocol being constitutional, we already assume the notion of the prevailing constitution.
-For now, assume $(P,\sigma,\Delta)$ is fixed arbitrarily for all blocks.
-In Section 6, we define the constitution prevailing in a block, and have each call (initial and recursive) to $\tau$ with a block $b$ employ the parameters of the constitution prevailing in $b$.
+During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send.
+During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader.
+Here, a long timeout is used ($9\Delta$), before which we try to inform the leader that the round has advanced to give it a chance to progress.
+Inform Leader: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
+Output: Once the protocol finalizes a block $b$, it orders the blocks approved by $b$ that have not been ordered yet using the function $\tau$.
+The definition of $\tau$ is essentially-identical to that of Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks.
+The output of $\tau(b)$ is a sequence that includes all blocks approved by $b$.
+Formally:
 Definition 3.8 ($\tau$).
-We assume a fixed topological sort function $\textit{xsort}(b,B)$ (exclude and sort) that takes a block $b$ and a blocklace $B$, and returns a sequence that preserves the ‘observes’ partial order and includes all the blocks in $B$ that are approved by $b$.
+Assume a safe constitution $(P,\sigma)$ and any fixed topological sort function $\textit{xsort}(b,B,B^{\prime})$ (exclude and sort) that takes a block $b$ and two blocklaces $B\supset B^{\prime}$, and returns a sequence that preserves the ‘observes’ partial order and includes all the non-empty blocks in $B\setminus B^{\prime}$ that are approved by $b$.
 The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
-$\tau(b):=\begin{cases}\tau(b^{\prime})\cdot\textit{xsort}(b,[b]\setminus[b^{\prime}])\text{\ \ if $b^{\prime}$ is the maximal-depth block ratified in $[b]$}\\ \ \ \ \ \ \ \ \ \ \textit{according to the constitution prevailing in $b$}\\ \textit{xsort}(b,[b])\text{\ \ \ \ \ if no such $b^{\prime}$ exists }\\ \end{cases}$
+[…]
 When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$.
@@
 In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
+Proposition 3.9.
+If the prevailing constitution is safe then $\tau(b)$ does not include equivocations.
+Proof 3.10.
+Consider an equivocation $b_{1},b_{2}\in[b]$.
+Consider the first call $\textit{xsort}(b^{\prime},[b^{\prime}],[b^{\prime\prime}])$ for which at least one of $b_{1}$ or $b_{2}$ is in $X:=[b^{\prime}]\setminus[b^{\prime\prime}]$, and the following two cases:
+  1. 1.
+    Both $b_{1},b_{2}\in X$: Then $b^{\prime}$ approves neither hence neither will be in the output of xsort and hence nor in the output of $\tau$.
+  2. 2.
+    Only one of them is in $X$, wlog $b_{1}\in X$, and since this is the first such call to xsort it must be that $b_{2}\in[b^{\prime\prime}]$.
+Then $b^{\prime}$ observes both $b_{1}$ and $b_{2}$ and hence will approve neither, whereas $b^{\prime\prime}$ observes only $b_{2}$, therefore will approve it and therefore only $b_{2}$ will occur in the output of $\tau$, if at all (assuming there are no other equivocations with $b_{2}$).
+In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$.
+Safety and liveness.
+We assume that the prevailing constitution is safe and live (Definition 2.2).
+In Section 4 we prove that $\tau$ is monotonic.
+This ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor.
+Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any.
+In addition, as we shall see, during a protocol run the outputs of calls to $\tau$ by correct agents always occur in the same order, which is the basis for protocol safety.
+Liveness is achieved after GST.
+Whenever there is a post-GST wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block.
+Therefore, no agent can produce a valid second round block that does not endorse the leader block.
+So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized.
+Once any agent has payload to send following a quiescent wave, the agent issues a block, and either the block is finalized in the same wave or the wave is non-quiescent and the protocol switches to high-throughput mode.
 Latency and complexity.
-Regarding latency, we note that after GST and in the good case a (formal or spontaneous) leader block is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
-Regarding communication complexity, during low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all.
-In Morpheus [28], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
+Regarding latency, we note that after GST and in the good case a leader block (formal or spontaneous) is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
+Regarding communication complexity: During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all.
+In Morpheus [23], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
 A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead.
 Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block.
-A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, should carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block.
+A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block.
 Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity.
-To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.k
+To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.
 During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
-3 Safety and Liveness of Grassroots Consensus
-Here we prove the safety and liveness of the Grassroots Consensus protocol.
-We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
-Definition 3.1 (Safety and Liveness of an Ordering Consensus Protocol).
-An ordering consensus protocol is:
-Safe if output sequences of correct agents are always consistent.
-Live if every non-empty block issued by a correct agent is eventually included in the output of every correct agent.
-Next we prove safety and liveness of the basic Grassroots Consensus protocol.
+4 Safety and Liveness of Single-Epoch Constitutional Consensus
+Here we prove the safety and liveness of the single-epoch Constitutional Consensus protocol presented in Algorithm 1.
+In this case, the members of the founding constitution $C=(P,\sigma,\Delta)$ participate in the protocol for the entire run.
+The output sequence at every correct participant begins with $C$ and does not include additional constitutions.
 4.1 Safety
+The safety properties of Definition 2.4 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
+{observation}
+If the output sequences of every two correct agents are consistent with each other, then single-epoch consensus ensures safety.
+Proof 4.1.
+Validity trivially holds because every agent participates in every epoch.
+Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency.
 Theorem 4.2.
-Grassroots Consensus is safe.
-To prove the theorem, we follow a condensed and simplified proof of the safety of Cordial Miners.
-First, observe that Grassroots Consensus produces only $f$-safe blocks; if a block created by the protocol is not $f$-safe it means that its closure includes equivocations by more than $f$ agents, a contradiction.
+If the founding constitution is safe then single-epoch Constitutional Consensus is safe.
+First, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
+We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
 For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
 Definition 4.3 ($\tau$ Safety).
-A blocklace $B$ is $\tau$-safe if every final block in $B$ is ratified by every higher-depth block ratified by some block in $B$.
+A blocklace $B$ is $\tau$-safe if for every wave $W\subset B$ with a block $b$ final in $W$ and any subsequent wave $W^{\prime}\subset B$ with a block $b^{\prime}$ ratified in $W^{\prime}$, $b$ is ratified in $W\cap[b^{\prime}]$.
 Proposition 4.4.
-An f-safe blocklace is $\tau$-safe.
-Proof 3.5.
-Let $B$ be an $f$-safe blocklace, let $b$ be a block of wave $k$ final in $B$ and $b^{\prime}$ a block of a later wave $k^{\prime}>k$ ratified by some block in $B$, which means that $b^{\prime}$ is endorsed by a $\sigma$-supermajority in $B$.
-We have to show that $b$ is ratified by $b^{\prime}$.
+A constitutional blocklace with a safe constitution is $\tau$-safe.
+Proof 4.5.
+Let $B$ be a constitutional blocklace with a safe constitution $(P,\sigma)$, $W\subset B$ a wave with a block $b$ final in $W$, and $W^{\prime}\subset B$ a subsequent wave with a block $b^{\prime}$ ratified in $W^{\prime}$.
+We have to show that $b$ is ratified in $W\cap[b^{\prime}]$.
 Since $b^{\prime}$ is ratified by some block it means that it is valid, since correct agents do not include invalid blocks in their blocklace, let alone endorse them, and a $\sigma$-supermajority includes (a majority of) correct agents.
 Hence, being a first-round block of a wave, $b^{\prime}$ must observe a $\sigma$-supermajority of the last round of the previous wave, let’s call it the blue supermajority.
-Then the blue supermajority and the finalizing $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection.
-Namely, there is a correct agent $p$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $k^{\prime}$ happens to be $k+1$), which in turn must observe a supermajority of blocks that endorse $b$.
+Then the blue supermajority and the finalising $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection.
+Namely, there is a correct agent $p\in P$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $W$ and $W^{\prime}$ are consecutive waves), which in turn must observe a $\sigma$-supermajority of blocks that endorse $b$.
 Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
 Proposition 4.6 ($\tau$ Monotonicity).
-$\tau$ is monotonic wrt $\supset$ over $f$-safe blocklaces.
-Proof 3.7.
-We have to show that if $B\subset B^{\prime}$ for $f$-safe blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
-By Observation 2.2, all the blocks final in $B$ are also final in $B^{\prime}$.
-By Observation 2.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their round number.
+$\tau$ is monotonic wrt $\supset$ over constitutionally-safe blocklaces.
+Proof 4.7.
+We have to show that if $B\subset B^{\prime}$ for blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
+By Observation 3.2, all the blocks final in $B$ are also final in $B^{\prime}$.
+By Observation 3.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth.
 If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially.
@@
 {observation}
-If the union $B\cup B^{\prime}$ of two blocklaces is $f$-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
-Proof 3.8.
-By monotonicity of $\tau$ over safe blocklaces, $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
-Proof 3.9 (Proof of Theorem 3.2).
-Since in a run of Grassroots Consensus there are at most $f$ faulty agents in total, the union of the blocklaces of two correct agents is always $f$-safe.
-Hence, by Observation 3.1, the outputs of any two correct agents running Grassroots Consensus are always consistent.
+If the union $B\cup B^{\prime}$ of two blocklaces is constitutionally-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
+Proof 4.8.
+By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition 4.6), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
+Proof 4.9 (Proof of Theorem 4.2).
+By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe.
+Hence, by Observation 4.1, the outputs of any two correct agents running Constitutional Consensus are always consistent.
+By Observation 4.1, the protocol is safe.
 4.2 Liveness
-Next, we discuss liveness. We first prove that Grassroots Consensus disseminates all correct agents’ blocks and the blocks they depend on.
+Next, we discuss liveness.
+Having proved safety, we know that all output sequences of correct agents are subsequences of a common sequence.
+Let $T$ be the minimal such sequence.
+Then the set of elements in $T$ is the union of all the elements in output sequences of correct agents.
+Because all members of $P$ participate in all epochs, Epoch Liveness requires that all their output sequences include all the transactions in $T$, which are all the transactions output by correct agents.
+Constitution Amendment Liveness trivially holds in the single-epoch case.
+We conclude with the following observation:
+{observation}
+[Single-epoch Liveness] A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
+Theorem 4.10.
+Single-epoch Constitutional Consensus is live.
+We first prove that Constitutional Consensus disseminates all correct agents’ blocks and the blocks they depend on.
 Proposition 4.11 (Dissemination).
-If a block $b$ is issued by a correct agent in Grassroots Consensus then $[b]$ is eventually received by all correct agents.
-Proof 3.11.
+If a block $b$ is issued by a correct agent in Constitutional Consensus then $[b]$ is eventually received by all correct agents.
+Proof 4.12.
 By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$.
-Since by assumption $b$ (and possibly other blocks in $[b_{p}]$) have not been received by $q$, then upon receipt of $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$.
-Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $[b_{q}]$, which includes $b$, to $q$.
+Since by assumption $b$ (and possibly other blocks in $b_{p}]$) have not been received by $q$, then after receiving $b_{p}$, $q$ issues a nack-block $b_{q}$ for which $b\in[b_{q}]$.
+Upon receipt of $b_{q}$ by $p$, $p$ will send (once) $b_{q}]$, which includes $b$, to $q$.
 Since the network is reliable, $q$ will eventually received $b$, a contradiction.
@@
 {observation}
-[Sequential Advance] If round $r$ is advanced in a blocklace $B$, then every round $0<r^{\prime}<r$ is also advanced in $B$.
+[Sequential Advance] If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
 We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress.
-Because Grassroots Consensus is quiescent, agents do not have to generate infinitely many blocks.
+Because Constitutional Consensus is quiescent, agents do not have to generate infinitely many blocks.
 But if some agent issues a block in a round, that round eventually advances.
 Proposition 4.13 (Advance).
-If a correct agent $p$ sends a block in some round $r$ in Grassroots Consensus, then $r$ eventually becomes advanced at every correct agent.
-Proof 3.13.
-Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, $r-1$ is advanced in $p_{1}$’s blocklace.
+If a correct agent $p$ sends a block in some round $r$ in Constitutional Consensus, then $r$ eventually becomes advanced at every correct agent.
+Proof 4.14.
+Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, r1r-1 is advanced in $p_{1}$’s blocklace.
 Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$.
-By Observation 3.11, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace.
-By Proposition 3.10, this means that no correct agent sends blocks in rounds $>r$.
+By Observation 4.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace.
+By Proposition 4.11, this means that no correct agent sends blocks in rounds >r>r.
 Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition 4.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
-By Proposition 3.10, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round $r-1$ is advanced in $p_{3}$’s blocklace.
+By Proposition 4.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
 In case $r$ is a first round following a quiescent wave at $p_{3}$, $r$ advances at $p_{3}$ upon receipt of $b$, and $p_{3}$ issues a round $r+1$ block, a contradiction.
@@
 Proposition 4.15 (Finite implies quiescence).
-Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
-Proof 3.15.
-Let $r$ be the last advanced round in $B$.
-(Note that round 00 is by definition advanced, so $r$ is always defined).
-Assume by way of contradiction that $B$ is not quiescent, then $p$ eventually issues a round $r+1$ block $b$.
-By Proposition 3.12, $r+1$ eventually becomes advanced at $p$.
+Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
+Proof 4.16.
+Assume by way of contradiction that $p$’s latest advanced round in the run, $r$, is not quiescent.
+Then, by the Issue rule, $p$ eventually issues a block in round $r+1$ (at most $9\Delta$ time after $r$ is advanced).
+Proposition 4.17 (Block issuance).
+If an agent $p$ has payload to send, then $p$ eventually issues a block with this payload.
+Proof 4.18.
+Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload.
+First, consider the case that $p$’s rounds do not advance in some suffix of the run.
+Then by Proposition 4.15, $p$’s blocklace is quiescent.
+Hence, condition 1 of the Issue rule is satisfied, and the payload is sent.
 A contradiction.
-Proposition 3.16 (Liveness in finite runs).
-Consider a run of Grassroots Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run.
-Then all non-empty ordinary blocks sent by correct agents in this run are ordered in $B$.
-Proof 3.17.
+Therefore, $p$’s advanced rounds increase indefinitely.
+Consider a time when $p$’s highest advanced round $r>r_{0}$ and the payload has not been sent.
+Then $p$ issues a block either in round $r+1$ (by the Issue rule) or in round $r$ (by the Issue backlog rule).
+A contradiction.
+Proposition 4.19 (Transaction Liveness in finite runs).
+Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all input transactions of correct agents are ordered in $B$.
+Proof 4.20.
 By Proposition 4.11, $B$ includes all ordinary blocks sent by correct agents.
@@
 Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$.
-Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalizing.
-Because quiescence is not monotonic, we consider here only waves $k$ for which the preceding wave, $k-1$, is not quiescent at any correct agent at any time.
+Moreover, by Proposition 4.17, no agents have payloads to send.
+It follows that all input transactions have been sent in ordinary blocks and ordered.
+Proposition 4.21 (Epoch Liveness in finite runs).
+Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all transactions output by correct agents are ordered in $B$.
+Proof 4.22.
+Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run.
+Because $t$ is at some point final in $p_{1}$’s blocklace, its blocklace eventually includes a block $b$ with $t$ in its payload and a super-majority of blocks that observe $b$.
+This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$.
+By Proposition 4.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$.
+Because $B_{2}$ is finite in a suffix of the run, by Proposition 4.15, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
+Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalising.
+Because quiescence is not monotonic, we consider here only waves $k$ for which the predecessor wave, k1k-1, is not quiescent at any correct agent at any time.
 Proposition 4.23 (First round progress after GST).
-Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round $r-1$ advances at time $t$>GST, and assume that the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time.
+Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt>GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time.
 Then $p$ sends a block in round $r$ if and only if it is the round’s leader.
 Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
-Proof 3.19.
+Proof 4.24.
 Let $p_{\ell}$ be the correct leader of round $r$.
-If $p=p_{\ell}$ then once $r-1$ is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
-Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ a nudge-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$.
-Because round $r-1$ is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$.
+If $p=p_{\ell}$ then once r1r-1 is advanced, $p$ immediately issues a leader block for round $r$, causing round $r$ to advance, and issues a block endorsing it in round $r+1$.
+Otherwise, $p$ waits for a block from $p_{\ell}$ until time $t+2\Delta$ and then sends $p_{\ell}$ an inform-block, which is received by $t+3\Delta$, causing $p_{\ell}$ to send $p$ a nack-block and get the blocks it is missing by time $t+5\Delta$.
+Because round r1r-1 is advanced and its wave is not quiescent, $p_{\ell}$ sends a leader block, which is received by $p$ by time $t+6\Delta$.
 Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
@@
 Let $r$ be a first round with a correct leader s.t.
-round $r-1$ advances after GST and the wave that ends in round $r-1$ is not quiescent in any correct agent’s blocklace at any time.
+round r1r-1 advances after GST and the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time.
 Then a round $r$ leader block is finalized in its wave at all correct agents.
-Proof 3.21.
+Proof 4.26.
 By Proposition 4.23, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$.
@@
 Proposition 4.27 (Infinite quiescent waves).
-Consider a run of Grassroots Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time.
+Consider a run of Constitutional Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time.
 Then all the blocks of correct agents are ordered by all agents in this run.
-Proof 3.23.
+Proof 4.28.
 Let $B$ be the (infinite) union of all blocklaces in the run.
@@
 Therefore, it is approved by all correct agents’ blocks from some wave onward.
-Because every quiescent wave is finalizing, and finalization is monotonic, there are infinitely many finalizing waves in $B$.
+Because every quiescent wave is finalising, and finalization is monotonic, there are infinitely many finalising waves in $B$.
 Let $b_{2}$ be a final block that approves $b_{1}$ in some wave $k$.
@@
 Observe that since $b_{2}$ is finalized in wave $k$, it is approved by a super-majority in wave $k$.
-Because every valid first round block approves a super-majority in the preceding round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
-Consider a finalizing wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$.
+Because every valid first round block approves a super-majority in the previous round and there is at least one correct agent at the intersection of any two super-majorities, every valid first round block in any round $k^{\prime}>k$ observes $b_{2}$.
+Consider a finalising wave $k^{\prime}>k$ in $B$ and let $b_{3}$ be the first-round block that is finalized in wave $k^{\prime}$.
 Then $b_{3}$ is endorsed by a super-majority in the second round.
 Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 4.11).
-If $b_{3}$ is not a leader block, then to be validly endorsed, $[b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
+If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
 By the same token, if any non-leader first round block $b$ is endorsed by a valid block of any agent in the second round of wave $k+1$, then $[b]$ is quiescent.
@@
 We are now ready to prove the liveness theorem.
-Theorem 3.24.
-Grassroots Consensus is live.
-Proof 3.25.
-Let $b$ be a non-empty ordinary $p$-block produced by a correct agent $p$ in a run of Grassroots Consensus.
-By Proposition 3.10, every correct agent receives and incorporates $b$ in its local blocklace.
+Proof 4.29 (Proof of Theorem 4.10).
+By Observation 4.2, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent.
+By Proposition 4.17, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks.
+Similarly, every output transaction of a correct agent is included in some ordinary block.
+Let $b$ be a non-empty ordinary $p$-block included in some correct agent’s blocklace in a run of Constitutional Consensus.
+If $p$ is correct, then by Proposition 4.11, every correct agent receives and incorporates $b$ in its local blocklace.
+Otherwise, $b$ is finalized in some blocklace, implying that it is approved by at least one $q$-block $b^{\prime}$ by a correct agent $q$.
+Again, by Proposition 4.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
 We consider three cases: First, assume the run is finite.
-By Proposition 3.16, all non-empty ordinary blocks sent by correct agents in this run are ordered at all correct agents, and the theorem holds.
+By Propositions 4.19 and 4.21, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds.
 Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time.
-By Proposition 3.22, $b$ is ordered.
+By Proposition 4.27, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
 Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time.
@@
 By Proposition 4.25, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
-4 UDP-Ready Grassroots Consensus for Eventually-Reliable Networks
-Grassroots platforms and protocols are geared for smartphone-based serverless implementation, and hence depend on smartphone-to-smartphone communication.
-Smartphones often reside behind NATs and Firewalls, which in turn are programmed with policies that make smartphones finding each other difficult.
-To overcome this, servers with known IP addresses (or domain names), running the STUN protocol help devices behind NATs find each other.
-However, once they do, establishing a direct TCP connection is difficult, and hence direct phone-to-phone communication typically employs UDP.
-Grassroots Consensus (Algorithm 1) was specified assuming a reliable network.
-Here, we relax the assumption to eventually-reliable networks, and extend the protocol to be UDP-ready, by adding ack-blocks and block resending.
-This protocol resends only ordinary blocks, as the reliable communication of dissemination-inducing blocks (nack, ack, and nudge) before GST is not required.
-Note that the protocol is very different from the trivial solution of “implementing TCP upon UDP”, as during correct execution the same block can arrive via multiple paths and blocks may arrive in many possible orders.
-The protocol does not need to to impose an ordering on agent-to-agent block communication.
-Definition 4.1 (ack-block).
-Given a block $b=(h,x,H)$, an ack-block for $b$ is the block $b^{\prime}=(h^{\prime},(\textsc{ack},h),\emptyset)$.
-The changes to make Grassroots Consensus UDP-ready are adding ack to the Receive rule, adding resending to the nack-rule, a new Resend rule, and revising the notion of judiciously send, as detailed in Algorithm 2 below (additions to Algorithm 1 are highlighted).
-In Algorithm 2, $k\geq 2$ is a parameter chosen to taste.
-In principle, it could be added as the fourth parameter of the constitution and amended similarly to $\Delta$.
-Receive: Upon receipt of $q$-block $b$, if $b$ a nack-block then judiciously send $[b]$ to $q$,
-else if $b$ a nudge-block such that $p=\emph{leader}(d(b))$ and $[b]\not\subseteq B\cup\{b\}$ then send once a nack-block for $b$ to $q$,
-else if $b$ is valid then add $b$ to $D$ and send an ack-block for $b$ to $q$.
-Accept & Nack: If $[b]\subseteq B\cup\{b\}$ for some $b\in D$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$, else if a $q$-block $b$ was added to $D$ more than $\Delta$ time ago, and no nack-block for $b$ was sent during the last $k\Delta$, then send a nack-block for $b$ to $q$.
-Resend: If $k\Delta$ has passed since the most-recent ordinary $p$-block $b$ was sent to $q$, no ack-block or nack-block was received for $b$, and $b\notin[b_{q}]$ for any $q$-block $b_{q}\in B\cup D$, then resend $b$ to $q$.
-judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) an ack-block or a nack-block was received for $b^{\prime}$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
-Algorithm 2 UDP-Ready Extension to Grassroots Consensus.
-Mobile devices.
-Smartphones are mobile, and when they move their IP address may change.
-Hence, even if a phone-to-phone connection is established, it has to be re-established every time one of the phones changes their IP address.
-Blocklace dissemination can support mobile agents recovering each other’s IP address via a joint stationary friend [41].
-The problem is already present at protocols of lower-level than consensus such as the grassroots social graph [44] and the blocklace-based solution presented there [41] is applicable here as well.
-Safety and liveness.
-The UDP-ready extension (Algorithm 2) to Grassroots Consensus only relates to dissemination-inducing blocks (an ack-block and its handing) and does not affect ordinary blocks.
-Hence the definitions, propositions, and proofs regarding the safety of Grassroots Consensus—all stated in terms of ordinary blocks only—remain valid verbatim.
-Safety of the UDP-ready extension is therefore a corollary of Theorem 3.2.
-Regarding liveness, prior to GST ordinary blocks issued by correct agents are resent to correct agents until received.
-Thus, there is a time $t\geq\text{GST}$ by which every block sent by a correct agent priot to GST either has been received, or has been resent after GST but before $t$.
-Thus, any message sent before $t$ would be received by $t+\delta$.
-After GST, the two models under consideration are the same.
-Thus, the time $t$ in the UDP-ready protocol has the same properties as GST in the Grassroots Consensus protocol: Any message between correct agents, if sent before $t$ arrives by $t+\delta$, and if sent after $t$ arrives within $\delta$.
-Hence, the liveness proofs of Grassroots Consensus hold for the UDP-ready protocol, replacing GST by $t$ thus defined, with the liveness of the UDP-ready protocol being a corollary of Theorem 3.24.
-5 Grassroots Consensus is Grassroots
-We recall the formal definition of a grassroots protocol from reference [44], instantiate it informally (bypassing the formalities of the definition) to the constitutional Grassroots Consensus protocol, and then argue that the protocol is grassroots.
-Definition 5.1 (Oblivious, Interactive, Grassroots).
-A protocol $\mathcal{F}$ is: 1.
-oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, a run of the protocol over $P$ can proceed just the same in the presence of agents in $P^{\prime}$.
-2.
-interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every run of the protocol over $P$, the protocol can proceed from any configuration $c$ in $r$ to interact with members in $P^{\prime}\setminus P$, namely produce a run that is not possible only over $P$.
-3.
-grassroots if it is oblivious and interactive.
-Note that the definition is stated in terms of any two populations $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$.
-Being oblivious, in the context of Grassroots Consensus, means that if the prevailing populations of the protocol during a run are always included in $P$, then the presence of additional agents hanging around, those in $P^{\prime}\setminus P$, which do not participate in the protocol, should not interfere with the protocol or affect its behaviour.
-Clearly, this is the case with Grassroots Consensus, as its prevailing populations may always choose not to amend the constitution to include members outside $P$.
-Hence Grassroots Consensus satisfies the requirement of being oblivious.
-Being interactive, in the context of Grassroots Consensus, means that if the prevailing populations of the protocol in some prefix of a run ending in a configuration $c$ are included in a set $P$, then the protocol has a computation starting from $c$ that interacts with members outside $P$ in a meaningful way, namely the protocol has a computation from $c$ that reaches a configuration $c^{\prime}$, in which members of $P^{\prime}\setminus P$ are included in the prevailing population, and thus $c^{\prime}$ cannot be reached if the prevailing populations are all restricted to $P$.
-Clearly, members of the prevailing population $P$, starting from configuration $c$, may amend the constitution to include in $P$ a new member $p^{\prime}\in(P^{\prime}\setminus P)$ and thus reach a configuration $c^{\prime}$ in which $p^{\prime}\notin P$ is a member.
-Hence Grassroots Consensus satisfies the requirement of being interactive.
-Together, these two properties imply that constitutional Grassroots Consensus is indeed grassroots.
-We thus conclude:
-Theorem 5.2.
-Grassroots Consensus is grassroots.
-Note that the two requirements cannot even be expressed in the context of the basic (non-constitutional) Grassroots Consensus protocol, as the population $P$ is given and fixed.
-Also note that the constitutional protocol being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
-6 Constitutional Grassroots Consensus
+5 Constitutional Consensus with Constitutional Amendments
 The problem of reconfiguration of a consensus protocol—namely changing the set of its participants—has engaged the distributed computing research community for decades [40, 42, 2, 40], ever since it was first formulated in Lamport’s Paxos paper [21].
-Here we explore the more general problem of constitutional amendment.
-Supporting constitutional amendments makes the protocol grassroots, as we formally prove in Section 5.
-6.1 The Constitution
-Generally, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself.
-The notion of a constitution and its amendment is different from the well-studied notion of reconfiguration in at least two respects:
+Here we explore the more general problem of constitutional amendments, explained and defined below.
+Supporting constitutional amendments makes the consensus protocol grassroots, as proven in Section 6.
+5.1 Constitutional Consensus: From General to Specific
+In its most general sense, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself.
+The notion of constitutional amendment is different from the well-studied notion of reconfiguration in at least two respects:
   1. 1.
-    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including of course the agents executing it but also the protocol itself and its parameters.
+    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including the agents executing it, but also the protocol itself and its parameters.
   2. 2.
-    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to be coming “out of the blue”.
-Amending the constitution of a protocol, on the other hand, is done by the agents executing the protocol themselves, in accordance with the prevailing constitution’s constitutional amendment rule.
-Here, we consider a restricted type of constitution $(P,\sigma,\Delta)$, with $\emptyset\subset P\subset\Pi$ being the population, $\frac{1}{2}\leq\sigma<1$ being a supermajority among the population the approval of which is needed to amend the constitution, which is also the supermajority used by the underlying consensus protocol, and $\Delta$ being the agreed-upon estimate of the prevailing population $P$ regarding the delay needed to ensure the liveness of the protocol.
-The three components of the constitution affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony.
-Even more so, $P$, $\sigma$, and $\Delta$ are key components of Grassroots Consensus, given that its primary aim is to execute digital social contracts, as explained next.
-The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc.
+    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to descend on obedient agents “out of the blue”.
+Amending the constitution of a protocol, on the other hand, is done by the volitional agents (e.g., people) executing the protocol, in accordance with the prevailing constitutional amendment rule.
+In this work we consider a restricted type of constitution $(P,\sigma,\Delta)$ (Definition 2.1), the components of which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony.
+Even more so, they are key to Constitutional Consensus, given that its primary aim is to execute digital social contracts [9, 38]:
+  1. 1.
+    $P$: The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc.
 Clearly, they should be sovereign to accept new members and remove existing members, namely to amend $P$.
-As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should $\sigma$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased.
-Adding new people to $P$, initially known to and trusted by only a subset of $P$, may require increasing $\sigma$, even if temporarily.
-The parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members are admitted with lower-capacity devices, $\Delta$ may need to be increased.
-In addition, technology advanced may allow decreasing $\Delta$.
-Thus, the prevailing population $P$ should be sovereign to amend the values of $\sigma$ and $\Delta$ as they see fit.
-6.2 Democratic Constitutional Amendment
-Here we describe meta-level democratic processes for deciding on amending each of the components of a constitution $(P,\sigma,\Delta)$.
+  2. 2.
+    $\sigma$: As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should the $\sigma$-supermajority among $P$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased.
+Adding new people to $P$, initially known to and trusted by only a fraction of $P$, may entail increasing $\sigma$, even if temporarily, in order to ensure that sufficiently many trusted agents are part of every decision.
+  3. 3.
+    $\Delta$: The delay parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members with lower-capacity devices are admitted, $\Delta$ may need to be increased.
+In addition, technology advances may allow decreasing $\Delta$.
+Thus, the prevailing population $P$ should be sovereign to amend the value $\Delta$ as they see fit.
+5.2 Democratic Constitutional Amendments
+The Constitutional Consensus protocol attempts to implement valid constitutional amendment decisions as soon as they are made, but is independent of how these decisions are made.
+Computational Social Choice research [6] has investigated democratic decision-making processes extensively.
+Here we describe possible democratic processes for amending each of the components of a constitution $(P,\sigma,\Delta)$.
 (1) Amending $P$.
-Regarding amending $P$ to $P^{\prime}$, requiring $\sigma$-supermajorities among both $P$ and $P^{\prime}$ to approve the amendment has been proposed before [17], to avoid, among other problems, the ‘dead sailors problem’ (in which all of $P^{\prime}$ are in fact dead, resulting in a constitutional stalemate) presented in the original Paxos paper [26].
-Once $P$ can be amended, the question of sybils (fake and duplicate identities) entering $P$, in addition to Byzantines, comes to the fore.
-First, since one Byzantine agent, having penetrated $P$ with multiple identities, can cause as much harm as multiple Byzantine agents.
-In fact, the principal goal of permissionless protocols, both Proof-of-Work [32] and Proof-of-Stake [25, 10], is to augment consensus with mechanisms that make it sybil-resilient.
-Second, since a person controlling multiple digital identities in $P$—even if all correct at the consensus level—will have undue influence on the meta-level democratic processes, violating one person – one vote.
-Existing large-scale digital communities suffer from extensive penetration of sybils.
-For example, Facebook removes around 1Bn fake accounts every quarter [48].
-Most work on sybil detection and elimination employs top-down approaches suitable for global platforms [50, 18, 37, 12, 4, 6], but by their nature do not provide support for the democratic formation and conduct of digital communities.
-For grassroots consensus, egalitarian sybil-repelling protocols for the grassroots formation of digital communities [33, 39] may be more appropriate.
-They are based on mutual trust among the people forming a community and their willingness to backup this trust with mutual sureties.
-These protocols ensure, under certain conditions, that a digital community can grow indefinitely while retaining a bounded fraction of sybils.
-A complementary line of work [38, 29], relevant to the meta-level democratic processes, shows how a community can employ sybil-resilient democratic governance: How the safety and liveness of digital democratic processes can be ensured despite a bounded penetration of sybils, even in the face of partial participation by its non-sybil members.
-Together, the two lines of work offer a comprehensive solution for (i) sybil-repellence during the egalitarian formation of digital communities and (ii) sybil-resilience in the democratic governance of digital communities, both needed during by the meta-level processes of constitutional amendment of Grassroots Consensus, as discussed next.
+For amending $P$ to $P^{\prime}$, we require that a $\sigma$-supermajority among $P$ and that all new members of $P^{\prime}$, namely those in $P^{\prime}\setminus P$, approve the decision.
+The purpose of the supermajority is to ensure that a majority of the correct agents in $P$ are willing to hand-over control to $P^{\prime}$.
+The purpose of requiring all new participants in $P^{\prime}$ to sign is, first, to respect their autonomy—an agent should agree to participate—and, second, to avoid the ‘dead sailors problem’ [21], in which all of $P^{\prime}$ are new but in fact dead, resulting in a constitutional stalemate.
+The respect for autonomy of new participants has a practical corollary: An honest person (who would participate as a correct agent) would not join a protocol instance if they are concerned that the prevailing population may have too many Byzantines, namely that the prevailing constitution is not safe.
+The flip side of this concern—that the prevailing population not add malicious agents—has been studied extensively, as discussed next.
+A constitutional consensus protocol shares a fundamental challenge with permissionless consensus protocols: Both allow the participating agent set to change dynamically, yet must prevent Sybil attacks where a single entity creates multiple digital agents to gain disproportionate influence.
+Permissionless protocols address this by requiring agents to prove control of scarce resources—computational power in Proof-of-Work [27] or financial stake in Proof-of-Stake [19, 8].
+We expect realization of constitutional consensus to take a people’s-based approach: To rely on trust relations among real people, and admit new agents only when existing members know, trust, and verify them as genuine individuals.
+This requires (i) sybil-resistant admission protocols [28, 33] during community formation to ensure that digital agents genuinely represent distinct people rather than fake or duplicate identities and (ii) sybil-resilient democratic governance [32, 24], to ensure that sybils cannot overpower the genuine agents in democratic decisions.
+These mechanisms are complementary to this work.
 (2) Amending $\sigma$.
-Amending the supermajority by which a constitution can be amended has been studies by Abramowitz et al.
+Amending the supermajority by which a constitution can be amended has been studied by Abramowitz et al.
 [1], who showed that the $h$-rule (similar to the $h$-index of author citation metrics) is the only rule that satisfies certain intuitive and self-evident axioms.
-According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$.
+According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$, in a process left unspecified.
 The value of $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which there is a $\sigma^{\prime}$-supermajority among the prevailing population $P$ who voted for a value greater or equal to $\sigma^{\prime}$.
@@
 Else $\sigma$ remains unchanged.
+Thus, to change $\sigma$ from $\frac{2}{3}$ to $\frac{3}{4}$, the approval of a $\frac{3}{4}$-supermajority among $P$ is needed, and to change to unanimity ($\sigma=\frac{2|P|-1}{2|P|}$), approval by all members of $P$ is needed.
+And if the prevailing constitution states unanimity, then changing back to a $\frac{2}{3}$-supermajority requires unanimous consent among $P$.
+The $h$-rule ensures that changing the supermajority threshold requires meeting that very threshold: To raise $\sigma$ to $\sigma^{\prime}$, you need a $\sigma^{\prime}$-supermajority supporting it (preventing minorities from imposing stricter requirements on the majority), while to lower $\sigma$ from the current value, you need the current $\sigma$-supermajority to agree (preventing minorities from weakening the constitution’s protection).
 (3) Amending $\Delta$.
 The value of the parameter $\Delta$ should be increased or decreased only if the majority of the correct agents wish so.
-To this end, we use the Suppress Outer-f parameter update rule by Shahaf et al.
-[38] [^1] In reference [38] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion., as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified.
-The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good”.
-If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median of the remaining votes $\Delta^{\prime}$ is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
-Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median of the remaining votes $\Delta^{\prime}$ is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
+To this end, we offer the Suppress Outer-f parameter update rule [^1] In reference [32] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion.
+by Shahaf et al.
+[32], as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified.
+The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good” [29].
+If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
+Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
 Else $\Delta$ remains unchanged.
@@
 If all correct agents unanimously wish to update $\Delta$ to $\Delta^{\prime}$ then $\Delta$ will change to $\Delta^{\prime}$ no matter what the Byzantine agents vote, provided $f<\frac{1}{3}n$.
-In particular, if the protocol is not making progress, it can be expected that all correct agents will vote to increase $\Delta$ and will succeed in doing so.
-6.3 Grassroots Consensus with Constitutional Amendment
-While we suggested specific meta-level democratic processes for amending the three components of the constitution, the safety and liveness of constitutional amendment does not depend on their specifics.
-[…]
-In a grassroots setup, the identifier of each instance of a grassroots consensus protocol should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents.
-This can be achieved, for example, by having the identifier Id of a protocol instance to consists of the founders $P_{1}$ together with the maximal serial number of any grassroots consensus protocol any founder has founded.
-This ensures that if $P_{1}$ has at least one correct agent, then Id would be unique.
-Definition 6.1 (Valid Constitutional Amendment).
-Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, $d_{i}$ is valid if $i=1$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid, and $Q_{i}$ is a $\max(\sigma_{i-1},\sigma_{i})$-supermajority in $P_{i-1}$ and a $\max(\sigma_{i-1},\sigma_{i})$-supermajority in $P_{i}$.
+In summary:
+A constitution $(P,\sigma,\Delta)$ could be amended democratically to $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ as follows: 1.
+$P$ to $P^{\prime}$ via the approval of both a $\sigma$-supermajority among $P$ and all members of $P^{\prime}\setminus P$.
+2.
+$\sigma$ to $\sigma^{\prime}$ via the h-rule for constitutional amendment [1].
+3.
+$\Delta$ to $\Delta^{\prime}$ via the suppress outer-f parameter update rule [32].
+5.3 Constitutional Amendments
+While we suggest above democratic decision-making processes for amending the three components of the constitution, their specifics are immaterial to the Constitutional Consensus protocol.
+The only thing that matters is the interface between the decision making process and the protocol, which takes the form of constitutional amendment decisions.
+The following definition assumes a Constitutional Consensus protocol instance identifier Id; we discuss its selection below.
+Definition 5.1 (Constitutional Amendment Decision).
+A constitutional amendment decision is a pair $d=(h,x)$, where $x=(\textsc{amend},\textit{Id},i,C,C^{\prime})$, Id is an instance identifier, $i\geq 1$ is the index of the decision within this instance, $C,C^{\prime}$ are constitutions, referred to as the old and new constitutions of $d$, respectively, and $h$ is a multisignature of the hash of $x$ by some $Q\subset\Pi$.
+If $i=1$ then $C=\bot$ and $P_{1}$ are referred to as the founders of the protocol instance Id—they are the signatories of its initial constitution and the participants of the first epoch.
+Note that the structure of $d$ is that of a constitutional genesis block.
+During a run of a protocol instance Id, agents produce a possibly-infinite ordered sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$ using democratic decision-making processes discussed above or others, where $d_{i}=(h_{i},x_{i})$ and $x_{i}=(\textsc{amend},\textit{Id},i,C_{i-1},C_{i})$.
+The protocol commences with $d_{1}$ as the constitutional genesis block of the first epoch, and with each subsequent $d_{i}$ serving as the constitutional genesis block of the $i^{th}$ epoch.
+Definition 5.2 (Valid Constitutional Amendment Decision and Block).
+Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, the amendment $d_{i}$ with prevailing constitution $C_{i-1}=(P_{i-1},\sigma_{i-1},\Delta_{i-1})$, new constitution $C_{i}=(P_{i},\sigma_{i},\Delta_{i})$, and signatories $Q_{i}$ is valid if $i=1$, $P_{0}=\emptyset$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid and the signatories $Q_{i}$ are (i) a $\sigma_{i-1}$-supermajority in $P_{i-1}$, (ii) a $\sigma_{i}$-supermajority in $P_{i}$, and (iii) $(P_{i}\setminus P_{i-1})\subseteq Q_{i}$.
 The sequence is valid if all its members are valid.
-Two non-identical constitutional amendment decisions are conflicting if they have the same index.
-Note that the two consecutive supermajorities are possibly overlapping, even identical.
-In the special case that only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed; if only $\sigma_{i}$ changes then a single $\max(\sigma_{i_{1}},\sigma_{i})$-supermajority among $P_{i}$ is needed; and if only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$ and a $\sigma_{i}$-supermajority among $P_{i}$ is needed.
-This definition ensures that the $h$-rule is followed and that the ‘dead sailors’ scenario of Paxos (that the new $P^{\prime}$ does not exists) [26] has not happened.
-We assume that correct agents do not sign invalid or two conflicting constitutional amendment decisions, and hence can make the following observation.
-{observation}[Safety of Constitutional Amendment] Given a valid sequence of constitutional amendment decisions $d_{1},\ldots d_{i}$, $i\geq 1$, if the number $f$ of faulty agents in $P_{i}$ among the total number of agents $n=|P_{i}|$ satisfies $\sigma_{i}\geq\frac{n+f}{2n}$, then the constitutional amendment process cannot extend this sequence with two conflicting valid constitutional amendment decisions.
-A constitutional amendment decision is enacted by the consensus protocol via a block that includes it as payload.
-To ensure orderly transition among constitutions, we assume that a correct agent would not sign $d_{i+1}$ before observing that the amendment decision $d_{i}$ has been finalized by the underlying consensus protocol, namely that a block containing $d_{i}$ as payload has been finalized (as we explain below, only leader blocks carry constitutional amendments as payload).
-We exploit this in the consensus protocol by assuming that a $d_{i+1}$ amendment decision cannot occur as a new payload of a correct agent $p$ before a block with $d_{i}$ is final in the blocklace of $p$.
-Definition 6.2 (Valid Constitutional Amendment Block).
-A block $b$ is a valid constitutional amendment block if it is a valid initial block with an initial constitution $d_{1}$ as payload, in which case it is called a founding constitution block, or with a valid constitution $d_{i}$ as payload, $i>1$, and for which there is a valid constitutional block $b^{\prime}$ with $d_{i-1}$ as payload and $b^{\prime}$ is final in $[b]$.
-Definition 6.3 (Constitutional, Founding, Prevailing, Pending).
-A blocklace $B$ is constitutional if it has a sole founding constitution block, with its constitution referred to as the founding constitution of $B$.
-Given a constitutional blocklace $B$, the prevailing constitution of $B$ is the constitution of the most-recent final valid constitutional amendment block in $B$.
-The sole pending constitution of $B$, if any, is the constitution of the most-recent non-final valid constitutional amendment block in $B$, if there is one.
-When these terms apply to a block $b$ they refer to the blocklace $[b]$.
-Importantly, we amend the definition of quiescence (Definition 2.2) to state that a blocklace with a pending constitution is not quiescent.
-Beyond that, Definitions 2.1, 2.2, and 2.3 are kept verbatim, with the prevailing constitution of each block interpreted as in Definition 6.3, and the $\sigma$-supermajority required for rounds to advance (in Definition 2.3) counted out of the prevailing constitution of the agent’s entire blockchain.
-The Grassroots Consensus protocol remains the same except in the way it assigns payloads to blocks, which is changed to facilitate constitutional amendments and avoid conflicting decisions during the transition between constitutions.
-This change is effected without modifying the Grassroots Consensus protocol (Algorithm 1), but through constraining the payload variable while there are pending amendments, as described in Algorithm 3 – an agent that knows of a constitutional amendment $d_{i}$ that is not prevailing does not send the usual payload.
-If the agent is a round leader sending a leader block, it sets its payload to $d_{i}$, and otherwise it sets it to $\bot$.
-Constitutional Amendment: If $p$ knows of a constitutional amendment decision $d$ not prevailing in $B$ then prior to issuing a block of round $r+1$, if $p$ is the leader of round $r+1$ then $\emph{payload}\leftarrow d$ else $\emph{payload}\leftarrow\bot$ (retaining the outstanding value of payload, if not $\bot$, for the future).
-Algorithm 3 Grassroots Consensus with Constitutional Amendment.
-The word knows in the Constitutional Amendment rule has a double meaning: Either $p$ knows that the meta-level process of constitutional amendment has produced a new valid constitutional amendment decision $d$ not yet prevailing in $B$, or $p$ knows that its blocklace $B$ includes a valid pending constitutional amendment block with payload $d$.
-To ensure liveness, an correct agent $p$ that is removed by a constitution amendment $d_{i}$ (namely, $p$ is in $d_{i-1}$ and not in $d_{i}$) must continue to participate in the protocol until it observes a final leader block $b$ whose prevailing constitution is $d_{i}$.
-In particular, the constitution amendment $d_{i}$ is final in $[b]$.
-Excluding faulty agents.
-One natural application of constitutional amendment is excluding exposed faulty agents, in particular equivocators or agents that produce invalid blocks.
-The protocol is amended so that a correct agent $p$ does not sign a constitutional amendment $i$ if $P_{i}$ includes an agent exposed as faulty in $p$’s blocklace $B$.
-As a block known to a correct agent is eventually known to every correct agent, every agent exposed as faulty by one correct agent will eventually be exposed as such by every correct agent, and all exposed faulty agents will eventually be excluded from the prevailing population.
+Two non-identical constitutional amendment decisions are conflicting if they have the same instance identifiers and the same index.
+Given a sequence of valid constitutional amendment decisions, a block is constitutional if it includes a valid constitutional amendment decision as payload.
+Note that for agents to determine that a constitutional amendment decision is valid, they need access to all its predecessors.
+Also note that any two consecutive supermajorities may overlap, even be identical.
+In particular, if only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed.
+If only $\sigma_{i}$ changes then a single $\max(\sigma_{i-1},\sigma_{i})$-supermajority among $P_{i}$ is needed; this ensures that the $h$-rule [1] is followed.
+If only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$, a $\sigma_{i}$-supermajority among $P_{i}$, and all new members in $P_{i-1}$, are needed.
+We assume that correct agents do not sign invalid or two conflicting constitutional amendments, and hence can make the following observation.
+{observation}
+[Safety of Constitutional Amendments] Given a valid finite sequence of constitutional amendment decisions $D=d_{1},\ldots d_{i}$, $i\geq 1$, if every $C_{i}$ in $D$ is safe, then the constitutional amendment process cannot extend $D$ with two conflicting valid constitutional amendments.
+Proof 5.3.
+By a counting argument, for this to happen a correct agent would have to sign two conflicting constitutional amendments, a contradiction.
+Note that while the safety depends only on $\sigma_{i}$ and $P_{i}$, validity depends also on $\sigma_{i-1}$ and $P_{i-1}$, in order to adhere to the $h$-rule.
+Grassroots selection of protocol instance identifier.
+Each constitutional amendment decision is associated with an instance Id of the Constitutional Consensus protocol.
+In a grassroots setup, the identifier of each instance should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents, and should be chosen without central authority, central resource, or coordination among disjoint instances.
+This can be achieved, for example, by having the identifier Id of a protocol instance consist of (the hash of) the founders (initial population) of the instance, each paired with the maximal serial number of any Constitutional Consensus protocol instance they have founded.
+This ensures Id to be unique if the founders include at least one correct agent, who would not sign the initial constitutional decision if its Id is not formed correctly.
+Epochs and agent responsiveness.
+The standard assumption in consensus protocols is that all correct agents are responsive throughout the run.
+However, this assumption is unreasonable in a protocol like Constitutional Consensus in which the set of participating agents changes dynamically, especially since the set $\Pi$ of potential participants is infinite.
+Hence, we only assume that a correct agent $p$ is responsive only during an epoch in which it participates.
+The one exception is the case of agents that are behind, still in “older” epochs.
+To complete the epoch they may need to receive such blocks from former correct participants in that epoch.
+Hence, agent may still need to respond to nack-blocks pertaining to blocks they have issued in past epoch.
+Still, an agent may determine that en epoch is ‘closed’ (and no further nack-blocks need to be served from it) if it has evidence that all participants in the epoch have ended it.
+In particular, an epoch can be ‘closed’ when all its participants have coronated the constitutional amendment decision that ended (see End Epoch below).
+Note that the interaction in which a non-responsive agent $p$ needs to be ‘woken up’ in order to sign a constitutional amendment in which $p$ is a new participant happens outside the Constitutional Consensus protocol.
+In the practical settings we envision this would happen through social interactions (in-person or digital) between existing participants and prospective new participants.
+5.4 Epoch Transitions via Constitutional Amendments
+Protocol concepts.
+Every epoch commences with a constitutional genesis block that specifies its prevailing constitution.
+If, during the epoch, members of the prevailing population become aware of a valid decision $d$ to amend the prevailing constitution to a new constitution, they cease normal operation and only endeavour to finalize this decision.
+Once an agent has a wave $W$ that finalizes a constitutional block with $d$, it coronates $d$ by producing a coronate block with payload $(\textsc{coronate},d)$, sending it to both the old and new populations of $d$, and ceasing participation in the epoch.
+Members of the old population that receive it and have yet to coronate $d$ can nack any missing blocks to obtain $W$ and coronate $d$ themselves.
+A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block.
+An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own (implying that $d$ is final in its blocklace).
+The protocol.
+The protocol produces blocks also with constitutional amendment decisions as payload; such a block is referred to as a constitutional block.
+When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution.
+Subsequent ordered blocks with the same amendment are suppressed from the output.
+The protocol is extended in Algorithm 2 as follows:
+  1. 1.
+    Amendment Finalization occurs whenever the agent $p$ knows of a valid constitutional amendment $d$ that has not been finalized, upon which it continues following the protocol but only producing constitutional blocks with $d$ as a payload, until $d$ is finalized.
+Here, knows has a dual meaning: Either $p$ knows that the external democratic decision making process resulted in a new valid constitutional amendment $d$, or $p$ has a constitutional block with $d$.
+In either case, no constitutional block with $d$ has been finalized yet in $B$.
+  2. 2.
+    End Epoch applies when a constitutional block with $d$ is finalised, upon which members of the old population create a coronate block with $d$, send it to all members of the old and new populations, and cease participation in the old epoch.
+Coronation also plays a role in the democratic decision process: A correct agent would not sign a new constitutional amendment decision before coronating the last constitutional amendment it was a signatory to.
+  3. 3.
+    Start Epoch applies when any agent $p$ in the new population receives from the old population a supermajority of coronate block with decision $d$ (and provided $p$ issued a coronate block if it is also a member of the old population), upon which $p$ creates a constitutional genesis block $b$ with $d$ and updates its epoch identifier to be $b$.
+  4. 4.
+    Respond to Belated nack applies when any agent $p$ receives a nack $q$-block $b$ for a $p$-block of a past epoch $e$, upon which $p$ responds to the nack to help $q$ complete this past epoch.
+Note that if $p$ had previously received a coronate $q$-block ending $e$, then $p$ knows that $q$ is not missing any epoch $e$ blocks; in this case “judicially send…” would be a no-op.
+Moreover, an agent may ‘write off’ an epoch $e$ as complete after receiving from all participants in $e$ blocks coronating the end of $e$.
+Amendment Finalization by $p$: Upon issuing an ordinary block, if $p$ knows of a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload.
+End Epoch by $p$: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to members of $P\cup P^{\prime}$, and cease participation in the epoch.
+Start Epoch by $q$: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
+Upon receipt by $q\in P^{\prime}$ of a $\sigma$-supermajority among $P$ of $(\textsc{coronate},d)$ blocks, and provided $q$ previously issued such a block if $q\in P$, commence participation in the new epoch by $e\leftarrow d$ as the new epoch constitutional genesis block, $B\leftarrow B\cup\{d\}$, and retaining $D$ and payload.
+Respond to Belated nack by $q$: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
+Algorithm 2 Incorporating Constitutional Amendments in Constitutional Consensus.
+Code for any participant $p$ in the prevailing constitution and for any $q\in\Pi$.
+In the good case, a new constitutional amendment $d$ is issued by the leader of a wave and finalized within that same wave.
+In the worst case, it might take arbitrarily many waves to finalize $d$ prior to GST or due to faulty leaders.
+After GST, a constitutional amendment block issued by a correct leader among $P$ is bound to be finalized (and such a leader is bound to be selected eventually due to the fairness of $\textit{leader}_{P}$).
+Once finalized, correct members of $P$ will issue a coronate block and correct members of $P^{\prime}$ will receive it, allowing the current epoch to end and the next one to begin.
+This argument informally establishes the liveness of the transition among constitutions due to constitutional amendments.
+5.5 Safety and Liveness of Constitutional Amendments
+Theorem 5.4 (Constitutional Amendment Correctness).
+The Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) is correct.
+We need to prove that in every run of the protocol, there exists a sequence $T$ s.t.
+the two safety properties and three liveness properties of Definition 2.4 hold w.r.t.
+$T$.
+Note that the $i^{th}$ epoch $e_{i}$ is in fact $d_{i}$.
+We have proven that Algorithm 1 is correct in each epoch.
+Therefore, there is a sequence $T^{i}$ so that the output sequences produced by all participants in epoch $e_{i}$ are consistent and valid w.r.t.
+$T^{i}$.
+Let $T$ be the concatenation of all these sequences, $T:=T^{1}\cdot T^{2}\cdots$. We next prove that the protocol satisfies safety and liveness properties w.r.t. $T$. The theorem will follow.
+Lemma 5.5 (Constitutional Amendment Safety).
+Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies safety w.r.t. $T$ defined above.
+Proof 5.6.
+Validity immediately follows from the fact that agents participate only in epochs of which they are members.
+Consistency follows from the fact that agents participate in epochs sequentially, namely, they stop running the protocol in epoch $i$ before starting epoch $j$.
+Lemma 5.7 (Constitutional Amendment Liveness).
+Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies liveness w.r.t. $T$ defined above.
+Proof 5.8.
+Consider an epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$, and the two liveness requirements.
+If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem 4.10.
+Else assume $e$ is not the last epoch, and therefore it must have ended with a coronation by a $\sigma$-supermajority among $P$, all observing a wave finalizing a constitutional block $b$ with payload $d$.
+Let $p\in P$ be a correct agent.
+We argue that $p$ must eventually coronate $d$, and while doing so, call $\tau(b)$.
+As all correct participants in the epoch $e$ end it by calling $\tau(b)$, all would output the same blocks for $e$.
+If $p$ is among the $\sigma$-supermajority coronating $d$ we are done.
+Else there is at least one correct agent $q\in P$ that coronated $d$ and sent to $p$ a coronate block $b^{\prime}$ observing a wave $W$ that finalizes $b$ .
+Since $q$ and $p$ are correct, eventually a $\textsc{nack}(b^{\prime})$-block by $q$ will be received by $p$, who will respond (even after the epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $q$ any blocks in $W$ and their predecessors that $q$ is missing, upon receipt of which $p$ would issue a $(\textsc{coronate},d)$ block and output $\tau(b)$.
+Considering Constitution Amendment Liveness, once a constitutional amendment $d$ is known, all correct participants follow the high-throughput protocol and produce only blocks with payload $d$.
+Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem 4.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks.
+6 Constitutional Consensus is Grassroots
+We recall the formal definition of a grassroots protocol from reference [38], instantiate it informally (bypassing the formalities of the definition in [38]) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
+Definition 6.1 (Oblivious, Interactive, Grassroots).
+A protocol $\mathcal{F}$ is:
+  1. 1.
+    oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, a run of the protocol over $P$ can proceed just the same in the presence of agents in $P^{\prime}$.
+  2. 2.
+    interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every run of the protocol over $P$, the protocol can proceed from any configuration $c$ in $r$ to interact with members in $P^{\prime}\setminus P$, namely produce a run that is not possible only over $P$.
+  3. 3.
+    grassroots if it is oblivious and interactive.
+Note that the definition is stated in terms of any two populations $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$.
+Being oblivious, in the context of Constitutional Consensus, means that if the prevailing populations of the protocol during a run are always included in $P$, then the presence of additional agents hanging around, those in $P^{\prime}\setminus P$, which do not participate in the protocol, should not interfere with the protocol or affect its behaviour.
+Clearly, this is the case with Constitutional Consensus, as its prevailing populations may always choose not to amend the constitution to include members outside $P$.
+Hence Constitutional Consensus satisfies the requirement of being oblivious.
+Being interactive, in the context of Constitutional Consensus, means that if the prevailing populations of the protocol in some prefix of a run ending in a configuration $c$ are included in a set $P$, then the protocol has a computation starting from $c$ that interacts with members outside $P$ in a meaningful way, namely the protocol has a computation from $c$ that reaches a configuration $c^{\prime}$, in which members of $P^{\prime}\setminus P$ are included in the prevailing population, and thus $c^{\prime}$ cannot be reached if the prevailing populations are all restricted to $P$.
+Clearly, members of the prevailing population $P$, starting from configuration $c$, may amend the constitution to include in $P$ a new member $p^{\prime}\in(P^{\prime}\setminus P)$ and thus reach a configuration $c^{\prime}$ in which $p^{\prime}\notin P$ is a member.
+Hence Constitutional Consensus satisfies the requirement of being interactive.
+Together, these two properties imply that Constitutional Consensus is indeed grassroots.
+We thus conclude:
+Theorem 6.2.
+Constitutional Consensus is grassroots.
+Note that the two requirements cannot even be expressed without in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed.
+Also note that Constitutional Consensus being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
 7 Conclusions
-The Grassroots Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts.
+The Constitutional Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts.
 Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve [39].
 Even small community banks and cooperatives [37] may still be, at least initially, on the low-throughput side.
-As communities federate to form larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow.
+As communities federate to form ever-larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow.
 Until then, simulating the various scenarios would be productive future work.

### v5

@@
-Constitutional Consensus
-Abstract
-Consider people with smartphones operating without external authorities or global resources other than the network itself.
-In this setting, high-end applications supporting sovereign democratic digital communities, community banks, and digital cooperatives require consensus executed by community members, which must be reconfigurable to support community dynamics.
-The Constitutional Consensus protocol aims to address this need by introducing constitutional self-governance to consensus: participants dynamically amend the participant set, supermajority threshold, and timeout parameter through the consensus protocol itself.
-We achieve this by enhancing a DAG-based protocol (like Cordial Miners) with participant-controlled reconfiguration, while also supporting both high- and low-throughput operation (like Morpheus), remaining quiescent when idle.
-This three-way synthesis uniquely combines: (1) constitutional amendments for self-governance, (2) a cryptographic DAG structure for simplicity, parallelism, and throughput, and (3) both high- and low-throughput operation.
-The protocol achieves consensus in $3\delta$, maintains $O(n)$ amortized communication complexity during high throughput, and seamlessly transitions between modes.
-The basic protocol (without constitutional amendments) realizes these features in 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
-keywords:
-Consensus Protocols, Grassroots Platforms, Blocklace, Digital Social Contracts, Democratic Digital Communities
-1 Introduction
-Participant-Controlled Consensus.
-Consider people who wish to form self-governing digital communities without relying on servers or external authorities.
-These participants—whether members of community banks, digital cooperatives, or democratic assemblies—need consensus protocols they can execute on their smartphones while maintaining complete control over membership and governance rules.
-Unlike traditional reconfiguration where external authorities determine changes [40, 2], these communities require constitutional self-governance: participants must collectively control who joins, how decisions are made, and how the protocol operates, all through the consensus protocol itself.
-This paper presents Constitutional Consensus, a protocol that enables such self-governance by allowing participants to dynamically amend the participant set $P$, supermajority threshold $\sigma$, and timeout parameter $\Delta$ through democratic decisions implemented within the protocol.
+Constitutional Consensus for Democratic Governance
+Abstract.
+Permissionless-consensus-based Decentralised Autonomous Organisations (DAOs) are the prevailing paradigm for participant-governed digital organisations.
+As participants have verified resources but no trusted identities, this ecosystem is necessarily plutocratic (one coin – one vote).
+Here we offer, for the first time, a democratic (one person – one vote) paradigm for the governance of digital communities and organisations, based on permissioned consensus and egalitarian decision processes.
+In line with Lamport’s vision of consensus as a self-governing parliament, in the democratic paradigm a constitution specifies both a decision making protocol as well as a consensus protocol, combined to let participants amend the constitution through constitutionally-valid decisions that are ratified by consensus.
+To meaningfully instantiate this paradigm we integrate the disciplines of distributed computing and computational social choice, with the goal of providing a practical and efficient smartphone-based solution for the democratic self-governance of grassroots sovereign digital communities and organisations.
+The resulting Constitutional Consensus protocol employs (1) state-of-the-art Sybil-resilient democratic decision processes for amending the set of participants, supermajority threshold, and timeout; and (2) a novel Byzantine-fault tolerant consensus protocol that is DAG-based (following Cordial Miners) thus eschewing reliable broadcast, with dual-mode operation (following Morpheus) that is quiescent when idle, has spontaneous leaders for isolated transactions, and formal round-robin leadership during high throughput.
+Constitutional Consensus operates in epochs, each governed by the prevailing constitution, with the next epoch commencing via the ratification of a constitutionally-valid constitutional-amendment decision.
+The protocol achieves consensus in $3\delta$ with $O(n)$ amortized communication complexity during high throughput.
+The single-epoch protocol is quite simple, requiring only 25 lines of pseudocode; the multi-epoch protocol is grassroots, as each member can initiate and participate in multiple independent instances that can form, operate and grow concurrently without coordination or dependence on global resources other than the net.
+Consensus, Paxos, constitutional amendment, digital social contracts, democratic digital communities, grassroots
+$c$s: Computer systems organization Peer-to-peer architecturesccs: Software and its engineering Distributed systems organizing principles
+1. Introduction
+Participant-governance of digital organisations.
+Permissionless-consensus-based Decentralised Autonomous Organisations (DAOs) (Ethereum, ndao; Ding et al., 2023; Hassan and De Filippi, 2021; Talmon, 2023) are the prevailing paradigm for realising participant-governed digital organisations.
+As participants have verified resources (e.g., staked tokens) but no trusted identities, this ecosystem is necessarily plutocratic: voting power is proportional to token holdings (one coin – one vote).
+Here we offer, for the first time, a democratic (one person – one vote) paradigm for the governance of digital communities and organisations, based on permissioned consensus and egalitarian decision processes.
+Constitutional self-governance.
+Reconfiguration—changing the set of participants in a consensus protocol—has been studied extensively (Lamport, 1998, 2001; Lamport et al., 2009; Aguilera et al., 2011; Spiegelman et al., 2017; Ongaro and Ousterhout, 2014).
+This literature addresses the mechanism of how membership changes execute safely, but not who decides on them: the decision authority is either left unspecified or delegated to an external entity.
+Moreover, prior work applies reconfiguration only to membership, leaving the rest of the protocol fixed.
+[^1] Tezos (Goodman, 2014) has no notion of membership but can self-modify the protocol via plutocratic governance.
+In line with Lamport’s vision of consensus as a self-governing parliament (Lamport, 1998), we propose a democratic paradigm in which a constitution specifies both a decision-making protocol and a consensus protocol, combined to let participants amend the constitution through constitutionally-valid decisions that are ratified by consensus.
+To meaningfully instantiate this paradigm we integrate the disciplines of distributed computing (Lynch, 1996) and computational social choice (Brandt et al., 2016), with the goal of providing a practical and efficient smartphone-based solution for the democratic self-governance of grassroots (Shapiro, 2023a) sovereign digital communities and organisations.
+The resulting Constitutional Consensus protocol presented here employs (1) state-of-the-art Sybil-resilient (Shahaf et al., 2020) democratic decision processes for amending the set of participants (Bulteau et al., 2021), supermajority threshold (Abramowitz et al., 2021a), and timeout (Shahaf et al., 2019); and (2) a novel Byzantine-fault tolerant consensus protocol that is DAG-based (following Cordial Miners (Keidar et al., 2023)) thus eschewing reliable broadcast, with dual-mode operation (following Morpheus (Lewis-Pye and Shapiro, 2025)) that is quiescent when idle, has spontaneous leaders for isolated transactions, and formal round-robin leadership during high throughput.
 Digital social contracts.
-These self-governing communities implement digital social contracts [9]—deterministic state machines that encode agreements among people who know each other.
-Like smart contracts, they provide computational guarantees about contract execution.
-However, while smart contracts [10] operate among anonymous accounts and run on consensus protocols executed by third-party servers, social contracts are among identified people and run on consensus protocols executed by the participants’ own devices.
-Existing consensus-based smart contracts, notably in Decentralized Finance (DeFi) platforms [10, 7, 25], have driven development of high-performance consensus protocols [43, 17, 18, 4, 16], but these operate among pseudonymous accounts on global networks rather than identified people in local communities.
-This fundamental difference enables democratic rather than plutocratic governance, as shown in Table 1.
-Table 1: Comparing Smart Contracts and Consensus-Based Digital Social Contracts
-Smart Contract [10] Digital Social Contract [9]
-Among: Accounts People
-Executed by: Third-party servers The people’s smartphones
-In return for: Gas Naught
-Applications: Decentralized Finance (DeFi) [31, 15],
-Decentralized Autonomous Organizations (DAOs) [12, 11],
-Non-Fungible Tokens (NFTs) [41]
-Grassroots democratic digital communities [28, 33] and their federation [38, 39],
-community banks and digital cooperatives [37]
-Amended via: Self-modifying code Constitutional amendments
-Governance: Token-weighted voting (plutocracy) One person, one vote (democracy)
-Amendment Authority: External governance tokens or DAOs Participants themselves via supermajority
-Parameter Control: Fixed in code or controlled by token holders Democratically adjustable by participants
-We note that while smart contracts are necessarily consensus-based, digital social contracts are not.
-Some grassroots platforms, notably grassroots social networks [35] and grassroots cryptocurrencies [37, 22], do not require consensus on the total ordering of the acts of their participants: Their safety depends only on the agents obeying the social contract [38].
-This can be ensured by mechanisms weaker then consensus, e.g.
-mutual attestations among agents operating within Trusted Execution Environments (TEEs) [30].
-However, applications requiring collective agreement—such as community banks with multiple signatories or democratic governance with voting—do require consensus to ensure all participants agree on the order of transactions and decisions.
-Technical Requirements.
-Communities implementing digital social contracts face unique technical challenges.
-Small communities may have transaction rates much lower than network latency, requiring a protocol that remains quiescent—sending no messages when idle.
-Yet the same protocol must efficiently handle high-throughput periods when the community is larger and/or more active.
-During low throughput, transactions arrive one-at-a-time with no conflicts to resolve, so ideally any participant should be able to finalize their transaction without designated leaders.
-During high throughput, the protocol should match the performance of modern consensus protocols.
-Specifically, an effective protocol for self-governing communities requires:
-  1. 1.
-    Quiescent operation: No activity when idle, spontaneous leader emergence for isolated transactions
-  2. 2.
-    Dual-mode efficiency: Seamless transition between low and high throughput modes
-  3. 3.
-    Responsiveness: Network-speed operation unaffected by conservative timeout estimates
-  4. 4.
-    DAG-based structure: Leveraging blocklace [3] parallelism and integration
-  5. 5.
-    Constitutional amendments: Participant-controlled governance of protocol parameters
-Technical Foundation.
-Constitutional Consensus achieves these requirements through a three-way synthesis.
-From Morpheus [23], it inherits dual-mode operation with spontaneous leaders during low throughput and formal leadership during high throughput.
-From Cordial Miners [18], it adopts the blocklace structure for parallelism and efficiency.
-To these, it adds constitutional amendment mechanisms enabling true self-governance.
-The protocol operates in waves of three rounds.
-During low throughput, any participant with a transaction can spontaneously lead a wave, achieving finality in $3\delta$.
-During high throughput, formal leaders are selected round-robin, maintaining $O(n)$ amortized communication complexity.
-The basic protocol (without amendments) requires only 25 lines of pseudocode, making it one of the most concise consensus protocols for eventual synchrony.
-Constitutional Self-Governance.
-The key innovation lies in how participants control their own governance.
-Unlike smart contracts that rely on external governance tokens or DAOs [12], constitutional amendments are decided by the participants themselves through supermajority vote.
-While Lamport’s original Paxos envisioned a self-governing parliament [21], subsequent reconfiguration protocols [42, 2, 40, 20] assume external authorities determine membership changes.
-Constitutional Consensus returns to Paxos’s original vision of true self-governance.
-Unlike Proof-of-Stake systems where validators must stake capital [8], participation requires only the democratic approval of existing members.
-This creates a fundamental architectural difference: while blockchain systems typically support only a single global instance whose security depends on token value, Constitutional Consensus enables unlimited independent communities to self-govern simultaneously.
-Connection to Grassroots Architecture.
-This capability for multiple independent instances operating without central coordination or global resources makes Constitutional Consensus inherently grassroots [38].
-A community in Bangladesh, a cooperative in Brazil, and a federation in Berlin can each run their own instance, potentially with overlapping members, without any coordination or competition for resources.
-This architecture—multiple local instances rather than a single global instance—enables the broader grassroots vision of egalitarian digital platforms that operate without central authorities [36, 34].
-Within this grassroots architecture, Constitutional Consensus provides the consensus layer for applications requiring collective agreement, complementing consensus-free grassroots platforms like social networks [35] and cryptocurrencies [37, 22].
-Together, these protocols enable communities to build sovereign democratic digital infrastructure from the ground up, controlled entirely by their members rather than external authorities or token holders.
-This distinguishes Constitutional Consensus from global platforms—whether centralized like Facebook or decentralized like Bitcoin [26] and IPFS [5]—that can support only a single instance.
-Example: Grassroots Communities, Banks, and Federations.
-As a concrete scenario for grassroots platforms, consider a small town whose residents wish to form a digital community.
-They begin by deploying independent instances of consensus-free grassroots platforms on their smartphones—creating their own grassroots social network [35] and trading via self-minted grassroots cryptocurrencies [37, 22].
-Crucially, they do this without coordination with any global network or central authority.
-When their needs evolve to require collective agreement—community banks with multiple signatories, constitutional governance, or federation with neighbouring towns—they deploy their own Constitutional Consensus instance.
-This instance operates independently yet can interoperate with other communities’ instances, embodying the grassroots principle of multiple local instances rather than a single global system.
-Through Grassroots Federation [39], these autonomous instances can scale democratically from local to potentially global governance.
-Small communities ($\approx 100$ members) govern themselves directly; larger ones use sortition-selected assemblies [14, 39].
-Each community’s governance runs as a digital social contract on their own Constitutional Consensus instance, executed by members on their smartphones—no external servers or authorities required.
-Model.
-As deterministic fault-tolerant consensus is not solvable in asynchronous systems [13], we assume an eventual synchrony model, where communication delays are at most $\delta$ after some Global Stabilization Time (GST) that is unknown to the agents.
-We consider Byzantine failures, as some malicious actors may infiltrate the network, but we assume a bounded ratio of Byzantine agents (to be specified later), since, by and large, the agents trust each other.
+A digital social contract (Cardelli et al., 2020) is a voluntary agreement among a set of people, specified, fulfilled, and enforced digitally.
+Many digital social contracts — including social graphs (Shapiro, 2026), social networks (Shapiro, 2023b), and digital currencies (Shapiro, 2024; Lewis-Pye et al., 2023)—require no consensus at all.
+Those that do require consensus among the contract’s participants—notably democratic governance and federation of sovereign communities (Shapiro and Talmon, 2025)—may execute the protocol on the participants’ own devices.
+Constitutional Consensus is designed for such social contracts.
+Unlike smart contracts (De Filippi et al., 2021), whose execution is enforced by global consensus among third-party validators on a blockchain, digital social contracts are among people with mutually-known identities and are enforced by the participants themselves.
+This distinction—summarised in Table 1 in Appendix J—has a direct architectural consequence: governance is democratic rather than plutocratic.
+Technical approach.
+Constitutional Consensus operates in epochs, each governed by the prevailing constitution, with the next epoch commencing via the ratification of a constitutionally-valid constitutional-amendment decision.
+The single-epoch protocol is quite simple, requiring only 25 lines of pseudocode.
+After GST, a leader block is finalised within $3\delta$, and the amortized communication complexity per transaction is $O(n)$ during high throughput.
+Grassroots.
+Because Constitutional Consensus requires no external authority, global ledger, or staking globally-controlled capital, the multi-epoch protocol is grassroots (Shapiro, 2023a, 2026): each member can initiate and participate in multiple independent instances that can form, operate and grow concurrently without coordination or dependence on global resources other than the net.
+Any group of people can bootstrap their own instance on their smartphones.
+This is in contrast with Proof-of-Stake protocols like Ethereum 2.0 (Buterin et al., 2020), where security depends on the staked token having real-world value, creating a winner-take-all dynamic that precludes independent instances.
 Paper structure.
 Section 2 provides the model and problem definitions.
-Section 3 presents a single-epoch Constitutional Consensus protocol, with only an initial constitution and no constitutional amendments.
-Section 4 proves the safety and liveness of single-epoch Constitutional Consensus.
-Section 5 adds support for constitutional amendments – for each of the three parameters $P,\sigma$, and $\Delta$, we present a democratic process for amending it.
-We then present the protocol extension needed to realize constitutional amendments and prove its safety and liveness.
-Section 6 proves that Constitutional Consensus is grassroots.
+Section 3 presents single-epoch Constitutional Consensus, with no constitutional amendments.
+Section 4 presents Sybil-resilient democratic decision processes for amending the set of participants, supermajority threshold, and timeout.
+Section 5 presents the formal amendment mechanism and epoch transitions.
+Section 6 summarises related work, with a full survey in Appendix J.
 Section 7 concludes.
-2 Model and Problem
-Model.
-We assume a set of agents $\Pi$, each endowed with a unique key-pair and identified by its public key $p\in\Pi$.
+Correctness proofs and other technicalities are deferred to appendices.
+2. Model and Problem
+We assume a set of agents $\Pi$, each has a unique key-pair and identified by its public key $p\in\Pi$.
 While $\Pi$ is potentially-infinite, we refer only to finite subsets $P\subset\Pi$ of it.
@@
 An agent that does not follow the protocol is faulty or Byzantine.
-We assume an eventual synchrony (aka partial synchrony) model in which every message sent between correct agents is eventually received.
+We assume an eventual synchrony (partial synchrony) model in which every message sent between correct agents is eventually received (this abstracts a dissemination layer that can use acks, gossip, or network coding, and is orthogonal to the consensus protocol).
 Message latency is unbounded, but in every run there is a time called the Global Stabilization Time (GST), which is unknown to the agents, after which all messages arrive within a least upper bound $\delta$ time for some $\delta>0$.
-Note that the abstraction of reliable messaging does not mean that agents must communicate via point-to-point TCP connections.
-Rather, it abstracts away a dissemination layer that can be implemented using acks/nacks, gossip, network coding, etc., and is orthogonal to the consensus protocol in this paper.
-A protocol proceeds through a sequence of epochs, each commences with a new constitution (defined next), referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution.
-The constitution defines the set of participants in the epoch, as well as key protocol parameters:
-Definition 2.1 (Constitution, $\sigma$-Supermajority).
-A constitution is a triple $(P,\sigma,\Delta)$ where:
-  1. 1.
-    $P\subset\Pi$ is a set of agents, with $n:=|P|$,
-  2. 2.
-    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the agents $P$,
-  3. 3.
-    $\Delta>0$ is a presumed upper-bound on message delay after GST among $P$.
-We may omit $\Delta$ when its value is immaterial (e.g., when discussing protocol safety) and refer to the constitution $(P,\sigma)$.
-A $\sigma$-supermajority among $P$ is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
-Supermajorities and faulty agents.
-Typically, a permissioned consensus protocol assumes that there are at most $f<n$ faulty agents among $P$, and with this assumption we define $\sigma:=\frac{n+f}{2n}$.
-Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty agents in $P$.
-For example, for $f=0$, $\sigma=\frac{1}{2}$; for $f=\frac{1}{2}n$, $\sigma=\frac{3}{4}$; and if everyone but one agent is faulty then $\sigma=\frac{2n-1}{2n}=1-\frac{1}{2n}$, namely unanimity is required.
-For $n=3f+1$ (the standard assumption), $\sigma=\frac{n+\frac{n-1}{3}}{2n}=\frac{2}{3}-\frac{1}{6n}$, namely a supermajority greater-or-equal to $\frac{2}{3}n$ is required.
+In general, the constitution specifies the entire set of protocols employed—decision making processes as well as the consensus protocol for ratifying them.
+In our case, these consist of known democratic decision processes recalled in Section 4 and the novel Constitutional Consensus protocol, which is the main subject of this paper.
+In this work the constitution assumes both of these protocols to be fixed except for three amendable components, the participant set, the supermajority threshold, and the timeout parameter.
+A protocol proceeds through a sequence of epochs, each commences with a new constitution, referred to as the prevailing constitution during the epoch, and ends when the constitution is amended, namely replaced with a new constitution:
+Definition 2.1 (Amendable component of the Constitution).
+The amendable component of the constitution is the triple $(P,\sigma,\Delta)$ where:
+  1. (1)
+    $P\subset\Pi$ is a set of agents, referred to as its participants, with $n:=|P|$,
+  2. (2)
+    $\frac{1}{2}\leq\sigma<1$ is a fraction that specifies a supermajority among the participants,
+  3. (3)
+    $\Delta>0$ is a presumed upper-bound on message delay after GST among the participants.
+We may omit $\Delta$ when discussing protocol safety and refer to the constitution $(P,\sigma)$. A $\sigma$-supermajority among the participants is a fraction $Q\subseteq P$ such that $|Q|>\sigma n$.
+In the following we may use the constitution as a shorthand, e.g. “the prevailing constitution” instead of “the amendable component of the prevailing constitution”.
+Supermajorities, Sybils and Byzantines.
+Our protocols are defined using a supermajority that aims to overcome participants that are (1) Byzantine (arbitrarily-behaving) (Lamport et al., 1982), which hamper consensus and (2) Sybil (with fake or duplicate identity) (Shahaf et al., 2020), which hamper equality.
+Assuming there are at most $f<n$ such faulty participants, we define $\sigma:=\frac{n+f}{2n}$.
+Equivalently, $\sigma$ is chosen with the assumption that there are at most $f=(2\sigma-1)n$ faulty participants.
+Both protocols ensure safety with any $f<n$ and liveness if $f<\frac{1}{3}n$.
 Typically, permissioned consensus protocols wait to hear from “at least nfn-f agents among PP”, which in the standard case of $n=3f+1$ (but not otherwise) happens to be equivalent to “a $\sigma$-supermajority among PP”.
+While Byzantines and Sybils are related (e.g.
+Byzantines may aim to create and control multiple Sybils to amplify their ability to cause harm) they are not the same, and a realisation of Constitutional Consensus may choose to incorporate a tighter supermajority for each, say $\sigma_{B}$ for consensus and $\sigma_{S}$ for democratic decisions; for simplicity, we employ here one supermajority $\sigma$ for both.
 Constitution safety and liveness.
-A constitution thus defines the protocol’s assumptions about the environment: $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$ and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent; $\Delta\geq\delta$ is an upper bound on the communication time after GST.
-Definition 2.2 (Safe and Live Constitution).
-A constitution $(P,\sigma,\Delta)$ is safe if $P$ includes at most $f=(2\sigma-1)n$ faulty agents, and live if after GST any message sent among correct agents in $P$ arrives within $\Delta$.
+The (amendable component of the) constitution thus defines the protocol’s assumptions about the environment: A constitution $(P,\sigma,\Delta)$ is safe if $P$ is the set of agents running the protocol; there is a correct $\sigma$-supermajority in $P$, and every $\sigma$-supermajority in $P$ includes a majority of correct agents, hence the intersection of any two $\sigma$-supermajorities contains a correct agent.
+A constitution is live if $f<\frac{1}{3}n$ and after GST any message sent among correct participants arrives within $\Delta$.
+Constitution safety and liveness attest to the validity of the assumption of the constitution regarding the participants and the network connecting them.
+Note that $\Delta$ bounds the message delay among correct (i.e., participating and online) agents after GST, not the time until every agent comes online.
+The dual-mode design means the protocol is quiescent when there are no pending transactions, so agents need not be continuously available.
+When a transaction arrives, only a correct $\sigma$-supermajority must respond within $\Delta$; agents that are temporarily offline are accounted for by the choice of $\sigma$.
 Constitutional Consensus problem definition.
 A constitutional consensus protocol may have multiple instances.
-Each instance commences with an initial constitution $(P,\sigma,\Delta)$ that defines its participating agents $P$ and the key protocol parameters $\sigma$ and $\Delta$.
+Each instance commences with an initial constitution $(P,\sigma,\Delta)$.
 As in standard consensus, participants spontaneously issue transactions (these are the inputs to the protocol) and every agent outputs a sequence of transactions.
-In addition, constitutional consensus assumes an that participants engage in an unspecified democratic decision process that may result in a decision $d$ to amend the prevailing constitution.
-The discussion of such democratic processes is deferred to Section 5.
-For the purpose of the constitutional consensus protocol, all is needed is to assume: (i) the existence of such a process that may result in constitutional amendment decisions, which become known to the participants once made; (ii) that participants can verify a constitutional amendment decisions $d$ to be valid according to the rules specified in Section 5.
-A valid constitutional amendment decision $d$ defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$, in which the participants $P^{\prime}$ and/or the parameters $\sigma^{\prime}$ and $\Delta^{\prime}$ are the result of amending the prevailing constitution.
-In the following, constitution refers to the new constitution of a valid constitutional amendment decision, and an initial constitution is the new constitution of a valid amendment of a constitution with an empty population.
-Valid amendments are provided to the consensus protocol as inputs by participants that know of them.
-Thus, the output sequence of a constitutional consensus protocol includes both transactions and valid constitutional amendment decisions, as follows:
+In addition, Constitutional Consensus assumes that participants engage in decision processes as specified by the constitution, which may result in decisions to amend the prevailing constitution.
+As an example, Section 5 shows how known democratic decision processes can be applied in this new context.
+Note that the parameters $P$ and $\sigma$ are shared: they govern both the consensus protocol and the democratic decision processes themselves.
+The interface between the Constitutional Consensus protocol and the constitutional democratic decision processes is as follows: (i) a democratic process may produce constitutional amendment decisions, which become known to the participants once made; (ii) participants can verify that a constitutional amendment decision is valid according to the prevailing constitution, as specified in Section 5.
+A valid constitutional amendment decision defines a new constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$.
+An initial constitution is the new constitution of a valid amendment of a constitution with an empty population.
+Constitutional amendment decisions are provided to the consensus protocol as inputs by participants that know of them.
+Thus, the protocol’s output sequence includes both transactions and valid constitutional amendment decisions, as follows:
 Definition 2.2 (Constitutional sequence, epoch, epoch sequence, prevailing constitution, participants).
@@
 An agent participates in an epoch if it is a member of the population of its prevailing constitution.
-The constitutional output sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
+The constitutional output-sequences produced by correct agents executing a constitutional consensus protocol instance must satisfy the following safety and liveness conditions:
 Definition 2.3 (Constitutional Consensus Correctness).
-In any run of an instance of constitutional consensus in which all valid constitutional amendment decisions made are safe and live, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following safety and liveness properties hold:
-  1. 1.
-    Safety:
-    Consistency
+A constitutional consensus protocol is correct if, in any run of it in which all valid constitutional amendments are safe, there exists a constitutional sequence $T=t_{1},t_{2},\dots$ for which the following properties hold:
+Consistency:
 Every correct agent’s output sequence is a subsequence of $T$.
-    Validity
-    If a correct agent $p$ outputs a transaction $t$, then $p$ is a participant in $t$’s epoch in $T$.
-  2. 2.
-    Liveness:
-    Epoch Liveness
-    If a correct agent $p$ is a participant in an epoch $d$ in $T$, then $p$’s output sequence eventually includes all the transactions in $d$’s epoch sequence.
-Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in the epoch $d$ is included in $T$.
-    Constitution Amendment Liveness
+Validity:
+If a correct agent $p$ outputs $t\in T$, then $p$ is a participant in $t$’s epoch.
+Epoch Liveness:
+For a correct participant $p$ in an epoch $d$ in $T$, $p$’s output eventually includes all the transactions in $d$’s epoch sequence.
+Furthermore, if $d$ is the last epoch in $T$, then every transaction input by a correct participant in epoch $d$ is included in $T$.
+Constitution Amendment Liveness:
 If a valid constitutional amendment decision $d$ is provided as input to the protocol by a correct agent, then $T$ eventually includes $d$.
-Note that if there is no decision to amend the initial constitution then the protocol run has a single epoch with a fixed set of participants.
-In this case, the definition above coincides with the standard consensus correctness definition.
-Note also that progress is guaranteed (after GST) even if the constitution changes infinitely many times.
-However, normal transactions might starve due to frequent amendments.
-This definition allows the implementation to prioritize constitution changes over regular transactions, allowing the latter to starve.
-In practice, amendments are expected to be rare compared to transactions, so we believe this modus operandi to be reasonable.
-3 Single Epoch Constitutional Consensus
-Here we present the behaviour of the Constitutional Consensus protocol during a single epoch.
-The protocol is based on the blocklace data structure, which we define in Section 3.1.
-The protocol’s concepts are detailed in Section 3.2 and its pseudo-code in Section 3.3.
-3.1 Blocklace Preliminaries
+Note that with no amendments, this reduces to the standard consensus correctness definition.
+Progress is guaranteed after GST even under infinitely many amendments, though frequent amendments may starve normal transactions; in practice, amendments are expected to be rare.
+3. Single Epoch Constitutional Consensus
+Here we present the operation of the Constitutional Consensus protocol during a single epoch.
+We overview the protocol’s operation, deferring its formal description and correctness proofs to Appendices A and E, resp.
+We begin (Section 3.1) by explaining the finalisation process, which ensures the protocol’s safety.
+We then discuss the conditions for generating new protocol blocks so as to guarantee liveness (Section 3.2).
+Finally, Section 3.3 summarises the protocol’s state, initialisation, and output.
+3.1. Safety: Finalising Transactions
+Blocklace-based ordering.
+The protocol is based on a blocklace (Keidar et al., 2023), a cryptographic DAG data structure consisting of blocks, each of which includes a payload and hash-based pointers to other (earlier) blocks.
+A block is signed by its creator; a block created by agent $p$ is called a $p$-block.
+An initial block has no pointers.
+The DAG structure defines a partial order on blocks: we say that a block $b$ observes a block $b^{\prime}$ in a blocklace $B$ if there is a path from $b$ to $b^{\prime}$ in $B$.
+(In particular, every block observes itself).
+The depth of a block $b$, $\textit{depth}(b)$, is 0 if $b$ is initial else $\textit{depth}(b)=\textit{depth}(b^{\prime})+1$ where $b^{\prime}$ is a maximal-depth block pointed to by $b$.
+The closure of a block $b$, denoted $[b]$, is the set of blocks observed by $b$.
+The protocol employs an evolving shared blocklace, of which each participant $p\in P$ maintains a local copy $B$.
+In each epoch, the protocol creates a blocklace that begins with a constitutional genesis block $b$, whose payload specifies a new constitution $C=(P,\sigma,\Delta)$.
+Every participant commences the epoch with its copy of $B$ holding $\{b\}$.
+During the epoch, participants create new blocks that observe their latest blocklace $B$ and disseminate them to other participants.
+Blocks’ payloads may include transactions the agents wish to order and also serve as protocol messages to drive agreement.
+While a blocklace is only partially ordered, the protocol totally orders the blocks in an epoch (and the transactions included therein) in a consistent way, producing the same output sequence at all participants in an epoch.
+Moreover, epochs are also totally ordered, so that the totality of all blocks of all epochs are also totally ordered, even though not always by the same participants.
+Waves, modes, and leaders.
+The blocklace is partitioned into a sequence of waves.
+For $k>0$, the $k^{th}$ wave in $B$ is the set of blocks of depths 3k23k-2, 3k13k-1, and $3k$ in $B$, referred to as the the first, second, and third round of the wave, respectively.
+The $0^{th}$ wave of a blocklace is defined to be the singleton holding its constitutional genesis block.
+The protocol has two operation modes, high and low throughput.
+In high throughput mode, agents continuously generate blocks in all rounds.
+To break the symmetry among agents, we use a deterministic (and fair) leader function that defines a unique formal leader $p_{\ell}\in P$ in the first round of each wave.
+Within each wave, the protocol strives to finalise a first-round $p_{\ell}$-block (if it exists).
+In low throughput mode, agents are quiescent as long as they have no new transactions to send.
+The protocol only quiesces at the end of a wave, after all previously sent transactions have been ordered.
+The quiescent state can last for seconds, minutes, or even hours, as in small social communities, transactions might be infrequent.
+Once any participant has a transaction to order, it initiates the next wave by sending a first round block.
+All other agents join this wave by issuing second, and later third, round blocks.
+In this scenario, the single first round block in the wave is considered a spontaneous leader, and the protocol strives to finalise it.
+Finalising blocks in waves.
+Finalisation within a wave proceeds in three steps.
+The first round of the wave includes candidates for finality, the second includes endorsements of candidates, and the third includes ratifications, as we now explain.
+But first, we need to disregard equivocations, namely, two blocks by the same agent that do not observe each other.
+We say that a block $b$ approves a block $b^{\prime}$ if $b$ observes $b^{\prime}$ and does not observe any block equivocating with $b^{\prime}$.
+A second-round block $b$ endorses at most one first-round block in its wave: if the preceding wave is quiescent, $b$ endorses the sole first-round block it approves (if unique); if non-quiescent, $b$ endorses the formal leader’s block (if approved).
+A first-round block is ratified by a third-round block that approves a $\sigma$-supermajority of endorsing second-round blocks, and final once a $\sigma$-supermajority of third-round blocks ratify it.
+Since correct agents do not equivocate, at most one block per wave can be ratified or final.
+Round $r+1$ has a single leader block (red) approving a supermajority in round $r$ (the last round of the previous wave).
+Round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block.
+Round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to all the second-round blocks, and thus ratifying the first-round leader block, which is thus finalised by the wave.
+Low-throughput good case: The wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after round $r$ has advanced.
+High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced.
+Figure 1. A good-case wave $w$.
+We are now ready to define quiescence; the definition is mutually recursive with the definitions of endorsement, ratification, and finality.
+For the base, the $0^{th}$-wave is quiescent and its constitutional genesis block final.
+A subsequent wave is quiescent if it includes a final leader block $b$, and all the blocks in the wave except $b$ have empty payloads.
+Thus, the protocol begins quiescent and remains in low throughput mode as long as there is at most one agent issuing transactions in each wave.
+It switches to high throughput mode whenever multiple agents send transactions in the same wave.
+Ordering.
+As in other DAG-based protocols (Keidar et al., 2021, 2023; Lewis-Pye and Shapiro, 2025), once a leader block $b$ is finalised, the protocol deterministically orders all the transactions approved by $b$ that have not yet been totally ordered, and outputs the resulting sequence.
+Thus, as long as new blocks are being finalised, all the transactions in the system become ordered.
+Ordering must be consistent at all agents.
+To this end, we leverage the property that at most one block is ratified or final in a wave.
+We define a recursive ordering function $\tau$.
+When called for a block $b$, $\tau$ locates the most recent ratified and yet unordered block in $b]\setminus\{b\}$ and recursively calls $\tau$ for this block.
+Once no such block exists, $\tau$ topologically sorts all unordered blocks approved by $b$.
+3.2. Liveness: Dissemination and Round Progress Rules
+In the previous section, we explained how transactions are finalised in the good case, when a first-round leader block is observed by a super-majority of endorsements in the second round and a super-majority of ratifications in the third.
+We now discuss how we continue to progress through rounds when the good case does not occur, while taking care not to rush to the next round before giving the leader a chance to be finalised in case we are after GST.
+To formalize these progress rules, we first define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely.
+In second and third rounds, all correct participants send blocks, and these rounds advance once they include a super-majority of blocks.
+In a first round, only a leader (formal or spontaneous) sends a block in the good case.
+Hence a round is advanced once it includes a leader block.
+Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own.
+Once the blocklace includes a super-majority of round-one blocks, the round is advanced despite not including a leader block.
+A second challenge is partial dissemination, whereby a faulty agent sends a block to some participants and not others, causing missing dependencies and potentially preventing some correct agents—including the formal leader—from advancing to the next round.
+Partial dissemination is handled using two special dissemination-inducing blocks, which trigger block forwarding and are not added to the blocklace: a nack-block for missing dependencies and an inform-block for an unresponsive leader.
+Before sending dissemination-inducing blocks, we wait sufficiently long for messages from correct participants to arrive after GST, so as not to induce extra traffic in the good case.
+Similarly, we have to carefully set the timeout for sending non-leader first-round blocks so as to give a correct leader sufficient time to make up any missing blocks and send its new block.
+The delays used in the protocol are explained in Appendix D.
+3.3. Single-Epoch Constitutional Consensus Pseudocode
+In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$.
+An epoch identifier $e$ points to the epoch’s constitutional genesis block.
+$B^{e}$ denotes the blocklace of epoch $e$ in $B$.
+Algorithm 1 (Appendix B) presents the single-epoch Constitutional Consensus pseudocode, in which the prevailing constitution is fixed throughout the run.
+We use the terminology send/issue once as a shorthand for send/issue if the block has not been sent/issued yet.
+The protocol makes judicious use of delays, to ensure both liveness and efficiency. The rationale for the various protocol delays is given in Appendix D.
+Initialisation and input.
+The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$.
+As input transactions arrive, they are added to payload.
+Output: Once the protocol finalises a block $b$, it orders the unordered blocks approved by $b$ using a recursive function $\tau$ defined in Appendix A.3, and outputs their transactions.
+Latency and complexity.
+After GST and in the good case, a leader block is finalised within $3\delta$.
+The amortized communication complexity per transaction is $O(n)$ during high throughput; details are in Appendix C.
+4. Sybil-Resilient Democratic Decision Processes
+The question of how a democratic constitution can be amended, democratically, is deep and fundamental (Buchanan and Tullock, 1962).
+Here we develop democratic decision rules for amending each component of the constitution $(P,\sigma,\Delta)$, grounding our approach in the theory of computational social choice (Brandt et al., 2016), using state-of-the-art machinery that fits to our setting, which could be summarised by the following three properties:
+  1. (1)
+    Reality-aware (Shapiro and Talmon, 2017): The only type of decision is to amend the status quo; in particular, to amend the prevailing constitution (Abramowitz et al., 2021b).
+Thus, the prevailing constitution $(P,\sigma,\Delta)$ is an ever-present reality: it governs the running consensus protocol regardless of whether anyone wishes to change it, with amendment rules treating it as the default and move away from it only when there is sufficient support.
+  2. (2)
+    Sybil-resilient (Shahaf et al., 2019; Meir et al., 2024): Change of the status-quo can be effected only by support large enough to necessarily reflect the will of the non-Sybil agents.
+[^2] Sybil-resilient admission protocols (Poupko et al., 2021; Shahaf et al., 2020) during community formation are complementary to this work.
+  3. (3)
+    Aggregation over metric spaces with single-peaked domains (Bulteau et al., 2021): We assume that: (1) each agent has an ideal point in each of the decision domains; (2) the domain has a metric, so that an agent prefers any point closer to their ideal point over any point further away from it; and (3) there exists an aggregation function that, given the ideal points of agents, produces a point that corresponds to the agent opinions.
+This allows unifying proposing and voting: Agents just state their ideal points and the aggregation function identifies a most supported point; if those that prefer this point over the status-quo are a supermajority, it becomes the new decision.
+These properties allow for unified proposing and voting, (see Appendix F).
+This means that participants only need only agree on a set of votes, and then each agent can independently apply the amendment rules (described below) to determine the next constitution.
+Constitutional amendment by consensus.
+At a fixed periodic deadline (e.g., every Tuesday at midnight GMT), [^3] Real-world realizations may wish to add response mechanisms for emergency situations.
+participants use the consensus protocol to disseminate votes: At the periodic deadline, each agent $p_{j}$ creates a transaction $T_{j}$ holding a set of signed votes including the latest vote it knows from each participant.
+The agent proposes this transaction as input to the single-epoch consensus protocol.
+Once ordered, these transactions are not output to users but instead provided as input to the amendment protocol, run independently by each agent.
+Once the amendment protocol gets, for the first time, a $\sigma$-supermajority of such ordered transactions $T_{j}$ in an epoch $i$ of an instance of the consensus protocol, it creates a set holding the latest vote from each participant included in any of the received sets.
+The amendment protocol determines the next constitution based on this set and produces a constitutional amendment decision.
+Note that the amendment protocol produces at most one amendment per protocol epoch.
+Amendment rules.
+We now specify the concrete amendment rules given the agreed upon set of votes and the prevailing $(P,\sigma,\Delta)$; the updated constitution $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ takes effect when the consensus protocol switches to a new epoch, as described in Section 5.
+(1) Amending $P$.
+Each agent $i\in P$ votes independently on each potential participant $p$: yes (in) or no (out).
+A candidate $p$ is added to (or retained in) $P^{\prime}$ if and only if a $\sigma$-supermajority of the prevailing $P$ votes yes.
+Additionally, every new member in $P^{\prime}\setminus P$ must consent to joining, to prevent deadlock from non-responsive members (Lamport, 1998).
+This per-member $\sigma$-supermajority rule follows directly from the Sybil-resilient setting of Shahaf et al.
+(Shahaf et al., 2019) (for the setting of voting in a binary domain-in/out), who show that it is safe with respect to honest agents: a minority of Sybils cannot force membership changes.
+(2) Amending $\sigma$.
+Each agent $i\in P$ states its preferred $\sigma_{i}\in[0.5,1)$.
+The value of $\sigma$ is updated according to the $h$-rule for constitutional amendment (Abramowitz et al., 2021b) (named after the $h$-index of citation metrics), as follows: (1) Raising $\sigma$: $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which a $\sigma^{\prime}$-supermajority of $P$ voted for a value $\geq\sigma^{\prime}$.
+(2) Lowering $\sigma$: $\sigma$ is decreased to the minimal $\sigma^{\prime}$ for which a $\sigma$-supermajority of $P$ voted for a value $\leq\sigma^{\prime}$.
+(3) If neither condition holds, $\sigma$ remains unchanged.
+For example, changing $\sigma$ from $\frac{2}{3}$ to $\frac{3}{4}$ requires the approval of a $\frac{3}{4}$-supermajority. Changing to unanimity requires approval by all members of $P$.
+The $h$-rule is self-referential by design: raising the threshold to $\sigma^{\prime}$ requires a $\sigma^{\prime}$-supermajority supporting it (preventing a minority from imposing stricter requirements), while lowering requires the current $\sigma$-supermajority to agree (preventing a minority from weakening the constitution).
+Abramowitz et al.
+(Abramowitz et al., 2021b) show that the $h$-rule is the unique rule satisfying a natural set of axioms for this self-referential setting.
+(3) Amending $\Delta$.
+Each agent $i\in P$ states its preferred $\Delta_{i}\in\mathbb{R}^{+}$.
+Since $\Delta$ is a one-dimensional, single-peaked domain with potential Sybil presence, we use the Suppress Outer-$f$ parameter update rule of Shahaf et al.
+(Shahaf et al., 2019): [^4] Equivalently, we can use the Reality-Enforcing rule of Meir et al.
+(Meir et al., 2022), which is defined differently but is mathematically equivalent.
+In the original reference (Shahaf et al., 2019) the rule refers to ‘outer-$\sigma$’; we use ‘outer-$f$’ here to avoid confusion with our use of $\sigma$.
+  • If the median vote $\Delta$: discard the maximal $f$ votes, compute the median $\Delta^{\prime}$ of the remaining votes, and update to $\Delta^{\prime}$ if $\Delta^{\prime}>\Delta$.
+  • Symmetrically, if the median vote $\Delta$: discard the minimal $f$ votes, compute the median $\Delta^{\prime}$ of the remaining votes, and update to $\Delta^{\prime}$ if $\Delta^{\prime}<\Delta$.
+  • Otherwise, $\Delta$ remains unchanged.
+The key properties: if all Sybil agents wish to increase $\Delta$ (to slow down the protocol) but the correct agents do not, $\Delta$ will not increase; symmetrically for decreasing.
+If all correct agents unanimously wish to update $\Delta$ to some $\Delta^{\prime}$, the update succeeds regardless of adversarial votes, provided $f<\frac{n}{3}$.
+A constitution $(P,\sigma,\Delta)$ is amended democratically to $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ via a periodic cycle, using the prevailing $(P,\sigma,\Delta)$ as status quo: (1) $P\to P^{\prime}$: per-member $\sigma$-supermajority among $P$, plus consent of all new members in $P^{\prime}\setminus P$ (Shahaf et al., 2019).
+(2) $\sigma\to\sigma^{\prime}$: the $h$-rule for constitutional amendment (Abramowitz et al., 2021b).
+(3) $\Delta\to\Delta^{\prime}$: the Suppress Outer-$f$ parameter update rule (Shahaf et al., 2019).
+5. Constitutional Consensus with Constitutional Amendments
+We now extend the consensus protocol with constitutional amendments, which generalise reconfiguration in both scope (any protocol parameter, not just membership) and authority (decided by participants through constitutionally-specified processes, not by a leader or external entity); a comparison with reconfiguration work is given in Appendix J.
+The protocol’s correctness is proved in Appendix H, and it being grassroots (Shapiro, 2023a) in Appendix I.
+The democratic decision processes described in Section 4 are part of the constitution: they specify how each amendable component may be changed.
+The result of the amendment protocol described above is a constitutional amendment decision:
+Definition 5.1 (Constitutional Amendment Decision).
+A constitutional amendment decision is a pair $d=(h,x)$, where $x=(\textsc{amend},\textit{Id},i,C,C^{\prime})$, Id is an instance identifier, $i\geq 1$ is the epoch within this instance, $C,C^{\prime}$ are constitutions, referred to as the old and new constitutions of $d$, resp., and $h$ is a set of signed votes.
+If $i=1$ then $C=\bot$ and $P_{1}$ are called the founders of the protocol instance Id—they are the signatories of its initial constitution and the participants of the first epoch.
+The decision $d$ is valid in epoch $i$ of instance $Id$ with prevailing constitution $C=(P,\sigma,\Delta)$ if (1) $h$ consists of a $\sigma$-supermajority in $P$ of signed votes for epoch $i$ in instance $Id$, (2) these are the first $\sigma$-supermajority of signed votes ordered in epoch $i$ of the constitutional consensus protocol instance $Id$, and (3) $C^{\prime}$ is the result of applying the amendment protocol to $h$ with prevailing constitution $C$.
+Thanks to the use of consensus to order the votes and the democratic decision process, there is a unique valid amendment decision in each epoch of each instance of the protocol.
+Constitutional amendments cause epoch transitions, which proceed in three phases.
+Amendment finalisation: Once the amendment protocol at a participant $p$ results in a valid amendment $d$, $p$ includes $d$ as payload in every block it issues until $d$ is finalised.
+End epoch: upon finalising a constitutional block with $d$, an agent produces a coronate block, sends it to both old and new populations, and ceases participation; it must still respond to belated nack requests from agents finalising the old epoch.
+Start epoch: a member of the new population commences the new epoch upon receiving a $\sigma$-supermajority of coronate blocks (and having sent one, if also in the old population).
+The full pseudocode is given in Algorithm 2 (Appendix G).
+Liveness of constitutional amendments is guaranteed after GST since agents periodically send sets of votes, which after GST all become ordered.
+6. Related Work
+A full survey of related work is provided in Appendix J, organised into five areas.
+It reviews smart contracts and digital social contracts (De Filippi et al., 2021; Goodman, 2014; Cardelli et al., 2020); DAOs and blockchain governance, including self-amending blockchains (Hassan and De Filippi, 2021; Wood, 2016; Corduan et al., 2023; Kwon and Buchman, 2016; Camenisch et al., 2022), DAO frameworks (Cuende and Izquierdo, 2017; Leshner, 2020; OpenZeppelin, 2021; Marino and Snapshot Labs, 2020; Optimism Foundation, 2022), critiques of plutocratic governance (Schneider, 2022; Siddarth et al., 2020; Talmon, 2023; Buterin, 2021), and consortium platforms (Androulaki et al., 2018; Brown et al., 2016); reconfiguration of consensus protocols (Lamport, 1998; Lamport et al., 2010, 2009; Aguilera et al., 2011; Spiegelman et al., 2017; Lorch et al., 2006; Ongaro and Ousterhout, 2014; Bessani et al., 2014; Duan and Zhang, 2022; Komatovic et al., 2025; Vizier and Gramoli, 2020; Bessani et al., 2020); BFT consensus protocols, both leader-based (Castro and Liskov, 1999; Yin et al., 2019; Buchman, 2016) and DAG-based (Keidar et al., 2021; Danezis et al., 2021; Giridharan et al., 2022; Gelashvili et al., 2022; Babel et al., 2023; Raikwar et al., 2024), including the two protocols on which Constitutional Consensus builds (Keidar et al., 2023; Lewis-Pye and Shapiro, 2025); and grassroots protocols and democratic identity, including centralised democratic platforms (Barandiaran et al., 2024; Loomio Cooperative Limited, 2024; Royo et al., 2020; Consul Democracy Foundation, 2025; Behrens et al., 2018) and proof-of-personhood systems (Ford, 2020; Borge et al., 2017; Kavazi et al., 2023; Idena Team, 2018; Brenzikofer, 2019).
+Constitutional Consensus combines three properties—democratic one-person-one-vote governance, decentralised BFT consensus, and self-governing constitutional amendment—and to the best of our knowledge, no existing system achieves all three.
+7. Conclusions
+The Constitutional Consensus protocol was devised to serve the needs of future grassroots platforms that operate digital social contracts.
+A key design principle is that governance is democratic: participants themselves decide on constitutional amendments—who participates, the supermajority threshold, and the timeout—through decision processes grounded in computational social choice.
+Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve.
+As communities federate to form ever-larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow.
+Until then, simulating the various scenarios would be productive future work.
+Another direction is making the democratic decision-making process an amendable component of the constitution, making the system reflexive: the amendment process would then be subject to amendment through the very process it specifies.
+Appendix A Formal Definitions
+A.1. Blocklace Preliminaries
 We use $a\neq b\in X$ as a shorthand for $a\in X\wedge b\in X\wedge a\neq b$.
@@
 To help avoid confusion, we use $B$ to denote a closed set of blocks (a blocklace) and $D$ or $X$ to denote any set of blocks, not necessarily closed.
-Also, we refer to a set of blocks by agents in $P\subset\Pi$ as $P$-blocks, the let the $P$-blocks in $B$ denote the set of all blocks in $B$ by agents in $P$.
+Also, we refer to a set of blocks by agents in $P\subset\Pi$ as $P$-blocks, and we let the $P$-blocks in $B$ denote the set of all blocks in $B$ by agents in $P$.
 Given a set of blocks $D$, a pointer to a block $b$ is dangling in $D$ if $b\notin D$.
@@
 A round $r\geq 0$ in $B$ is the set of all blocks $b\in B$ with $\textit{depth}(b)=r$, and the $r$-prefix of $B$, $B_{r}$, consists of all blocks $b\in B$ with $\textit{depth}(b)\leq r$, equivalently all rounds in $B$ up to and including $r$.
-Note that, by definition, blocks of the same round do not observe each other.
-3.2 Protocol Concepts
-The blocklace created during an epoch begins with a constitutional genesis block $(h,d)$ where $d$ specifies a new constitution $C=(P,\sigma,\Delta)$.
-In fact $d$ also records the constitutional decision to amend the old constitution (if any) to the new constitution and is signed by multiple agents; these are discussed in Section 5.
-Here, we focus on the behaviour of the protocol during a single epoch.
-For now we refer to a blocklace in which all initial blocks are constitutional genesis blocks as a constitutional blocklace.
-Every participant $p\in P$ commences an epoch $b$ with its blocklace including $b$ and its local epoch identifier $e$ set to $b$, $e\leftarrow b$.
-The protocol employs an evolving shared blocklace.
-Each participant $p\in P$ maintains a local copy of the constitutional blocklace.
-In the course of an epoch, agents create new blocks that observe the epoch’s genesis block and and disseminate them to other agents.
-Blocks may include transactions the agents wish to order, if they have any.
-While a blocklace is a partially ordered set of blocks, the protocol totally orders the blocks in an epoch in a consistent way, producing the same output sequence at all participants in an epoch.
-Moreover, epochs are also totally ordered, so that the totality of all blocks of all epochs are also totally ordered, even though not always by the same participants
-Because correct agents follow the protocol, they do not equivocate.
-Hence, for a blocklace $B$ created by the protocol with constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
-Waves.
-Given a constitutional blocklace $B$, a wave $W\subseteq B$ in a blocklace is either its constitutional genesis block, referred to as its $0^{th}$-wave, or the set of blocks of depths $r+1$, $r+2$, and $r+3$ in $B$, referred to as the the first, second, and third round of the wave, respectively, for any $r\geq 0$ such that $r\mod 3=0$.
-Such a wave is the $k^{th}$ wave of $B$, $k:=(r+3)/3$.
+Note that blocks of the same round do not observe each other.
+A.2. Protocol concepts
+Overview.
+We begin by disregarding equivocations, namely, two blocks by the same agent that do not observe each other.
+We say that a block $b$ approves a block $b^{\prime}$ if $b$ observes $b^{\prime}$ and does not observe any block equivocating with $b^{\prime}$.
+We use this notion to define endorsements and ratification.
+A second-round block $b$ of a wave may approve multiple first-round blocks in its wave, but endorses (towards finality) at most one of them, depending on the mode.
+The mode is determined according the quiescence property of the wave $W$ that precedes $b$ in $[b]$, which we shall define later.
+If $W$ is quiescent and $b$ approves exactly one round-one block in its wave, $b^{\prime}$, then $b$ endorses $b^{\prime}$.
+If $W$ is non-quiescent and $b$ approves a formal leader’s round-one block, $b^{\prime}$, in its wave, then $b$ endorses $b^{\prime}$.
+In case neither condition is satisfied, $b$ does not endorse any block.
+A first-round block $b$ is ratified by a third-round block approving a $\sigma$-supermajority in $P$ of second-round blocks that endorse $b$.
+And a first-round block $b$ is final once there is a $\sigma$-supermajority in $P$ of third-round blocks that ratify $b$.
+Figure 1 illustrates a “good case” wave in which the leader block is finalised.
+Observe that a wave can have at most one ratified or final block.
+This is because a second-round block endorses at most one first-round block, and since correct agents do not equivocate, there cannot be two $\sigma$-supermajorities in the same round endorsing different blocks.
 For a given $P\subset\Pi$, the function leaderP maps integers to $P$.
 For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$.
-We assume the leader function is defined so that it is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
-Finalizing blocks in waves.
-The consensus protocol commences with a constitutional blocklace and progresses in waves.
-The protocol strives to finalize a new block in each wave.
-The first round of a wave includes candidates for finality, the second round includes endorsements of candidates, and the third round includes ratifications.
-A first round block is a candidate for finality in two cases—if it is a formal leader block or the only block in its round.
-These two possibilities typically arise in high- and low-throughput scenarios, respectively.
-During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with any others, all agents produce second and third round blocks that finalize $b$, in which case $b$ is referred to as spontaneous leader.
-During high throughput, multiple agents may have transactions to send, and the leader function is used to break symmetry, so that only the formal leader sends a leader block in the first round.
-Finalized leader blocks of both types induce the total ordering of all blocks, as in other DAG-based protocols [17, 18, 23].
-A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows.
-Note that the next two definitions are mutually-recursive.
-Definition 3.1 (Endorse).
+We assume the leader function is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
+Definition A.1 (Endorse).
 Let $b^{\prime}$ be a second round block that approves a first round block $b$ in a wave $W$ in the blocklace $B=[b^{\prime}]$. Let $(P,\sigma)$ be the prevailing constitution.
 Then $b^{\prime}$ endorses $b$ if the most-recent wave $W^{\prime}\subset B$ that precedes $b$ is:
-  1. 1.
+  1. (1)
     quiescent and $b$ is the only first-round block approved by $b^{\prime}$, or
-  2. 2.
+  2. (2)
     non-quiescent and $b$ is a $\textit{leader}_{P}$ block approved by $b^{\prime}$.
@@
 Formally:
-Definition 3.2 (Ratified, Final, Ordered, Quiescent).
+Definition A.2 (Ratified, Final, Ordered, Quiescent).
 Assume a blocklace $B$ and a wave $W\subseteq B$. A first-round block $b\in W$ with prevailing constitution $(P,\sigma)$ is:
-  1. 1.
+  1. (1)
     ratified by a third-round block $b^{\prime}\in W$ if $b^{\prime}$ approves a $\sigma$-supermajority in $P$ that endorses $b$, in which case $b$ is also ratified in $W$ and in $B$; and
-  2. 2.
+  2. (2)
     final in $W$ and in $B$ if the third-round of $W$ includes a $\sigma$-supermajority in $P$, referred to as a finalising $\sigma$-supermajority, in which each block ratifies $b$.
 The wave $W$ is:
-  1. 1.
+  1. (1)
     finalising if it includes a block $b$ that is final in $W$; and
-  2. 2.
+  2. (2)
     quiescent if, in addition, all blocks in $W$ except perhaps $b$ are empty, and no block in $B$ conflicts with $b$.
-In addition, the $0^{th}$-wave is defined to be quiescent and its constitutional genesis block to be final.
+In addition, the $0^{th}$-wave is quiescent and its constitutional genesis block final.
 Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final.
 Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered.
-We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$.
-Note, however, that quiescence is not monotonic, because conflicting and non-empty blocks may be added to a round.
-The next observation asserts that a wave’s final block is unique. {observation}[Uniqueness] A wave can have at most one ratified or final block.
-Proof 3.3.
-The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
-Quiescence, low- and high-throughput scenarios.
-The protocol is initially quiescent.
-It begins operations and continues to operate as long as agents have payloads to send.
-If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode.
-It returns to being quiescent following a quiescent wave.
-Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block, which is thus finalized by the wave.
-Low-throughput good case: The previous wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after the previous round $r$ has advanced, followed promptly by two rounds of empty blocks, resulting in a finalising and quiescent wave.
-High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced and all (or many) of the wave’s blocks are non-empty.
-Figure 1: A good-case wave $w$.
-We first consider the good-case scenarios summarized in Figure 1.
-During low-throughput all waves are quiescent.
-A wave commences when some agent has a new transaction.
-This agent issues a (first-round) non-empty block, and all agents follow promptly with two rounds of empty blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final, with the new wave also being quiescent.
-During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty, with transactions to be ordered by the the next final leader block.
-Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own.
-Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block, with blocks by correct agents to be ordered by the leader of some subsequent finalising wave.
-Similarly, if during low-throughput there are multiple agents issuing first-round blocks, then typically none of the first-round blocks will become final.
-In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
-To formalize these progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
-Definition 3.4 (Advanced Round).
+We observe that the block properties defined above are monotonic:
+Observation A.3 (Monotonicity).
+If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$.
+To formalize the round progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
+Definition A.4 (Advanced Round).
 Assume a constitutional blocklace $B$ with prevailing constitution $(P,\sigma)$. A round $r$ in a wave $W\subseteq B$ is advanced if it is a:
-  1. 1.
+  1. (1)
     first round that either (i) includes a $\sigma$-supermajority of blocks among $P$, or (ii) includes a formal leader block or (iii) is preceded by a quiescent wave in $B$ and includes at least one block, or a
-  2. 2.
+  2. (2)
     second or third round that includes blocks by a $\sigma$-supermajority among $P$.
@@
 Agents incorporate blocks in the blocklace only if they are valid, namely could have been produced by a correct agent during the run of the protocol. Formally:
-Definition 3.5 (Valid Block).
-An ordinary block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
+Definition A.5 (Valid Block).
+A block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
 Dissemination.
@@
 Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
-Protocols such as TCP achieve the reliable transfer of a sequence of bytes from the sender to the recipient by employing ack and nack messages for transmission control.
-The purpose of an ack message is to acknowledge the receipt of a datagram, and of a nack message to report an out-of-order datagram to its sender, e.g., in the case of receiving the $(k+1)^{th}$-datagram before receiving the $k^{th}$-datagram.
-In the context of the blocklace, an ‘out-of-order’ block is a block received before some block it points to, and the purpose of a nack-block is indeed to report of such an out-of-order block to its creator (which is not necessarily its sender).
 nack blocks are used to achieve reliable broadcast of ordinary blocks.
@@
 These are the tips of $b]\setminus(B\cup\{b\})$.
-Hence the following definition:
-Definition 3.6 (nack-block).
+Definition A.6 (nack-block).
 Given a blocklace $B$, a nack-block by $p$ for a block $b=(h,x,H)$ such that $b]\not\subseteq B\cup\{b\}$ is the $p$-block $b^{\prime}=(h^{\prime},x^{\prime},H^{\prime})$ where $x^{\prime}=(\textsc{nack},h)$ and $H^{\prime}\neq\emptyset$ contains the tips of $b]\setminus(B\cup\{b\})$.
@@
 If the nack-block $b^{\prime}$ is defined, then its pointers $H^{\prime}$ are precisely the non-empty subset of $H$ that is dangling in $B$.
-The reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block functions as an ack-block, as it attests to all the blocks known to $p$ at the time of its creation.
-nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it.
-In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
-Definition 3.7 (inform-Block).
+nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
+Definition A.7 (inform-Block).
 Given a blocklace $B$, an inform-block $(h,x,H)$ for round $r+1>1$ has payload $x=\textsc{inform}$ and pointers $H$ to the blocks of depth $r$ in $B$.
@@
 If the leader indeed misses any of these blocks, which may happen even after GST due to Byzantine partial dissemination, it will respond to $p$ with a nack-block, which will in turn cause $p$ to send to the leader the blocks it needs.
-3.3 Single-Epoch Constitutional Consensus Pseudocode
-In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$.
-An epoch identifier $e$ points to the epoch’s constitutional genesis block.
-We use $B^{e}$ to denote the blocklace of epoch $e$ in $B$, which is the subset of $B$ that observes $e$.
-The protocol is event-driven.
-Algorithm 1 describes a single-epoch Constitutional Consensus, in which the prevailing constitution is fixed throughout the run.
-Initialisation.
-The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$.
-The initialisation of subsequent epochs is discussed in Algorithm 2.
-Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Receive: Upon receipt of a $q$-block $b$, $q\in P$
+A.3. The Ordering Function $\tau$
+Definition A.8 ($\tau$).
+Assume a safe constitution $(P,\sigma)$ and any fixed topological sort function $\textit{xsort}(b,B,B^{\prime})$ (exclude and sort) that takes a block $b$ and two blocklaces $B\supset B^{\prime}$, and returns a sequence that preserves the ‘observes’ partial order and includes all the non-empty blocks in $B\setminus B^{\prime}$ that are approved by $b$.
+The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
+[…]
+When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$.
+In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
+Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$.
+Input:: Upon input a transaction $T$, $\emph{payload}\leftarrow\emph{payload}\cup\{T\}$ Receive:: Upon receipt of a $q$-block $b$, $q\in P$
 if $b$ is a nack-block for a $p$-block then judiciously send $[b]$ to $q$,
@@
 else if $b$ is a valid ordinary block then $D\leftarrow D\cup\{b\}$.
-Accept or Nack: For every $q$-block $b\in D$, $q\in P$,
+Accept or Nack:: For every $q$-block $b\in D$, $q\in P$,
 if $b]\subseteq B\cup\{b\}$ then $B\leftarrow B\cup\{b\}$, $D\leftarrow D\setminus\{b\}$,
 else if a $b$ was added to $D$ more than $\Delta$ time ago
-——- then send once a nack-block for $b^{\prime}$ to $q$.
-Issue: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: 1.
-quiescent and $\emph{payload}\neq\bot$, or 2.
-non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$.
-Issue Backlog: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block.
-Inform Leader: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$.
-Output: If $B$ contains a final block $b$ of depth greater than the previous final block then output every block in $\tau(b)$ (Definition 3.8) unless it has already been output.
-issue a round $k$ block means create a block $b$ with payload and pointers to the $P$-blocks that are tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$.
+——- then send once a nack-block for $b$ to $q$.
+Issue:: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: (1): quiescent and $\emph{payload}\neq\bot$, or (2): non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$.
+Issue Backlog:: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block.
+Inform Leader:: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$.
+Output:: If $B$ contains a final block $b$ of depth greater than the previous final block then output all the transactions in every block in $\tau(b)$ (Definition A.8) that have not already been output.
+issue a round $k$ block means create a block $b$ with payload and pointers to the tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$.
 judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
 Algorithm 1 Constitutional Consensus for a single epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$. Code for agent $p\in P$.
+Appendix B Pseudocode Walkthrough
+Receive: valid incoming ordinary blocks are inserted to $D$.
+Subsequently, the Accept or Nack is applied: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
+Issue: The variable $r$ always holds the highest advanced round in $B$, and an agent typically issues blocks in round $r+1$.
+This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced.
+This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
+The only exception (to sending in round $r+1$) is the Issue Backlog rule: In case $p$ has payload to send, has not sent a round $r$ block, and is not supposed to send a round $r+1$ block, $p$ issues a round $r$ block to ensure its payload is disseminated.
+If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block.
+If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode.
+During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send.
+During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader.
+Here, a long timeout is used ($9\Delta$), before which we inform the leader that the round has advanced through the Inform Leader rule: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
+Appendix C Latency and Communication Complexity
+Latency. After GST and in the good case a leader block (formal or spontaneous) is finalised by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
+Communication complexity.
+During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all.
+In Morpheus (Lewis-Pye and Shapiro, 2025), this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
+A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead.
+Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block.
+A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block.
+Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity.
+During high-throughput, all blocks carry payloads.
+If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
+Appendix D Single-Epoch Protocol: Delays
 Delays. The protocol makes judicious use of delays, to ensure both liveness and efficiency. Here is the rationale for the various protocol delays, all expressed as multiples of $\Delta$:
@@
 If $p$ does not accept any leader-block by $t+9\Delta$, it concludes that either $p_{\ell}$ is faulty or GST has not arrived yet, and sends a first-round block of its own in order to allow the protocol to proceed to the next round.
-Protocol walk-through.
-Receive: valid incoming ordinary blocks are inserted to $D$.
-Accept or Nack: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
-Issue: The variable $r$ always holds the highest advanced round in $B$, and an agent typically issues blocks in round $r+1$.
-This means that rounds can be skipped if the agent has been slow sending while the blocklace has advanced.
-This is useful when an agent that has been offline for some time catches up with the current blocklace before beginning to send blocks of its own.
-The only exception (to sending in round $r+1$) is the Issue Backlog rule: In case $p$ has payload to send, has not sent a round $r$ block, and is not supposed to send a round $r+1$ block, $p$ issues a round $r$ block to ensure its payload is disseminated.
-If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block.
-If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode.
-During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send.
-During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader.
-Here, a long timeout is used ($9\Delta$), before which we try to inform the leader that the round has advanced to give it a chance to progress.
-Inform Leader: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
-Output: Once the protocol finalizes a block $b$, it orders the blocks approved by $b$ that have not been ordered yet using the function $\tau$.
-The definition of $\tau$ is essentially-identical to that of Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks.
-The output of $\tau(b)$ is a sequence that includes all blocks approved by $b$.
-Formally:
-Definition 3.8 ($\tau$).
-Assume a safe constitution $(P,\sigma)$ and any fixed topological sort function $\textit{xsort}(b,B,B^{\prime})$ (exclude and sort) that takes a block $b$ and two blocklaces $B\supset B^{\prime}$, and returns a sequence that preserves the ‘observes’ partial order and includes all the non-empty blocks in $B\setminus B^{\prime}$ that are approved by $b$.
-The function $\tau$ takes a block as input and returns a sequence of blocks in $[b]$ as output, as follows:
-[…]
-When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$.
-Note that $b^{\prime}$ is not necessarily final in $[b]$, nor final in the closure of the final block with which $\tau$ was called initially.
-In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
-Proposition 3.9.
-If the prevailing constitution is safe then $\tau(b)$ does not include equivocations.
-Proof 3.10.
-Consider an equivocation $b_{1},b_{2}\in[b]$.
-Consider the first call $\textit{xsort}(b^{\prime},[b^{\prime}],[b^{\prime\prime}])$ for which at least one of $b_{1}$ or $b_{2}$ is in $X:=[b^{\prime}]\setminus[b^{\prime\prime}]$, and the following two cases:
-  1. 1.
-    Both $b_{1},b_{2}\in X$: Then $b^{\prime}$ approves neither hence neither will be in the output of xsort and hence nor in the output of $\tau$.
-  2. 2.
-    Only one of them is in $X$, wlog $b_{1}\in X$, and since this is the first such call to xsort it must be that $b_{2}\in[b^{\prime\prime}]$.
-Then $b^{\prime}$ observes both $b_{1}$ and $b_{2}$ and hence will approve neither, whereas $b^{\prime\prime}$ observes only $b_{2}$, therefore will approve it and therefore only $b_{2}$ will occur in the output of $\tau$, if at all (assuming there are no other equivocations with $b_{2}$).
-In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$.
-Safety and liveness.
-We assume that the prevailing constitution is safe and live (Definition 2.2).
-In Section 4 we prove that $\tau$ is monotonic.
-This ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor.
-Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any.
-In addition, as we shall see, during a protocol run the outputs of calls to $\tau$ by correct agents always occur in the same order, which is the basis for protocol safety.
-Liveness is achieved after GST.
-Whenever there is a post-GST wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block.
-Therefore, no agent can produce a valid second round block that does not endorse the leader block.
-So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized.
-Once any agent has payload to send following a quiescent wave, the agent issues a block, and either the block is finalized in the same wave or the wave is non-quiescent and the protocol switches to high-throughput mode.
-Latency and complexity.
-Regarding latency, we note that after GST and in the good case a leader block (formal or spontaneous) is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
-Regarding communication complexity: During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all.
-In Morpheus [23], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
-A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead.
-Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block.
-A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block.
-Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity.
-To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.
-During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
-4 Safety and Liveness of Single-Epoch Constitutional Consensus
+Appendix E Safety and Liveness of Single-Epoch Constitutional Consensus
 Here we prove the safety and liveness of the single-epoch Constitutional Consensus protocol presented in Algorithm 1.
@@
 The output sequence at every correct participant begins with $C$ and does not include additional constitutions.
-4.1 Safety
-The safety properties of Definition 2.4 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
-{observation}
+E.1. Safety
+In this section, we prove the single-epoch protocol’s safety.
+Theorem E.1.
+If the founding constitution is safe then single-epoch Constitutional Consensus is safe.
+The safety properties (Consistency and Validity) in Definition 2.3 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
+Observation E.2.
 If the output sequences of every two correct agents are consistent with each other, then single-epoch consensus ensures safety.
-Proof 4.1.
+Proof.
 Validity trivially holds because every agent participates in every epoch.
 Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency.
-Theorem 4.2.
-If the founding constitution is safe then single-epoch Constitutional Consensus is safe.
-First, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
+∎
+The next observation asserts that a wave’s final block is unique.
+Observation E.3 (Uniqueness).
+A wave can have at most one ratified or final block.
+Proof.
+The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
+∎
+To prove consistency, first, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents.
 We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
+Proposition E.4.
+If the prevailing constitution is safe then $\tau(b)$ does not include equivocations.
+Proof.
+Consider an equivocation $b_{1},b_{2}\in[b]$.
+Consider the first call $\textit{xsort}(b^{\prime},[b^{\prime}],[b^{\prime\prime}])$ for which at least one of $b_{1}$ or $b_{2}$ is in $X:=[b^{\prime}]\setminus[b^{\prime\prime}]$, and the following two cases:
+  1. (1)
+    Both $b_{1},b_{2}\in X$: Then $b^{\prime}$ approves neither hence neither will be in the output of xsort and hence nor in the output of $\tau$.
+  2. (2)
+    Only one of them is in $X$, wlog $b_{1}\in X$, and since this is the first such call to xsort it must be that $b_{2}\in[b^{\prime\prime}]$.
+Then $b^{\prime}$ observes both $b_{1}$ and $b_{2}$ and hence will approve neither, whereas $b^{\prime\prime}$ observes only $b_{2}$, therefore will approve it and therefore only $b_{2}$ will occur in the output of $\tau$, if at all (assuming there are no other equivocations with $b_{2}$).
+In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$. ∎
 For the following proofs we generalize $\tau$ to blocklaces, so that $\tau(B)$ is $\tau(b)$ if $b$ is the most-recent final block in $B$ if there is one, else it is the empty sequence $\Lambda$.
-Definition 4.3 ($\tau$ Safety).
+Definition E.5 ($\tau$ Safety).
 A blocklace $B$ is $\tau$-safe if for every wave $W\subset B$ with a block $b$ final in $W$ and any subsequent wave $W^{\prime}\subset B$ with a block $b^{\prime}$ ratified in $W^{\prime}$, $b$ is ratified in $W\cap[b^{\prime}]$.
-Proposition 4.4.
+Proposition E.6.
 A constitutional blocklace with a safe constitution is $\tau$-safe.
-Proof 4.5.
+Proof.
 Let $B$ be a constitutional blocklace with a safe constitution $(P,\sigma)$, $W\subset B$ a wave with a block $b$ final in $W$, and $W^{\prime}\subset B$ a subsequent wave with a block $b^{\prime}$ ratified in $W^{\prime}$.
@@
 Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
-Proposition 4.6 ($\tau$ Monotonicity).
+∎
+Proposition E.7 ($\tau$ Monotonicity).
 $\tau$ is monotonic wrt $\supset$ over constitutionally-safe blocklaces.
-Proof 4.7.
+Proof.
 We have to show that if $B\subset B^{\prime}$ for blocklaces $B,B^{\prime}$ then $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
-By Observation 3.2, all the blocks final in $B$ are also final in $B^{\prime}$.
-By Observation 3.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth.
+By Observation A.3, all the blocks final in $B$ are also final in $B^{\prime}$.
+By Observation E.3, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth.
 If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially.
@@
 Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$.
-By Proposition 4.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
-{observation}
+By Proposition E.6 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
+∎
+Observation E.8.
 If the union $B\cup B^{\prime}$ of two blocklaces is constitutionally-safe then $\tau(B)$ and $\tau(B^{\prime})$ are consistent.
-Proof 4.8.
-By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition 4.6), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
-Proof 4.9 (Proof of Theorem 4.2).
+Proof.
+By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition E.7), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
+∎
+Proof of Theorem E.1.
 By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe.
-Hence, by Observation 4.1, the outputs of any two correct agents running Constitutional Consensus are always consistent.
-By Observation 4.1, the protocol is safe.
-4.2 Liveness
+Hence, by Observation E.8, the outputs of any two correct agents running Constitutional Consensus are always consistent.
+By Observation E.2, the protocol is safe.
+∎
+E.2. Liveness
 Next, we discuss liveness.
@@
 We conclude with the following observation:
-{observation}
-[Single-epoch Liveness] A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
-Theorem 4.10.
+Observation E.9 (Single-epoch Liveness).
+A single-epoch consensus protocol is live if and only if the output sequences of all correct participants include all the transactions input or output by correct participants.
+Theorem E.10.
 Single-epoch Constitutional Consensus is live.
 We first prove that Constitutional Consensus disseminates all correct agents’ blocks and the blocks they depend on.
-Proposition 4.11 (Dissemination).
+Proposition E.11 (Dissemination).
 If a block $b$ is issued by a correct agent in Constitutional Consensus then $[b]$ is eventually received by all correct agents.
-Proof 4.12.
+Proof.
 By way of contradiction assume that $p,q\in P$ are correct, $b_{p}$ is issued by $p$ (and received by $q$ since the network is reliable) and some block $b\in[b_{p}]$ is never received by $q$.
@@
 Since the network is reliable, $q$ will eventually received $b$, a contradiction.
+∎
 The following observation follows directly from the definition of block validity:
-{observation}
-[Sequential Advance] If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
+Observation E.12 (Sequential Advance).
+If a blocklace $B$ contains a round $r$ block, then every round $0<r^{\prime}<r$ is advanced in $B$.
 We next show that the protocol does not get “stuck”, i.e., agents continue to produce blocks when they are needed for progress.
@@
 But if some agent issues a block in a round, that round eventually advances.
-Proposition 4.13 (Advance).
+Proposition E.13 (Advance).
 If a correct agent $p$ sends a block in some round $r$ in Constitutional Consensus, then $r$ eventually becomes advanced at every correct agent.
-Proof 4.14.
+Proof.
 Let $p_{1}$ be an agent that sends a block $b$ in round $r$. Then when $p_{1}$ does so, r1r-1 is advanced in $p_{1}$’s blocklace.
 Assume by way of contradiction that $r$ does not advance at some correct agent $p_{2}$.
-By Observation 4.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace.
-By Proposition 4.11, this means that no correct agent sends blocks in rounds >r>r.
-Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition 4.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
-By Proposition 4.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
+By Observation E.12, no blocks of depth greater than $r$ are added to $p_{2}$’s blocklace.
+By Proposition E.11, this means that no correct agent sends blocks in rounds >r>r.
+Moreover, because $r$ does not advance at $p_{2}$, it must be the case that $p_{2}$ does not accept a supermajority of blocks in round $r$, which by Proposition E.11 implies that there is some correct agent $p_{3}$ that does not send a round $r$ block.
+By Proposition E.11, $p_{3}$ receives all the blocks in $[b]$. Therefore, $p_{3}$ accepts all these blocks into its blocklace. Thus, round r1r-1 is advanced in $p_{3}$’s blocklace.
 In case $r$ is a first round following a quiescent wave at $p_{3}$, $r$ advances at $p_{3}$ upon receipt of $b$, and $p_{3}$ issues a round $r+1$ block, a contradiction.
-Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction.
+Otherwise, $r$ is the maximum advanced round at $p_{3}$ and is not a first round following a quiescent wave at $p_{3}$. Therefore, $p_{3}$ issues a round $r$ block, a contradiction. ∎
 We next prove liveness for finite runs.
-Proposition 4.15 (Finite implies quiescence).
+Proposition E.14 (Finite implies quiescence).
 Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent $p$ is finite in a suffix of the run. Then $B$ is quiescent.
-Proof 4.16.
+Proof.
 Assume by way of contradiction that $p$’s latest advanced round in the run, $r$, is not quiescent.
 Then, by the Issue rule, $p$ eventually issues a block in round $r+1$ (at most $9\Delta$ time after $r$ is advanced).
-Proposition 4.17 (Block issuance).
+Proposition E.15 (Block issuance).
 If an agent $p$ has payload to send, then $p$ eventually issues a block with this payload.
-Proof 4.18.
+Proof.
 Assume by contradiction that $p$ has a non-empty payload from some time when $p$’s most advanced round is $r_{0}$ onward, and $p$ never issues a block with this payload.
 First, consider the case that $p$’s rounds do not advance in some suffix of the run.
-Then by Proposition 4.15, $p$’s blocklace is quiescent.
+Then by Proposition E.14, $p$’s blocklace is quiescent.
 Hence, condition 1 of the Issue rule is satisfied, and the payload is sent.
@@
 A contradiction.
-Proposition 4.19 (Transaction Liveness in finite runs).
+∎
+Proposition E.16 (Transaction Liveness in finite runs).
 Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all input transactions of correct agents are ordered in $B$.
-Proof 4.20.
-By Proposition 4.11, $B$ includes all ordinary blocks sent by correct agents.
-By Proposition 4.15, $B$ is quiescent.
+Proof.
+By Proposition E.11, $B$ includes all ordinary blocks sent by correct agents.
+By Proposition E.14, $B$ is quiescent.
 This means that its last wave includes a first round block $b$ that is final in $B$, and all other blocks in the wave are empty.
@@
 Also by assumption of $B$ being quiescent, all non-empty ordinary blocks in $B$ are in $[b]$, and hence ordered by $\tau(b)$.
-Moreover, by Proposition 4.17, no agents have payloads to send.
+Moreover, by Proposition E.15, no agents have payloads to send.
 It follows that all input transactions have been sent in ordinary blocks and ordered.
-Proposition 4.21 (Epoch Liveness in finite runs).
+∎
+Proposition E.17 (Epoch Liveness in finite runs).
 Consider a run of Constitutional Consensus in which the blocklace $B$ of a correct agent is finite in a suffix of the run. Then all transactions output by correct agents are ordered in $B$.
-Proof 4.22.
+Proof.
 Consider a transaction $t$ output by a correct agent $p_{1}$ and another correct agent $p_{2}$ whose blocklace $B_{2}$ is finite in a suffix of the run.
@@
 This super-majority includes at least one $p_{3}$-block $b^{\prime}$ by a correct agent $p_{3}$.
-By Proposition 4.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$.
-Because $B_{2}$ is finite in a suffix of the run, by Proposition 4.15, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
+By Proposition E.11, $p_{2}$ $b^{\prime}$ is included in $B_{2}$, and because $b\in[b]$, so is $b$.
+Because $B_{2}$ is finite in a suffix of the run, by Proposition E.14, it is quiescent, which means that all non-empty blocks including $b$ are final in $B_{2}$, as needed.
+∎
 Next, we show that after GST, all waves that have correct leaders in high throughput mode are finalising.
 Because quiescence is not monotonic, we consider here only waves $k$ for which the predecessor wave, k1k-1, is not quiescent at any correct agent at any time.
-Proposition 4.23 (First round progress after GST).
-Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt>GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time.
+Proposition E.18 (First round progress after GST).
+Let $r$ be a first round with a correct leader, let $p$ be a correct agent at which round r1r-1 advances at time tt¿GST, and assume that the wave that ends in round r1r-1 is not quiescent in any correct agent’s blocklace at any time.
 Then $p$ sends a block in round $r$ if and only if it is the round’s leader.
 Moreover, $p$ endorses a round $r$ leader block in round $r+1$.
-Proof 4.24.
+Proof.
 Let $p_{\ell}$ be the correct leader of round $r$.
@@
 Some predecessors of this block might be missing, causing $p$ to wait $\Delta$ and then send a nack-block and get the missing blocks allowing it to accept the leader block by time $t+9\Delta$.
-When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block.
+When the leader block is received, the round advances and $p$’s round $r+1$ block endorses the leader block. ∎
 Below, we say that a round $r$ advances at time $t$ if there is some correct agent $p$ such that round $r$ is advanced in $p$’s blocklace at time $t$ and $r$ is not advanced at any correct agent before time $t$.
-Proposition 4.25 (Progress after GST).
+Proposition E.19 (Progress after GST).
 Let $r$ be a first round with a correct leader s.t.
@@
 Then a round $r$ leader block is finalized in its wave at all correct agents.
-Proof 4.26.
-By Proposition 4.23, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$.
+Proof.
+By Proposition E.18, the only correct agent that sends a round $r$ block is the round’s leader and all correct agents endorse this block in round $r+1$.
 This means that it is impossible to create a valid round $r+1$ block that does not endorse the leader, as such a block would have to point to a super-majority of round $r$ blocks excluding the leader block.
 Therefore, every valid round $r+2$ block ratifies the leader block, and the block is finalized in round $r+2$ as soon as it advances at any correct agent.
+∎
 Next, we consider runs with infinitely many quiescent waves.
-Proposition 4.27 (Infinite quiescent waves).
+Proposition E.20 (Infinite quiescent waves).
 Consider a run of Constitutional Consensus in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time.
 Then all the blocks of correct agents are ordered by all agents in this run.
-Proof 4.28.
+Proof.
 Let $B$ be the (infinite) union of all blocklaces in the run.
 Let $b_{1}$ be a block of a correct agent in $B$.
-By Proposition 4.11, $b_{1}$ is eventually included in all correct agents’ blocklaces.
+By Proposition E.11, $b_{1}$ is eventually included in all correct agents’ blocklaces.
 Therefore, it is approved by all correct agents’ blocks from some wave onward.
@@
 Then $b_{3}$ is endorsed by a super-majority in the second round.
-Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition 4.11).
+Because this super-majority includes correct agents, all correct agents eventually accept $b_{3}$ (by Proposition E.11).
 If $b_{3}$ is not a leader block, then to be validly endorsed, $b_{3}]$ must be quiescent, and so upon receiving $b_{3}$, all the correct agents learn that $b_{2}$ is finalized.
@@
 In the former case the leader block is finalized, ordering all preceding ordinary blocks including $b_{2}$, and in the latter $b_{2}$ is already final in the block’s closure.
+∎
 We are now ready to prove the liveness theorem.
-Proof 4.29 (Proof of Theorem 4.10).
-By Observation 4.2, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent.
-By Proposition 4.17, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks.
+Proof of Theorem E.10.
+By Observation E.9, it suffices to show that every input or output transaction of a correct agent is eventually included in the output sequence of every correct agent.
+By Proposition E.15, all input transactions in agents’ payloads are eventually sent in (ordinary) blocks.
 Similarly, every output transaction of a correct agent is included in some ordinary block.
 Let $b$ be a non-empty ordinary $p$-block included in some correct agent’s blocklace in a run of Constitutional Consensus.
-If $p$ is correct, then by Proposition 4.11, every correct agent receives and incorporates $b$ in its local blocklace.
+If $p$ is correct, then by Proposition E.11, every correct agent receives and incorporates $b$ in its local blocklace.
 Otherwise, $b$ is finalized in some blocklace, implying that it is approved by at least one $q$-block $b^{\prime}$ by a correct agent $q$.
-Again, by Proposition 4.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
+Again, by Proposition E.11, every correct agent receives and incorporates every block in $b^{\prime}]$, including $b$, in its local blocklace.
 We consider three cases: First, assume the run is finite.
-By Propositions 4.19 and 4.21, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds.
+By Propositions E.16 and E.17, all non-empty ordinary blocks sent or output by correct agents in this run are ordered at all correct agents, and the theorem holds.
 Second, consider a run in which infinitely many waves are quiescent in at least one correct agent’s blocklace at some point in time.
-By Proposition 4.27, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
+By Proposition E.20, if $p$ is correct then $b$ is ordered, and otherwise $b^{\prime}$ is ordered, implying that $b$ is ordered as well because $b\in[b^{\prime}]$.
 Finally, assume the run is infinite and contains only finitely many waves that are quiescent in at least one correct agent’s blocklace at some point in time.
@@
 By fairness of leader selection, there are inifinitely many waves $k^{\prime}>k$ with a correct leader.
-By Proposition 4.25, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
-5 Constitutional Consensus with Constitutional Amendments
-The problem of reconfiguration of a consensus protocol—namely changing the set of its participants—has engaged the distributed computing research community for decades [40, 42, 2, 40], ever since it was first formulated in Lamport’s Paxos paper [21].
-Here we explore the more general problem of constitutional amendments, explained and defined below.
-Supporting constitutional amendments makes the consensus protocol grassroots, as proven in Section 6.
-5.1 Constitutional Consensus: From General to Specific
-In its most general sense, a constitutional consensus protocol is provided with an initial constitution that specifies an initial population, an initial consensus protocol, a rule for amending both, and a rule for amending the constitution itself.
-The notion of constitutional amendment is different from the well-studied notion of reconfiguration in at least two respects:
-  1. 1.
-    Reconfiguration typically refers to changing the set of active agents (servers, processes), whereas a constitution of a consensus protocol may address any aspect of it, including the agents executing it, but also the protocol itself and its parameters.
-  2. 2.
-    Reconfiguration studies are not concerned with who makes reconfiguration decisions: They are typically assumed to descend on obedient agents “out of the blue”.
-Amending the constitution of a protocol, on the other hand, is done by the volitional agents (e.g., people) executing the protocol, in accordance with the prevailing constitutional amendment rule.
-In this work we consider a restricted type of constitution $(P,\sigma,\Delta)$ (Definition 2.1), the components of which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony.
-Even more so, they are key to Constitutional Consensus, given that its primary aim is to execute digital social contracts [9, 38]:
-  1. 1.
-    $P$: The people $P$ are the parties to the contract—members of the democratic digital community, community bank, cooperative, etc.
-Clearly, they should be sovereign to accept new members and remove existing members, namely to amend $P$.
-  2. 2.
-    $\sigma$: As the level of knowledge and trust among the people $P$ evolves, possibly as a result of amending $P$, so should the $\sigma$-supermajority among $P$ be amended to reflect that: If the people in $P$ conclude that they can trust each other more than initially anticipated, $\sigma$ could be decreased.
-Adding new people to $P$, initially known to and trusted by only a fraction of $P$, may entail increasing $\sigma$, even if temporarily, in order to ensure that sufficiently many trusted agents are part of every decision.
-  3. 3.
-    $\Delta$: The delay parameter $\Delta$ is also highly-sensitive to the composition of the population $P$: If all have high-performance smartphones that are well-connected, $\Delta$ can be decreased; if new members with lower-capacity devices are admitted, $\Delta$ may need to be increased.
-In addition, technology advances may allow decreasing $\Delta$.
-Thus, the prevailing population $P$ should be sovereign to amend the value $\Delta$ as they see fit.
-5.2 Democratic Constitutional Amendments
-The Constitutional Consensus protocol attempts to implement valid constitutional amendment decisions as soon as they are made, but is independent of how these decisions are made.
-Computational Social Choice research [6] has investigated democratic decision-making processes extensively.
-Here we describe possible democratic processes for amending each of the components of a constitution $(P,\sigma,\Delta)$.
-(1) Amending $P$.
-For amending $P$ to $P^{\prime}$, we require that a $\sigma$-supermajority among $P$ and that all new members of $P^{\prime}$, namely those in $P^{\prime}\setminus P$, approve the decision.
-The purpose of the supermajority is to ensure that a majority of the correct agents in $P$ are willing to hand-over control to $P^{\prime}$.
-The purpose of requiring all new participants in $P^{\prime}$ to sign is, first, to respect their autonomy—an agent should agree to participate—and, second, to avoid the ‘dead sailors problem’ [21], in which all of $P^{\prime}$ are new but in fact dead, resulting in a constitutional stalemate.
-The respect for autonomy of new participants has a practical corollary: An honest person (who would participate as a correct agent) would not join a protocol instance if they are concerned that the prevailing population may have too many Byzantines, namely that the prevailing constitution is not safe.
-The flip side of this concern—that the prevailing population not add malicious agents—has been studied extensively, as discussed next.
-A constitutional consensus protocol shares a fundamental challenge with permissionless consensus protocols: Both allow the participating agent set to change dynamically, yet must prevent Sybil attacks where a single entity creates multiple digital agents to gain disproportionate influence.
-Permissionless protocols address this by requiring agents to prove control of scarce resources—computational power in Proof-of-Work [27] or financial stake in Proof-of-Stake [19, 8].
-We expect realization of constitutional consensus to take a people’s-based approach: To rely on trust relations among real people, and admit new agents only when existing members know, trust, and verify them as genuine individuals.
-This requires (i) sybil-resistant admission protocols [28, 33] during community formation to ensure that digital agents genuinely represent distinct people rather than fake or duplicate identities and (ii) sybil-resilient democratic governance [32, 24], to ensure that sybils cannot overpower the genuine agents in democratic decisions.
-These mechanisms are complementary to this work.
-(2) Amending $\sigma$.
-Amending the supermajority by which a constitution can be amended has been studied by Abramowitz et al.
-[1], who showed that the $h$-rule (similar to the $h$-index of author citation metrics) is the only rule that satisfies certain intuitive and self-evident axioms.
-According to the $h$-rule, each agent in the prevailing population $P$ votes by stating their most-preferred new value of $\sigma$, in a process left unspecified.
-The value of $\sigma$ is increased to the maximal $\sigma^{\prime}>\sigma$ for which there is a $\sigma^{\prime}$-supermajority among the prevailing population $P$ who voted for a value greater or equal to $\sigma^{\prime}$.
-Else $\sigma$ is decreased to the minimal $\sigma^{\prime}$ for which there is a $\sigma$-supermajority among the prevailing population $P$ who voted for a value less than or equal to $\sigma^{\prime}$.
-Else $\sigma$ remains unchanged.
-Thus, to change $\sigma$ from $\frac{2}{3}$ to $\frac{3}{4}$, the approval of a $\frac{3}{4}$-supermajority among $P$ is needed, and to change to unanimity ($\sigma=\frac{2|P|-1}{2|P|}$), approval by all members of $P$ is needed.
-And if the prevailing constitution states unanimity, then changing back to a $\frac{2}{3}$-supermajority requires unanimous consent among $P$.
-The $h$-rule ensures that changing the supermajority threshold requires meeting that very threshold: To raise $\sigma$ to $\sigma^{\prime}$, you need a $\sigma^{\prime}$-supermajority supporting it (preventing minorities from imposing stricter requirements on the majority), while to lower $\sigma$ from the current value, you need the current $\sigma$-supermajority to agree (preventing minorities from weakening the constitution’s protection).
-(3) Amending $\Delta$.
-The value of the parameter $\Delta$ should be increased or decreased only if the majority of the correct agents wish so.
-To this end, we offer the Suppress Outer-f parameter update rule [^1] In reference [32] the rule refers to ‘outer-$\sigma$’, however the definition of $\sigma$ there is $\frac{f}{n}$, so we changed the name here to avoid confusion.
-by Shahaf et al.
-[32], as follows: Each member of the prevailing population $P$ votes by stating their most-preferred new value of $\Delta$, through a process left unspecified.
-The vote on $\Delta$ should not be based on the subjective experience of each member of $P$, but rather on the perception of each member of $P$ of the “common good” [29].
-If the median of the votes is larger than $\Delta$, then the maximal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}>\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
-Symmetrically, if the median of the votes is smaller than $\Delta$, then the minimal $f$ votes are discarded, the median $\Delta^{\prime}$ of the remaining votes is calculated, and if $\Delta^{\prime}<\Delta$ then the value of $\Delta$ is updated to $\Delta^{\prime}$.
-Else $\Delta$ remains unchanged.
-We note several desirable properties of the Suppress Outer-f rule: If all Byzantines wish to increase $\Delta$ to slow-down the protocol, and the correct agents do not, then $\Delta$ will not be increased.
-Similarly, if all Byzantines wish to decrease $\Delta$ to undermine the liveness of the protocol, and the correct agents do not, then $\Delta$ will not be decreased.
-If all correct agents unanimously wish to update $\Delta$ to $\Delta^{\prime}$ then $\Delta$ will change to $\Delta^{\prime}$ no matter what the Byzantine agents vote, provided $f<\frac{1}{3}n$.
-In summary:
-A constitution $(P,\sigma,\Delta)$ could be amended democratically to $(P^{\prime},\sigma^{\prime},\Delta^{\prime})$ as follows: 1.
-$P$ to $P^{\prime}$ via the approval of both a $\sigma$-supermajority among $P$ and all members of $P^{\prime}\setminus P$.
-2.
-$\sigma$ to $\sigma^{\prime}$ via the h-rule for constitutional amendment [1].
-3.
-$\Delta$ to $\Delta^{\prime}$ via the suppress outer-f parameter update rule [32].
-5.3 Constitutional Amendments
-While we suggest above democratic decision-making processes for amending the three components of the constitution, their specifics are immaterial to the Constitutional Consensus protocol.
-The only thing that matters is the interface between the decision making process and the protocol, which takes the form of constitutional amendment decisions.
-The following definition assumes a Constitutional Consensus protocol instance identifier Id; we discuss its selection below.
-Definition 5.1 (Constitutional Amendment Decision).
-A constitutional amendment decision is a pair $d=(h,x)$, where $x=(\textsc{amend},\textit{Id},i,C,C^{\prime})$, Id is an instance identifier, $i\geq 1$ is the index of the decision within this instance, $C,C^{\prime}$ are constitutions, referred to as the old and new constitutions of $d$, respectively, and $h$ is a multisignature of the hash of $x$ by some $Q\subset\Pi$.
-If $i=1$ then $C=\bot$ and $P_{1}$ are referred to as the founders of the protocol instance Id—they are the signatories of its initial constitution and the participants of the first epoch.
-Note that the structure of $d$ is that of a constitutional genesis block.
-During a run of a protocol instance Id, agents produce a possibly-infinite ordered sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$ using democratic decision-making processes discussed above or others, where $d_{i}=(h_{i},x_{i})$ and $x_{i}=(\textsc{amend},\textit{Id},i,C_{i-1},C_{i})$.
-The protocol commences with $d_{1}$ as the constitutional genesis block of the first epoch, and with each subsequent $d_{i}$ serving as the constitutional genesis block of the $i^{th}$ epoch.
-Definition 5.2 (Valid Constitutional Amendment Decision and Block).
-Given a sequence of constitutional amendment decisions $d_{1},d_{2},\ldots$, the amendment $d_{i}$ with prevailing constitution $C_{i-1}=(P_{i-1},\sigma_{i-1},\Delta_{i-1})$, new constitution $C_{i}=(P_{i},\sigma_{i},\Delta_{i})$, and signatories $Q_{i}$ is valid if $i=1$, $P_{0}=\emptyset$ and $Q_{1}=P_{1}$, or $i>1$, $d_{i-1}$ is valid and the signatories $Q_{i}$ are (i) a $\sigma_{i-1}$-supermajority in $P_{i-1}$, (ii) a $\sigma_{i}$-supermajority in $P_{i}$, and (iii) $(P_{i}\setminus P_{i-1})\subseteq Q_{i}$.
-The sequence is valid if all its members are valid.
-Two non-identical constitutional amendment decisions are conflicting if they have the same instance identifiers and the same index.
-Given a sequence of valid constitutional amendment decisions, a block is constitutional if it includes a valid constitutional amendment decision as payload.
-Note that for agents to determine that a constitutional amendment decision is valid, they need access to all its predecessors.
-Also note that any two consecutive supermajorities may overlap, even be identical.
-In particular, if only $\Delta_{i}$ changes, a $\sigma_{i}$-supermajority among $P_{i}$ is needed.
-If only $\sigma_{i}$ changes then a single $\max(\sigma_{i-1},\sigma_{i})$-supermajority among $P_{i}$ is needed; this ensures that the $h$-rule [1] is followed.
-If only $P_{i}$ changes then a $\sigma_{i}$-supermajority among $P_{i-1}$, a $\sigma_{i}$-supermajority among $P_{i}$, and all new members in $P_{i-1}$, are needed.
-We assume that correct agents do not sign invalid or two conflicting constitutional amendments, and hence can make the following observation.
-{observation}
-[Safety of Constitutional Amendments] Given a valid finite sequence of constitutional amendment decisions $D=d_{1},\ldots d_{i}$, $i\geq 1$, if every $C_{i}$ in $D$ is safe, then the constitutional amendment process cannot extend $D$ with two conflicting valid constitutional amendments.
-Proof 5.3.
-By a counting argument, for this to happen a correct agent would have to sign two conflicting constitutional amendments, a contradiction.
-Note that while the safety depends only on $\sigma_{i}$ and $P_{i}$, validity depends also on $\sigma_{i-1}$ and $P_{i-1}$, in order to adhere to the $h$-rule.
-Grassroots selection of protocol instance identifier.
-Each constitutional amendment decision is associated with an instance Id of the Constitutional Consensus protocol.
-In a grassroots setup, the identifier of each instance should be unique in order to allow for multiple instances to operate concurrently without interference, even if they share some or all of the agents, and should be chosen without central authority, central resource, or coordination among disjoint instances.
-This can be achieved, for example, by having the identifier Id of a protocol instance consist of (the hash of) the founders (initial population) of the instance, each paired with the maximal serial number of any Constitutional Consensus protocol instance they have founded.
-This ensures Id to be unique if the founders include at least one correct agent, who would not sign the initial constitutional decision if its Id is not formed correctly.
-Epochs and agent responsiveness.
-The standard assumption in consensus protocols is that all correct agents are responsive throughout the run.
-However, this assumption is unreasonable in a protocol like Constitutional Consensus in which the set of participating agents changes dynamically, especially since the set $\Pi$ of potential participants is infinite.
-Hence, we only assume that a correct agent $p$ is responsive only during an epoch in which it participates.
-The one exception is the case of agents that are behind, still in “older” epochs.
-To complete the epoch they may need to receive such blocks from former correct participants in that epoch.
-Hence, agent may still need to respond to nack-blocks pertaining to blocks they have issued in past epoch.
-Still, an agent may determine that en epoch is ‘closed’ (and no further nack-blocks need to be served from it) if it has evidence that all participants in the epoch have ended it.
-In particular, an epoch can be ‘closed’ when all its participants have coronated the constitutional amendment decision that ended (see End Epoch below).
-Note that the interaction in which a non-responsive agent $p$ needs to be ‘woken up’ in order to sign a constitutional amendment in which $p$ is a new participant happens outside the Constitutional Consensus protocol.
-In the practical settings we envision this would happen through social interactions (in-person or digital) between existing participants and prospective new participants.
-5.4 Epoch Transitions via Constitutional Amendments
-Protocol concepts.
-Every epoch commences with a constitutional genesis block that specifies its prevailing constitution.
-If, during the epoch, members of the prevailing population become aware of a valid decision $d$ to amend the prevailing constitution to a new constitution, they cease normal operation and only endeavour to finalize this decision.
-Once an agent has a wave $W$ that finalizes a constitutional block with $d$, it coronates $d$ by producing a coronate block with payload $(\textsc{coronate},d)$, sending it to both the old and new populations of $d$, and ceasing participation in the epoch.
-Members of the old population that receive it and have yet to coronate $d$ can nack any missing blocks to obtain $W$ and coronate $d$ themselves.
-A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block.
-An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own (implying that $d$ is final in its blocklace).
-The protocol.
-The protocol produces blocks also with constitutional amendment decisions as payload; such a block is referred to as a constitutional block.
-When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution.
-Subsequent ordered blocks with the same amendment are suppressed from the output.
-The protocol is extended in Algorithm 2 as follows:
-  1. 1.
-    Amendment Finalization occurs whenever the agent $p$ knows of a valid constitutional amendment $d$ that has not been finalized, upon which it continues following the protocol but only producing constitutional blocks with $d$ as a payload, until $d$ is finalized.
-Here, knows has a dual meaning: Either $p$ knows that the external democratic decision making process resulted in a new valid constitutional amendment $d$, or $p$ has a constitutional block with $d$.
-In either case, no constitutional block with $d$ has been finalized yet in $B$.
-  2. 2.
-    End Epoch applies when a constitutional block with $d$ is finalised, upon which members of the old population create a coronate block with $d$, send it to all members of the old and new populations, and cease participation in the old epoch.
-Coronation also plays a role in the democratic decision process: A correct agent would not sign a new constitutional amendment decision before coronating the last constitutional amendment it was a signatory to.
-  3. 3.
-    Start Epoch applies when any agent $p$ in the new population receives from the old population a supermajority of coronate block with decision $d$ (and provided $p$ issued a coronate block if it is also a member of the old population), upon which $p$ creates a constitutional genesis block $b$ with $d$ and updates its epoch identifier to be $b$.
-  4. 4.
-    Respond to Belated nack applies when any agent $p$ receives a nack $q$-block $b$ for a $p$-block of a past epoch $e$, upon which $p$ responds to the nack to help $q$ complete this past epoch.
-Note that if $p$ had previously received a coronate $q$-block ending $e$, then $p$ knows that $q$ is not missing any epoch $e$ blocks; in this case “judicially send…” would be a no-op.
-Moreover, an agent may ‘write off’ an epoch $e$ as complete after receiving from all participants in $e$ blocks coronating the end of $e$.
-Amendment Finalization by $p$: Upon issuing an ordinary block, if $p$ knows of a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload.
-End Epoch by $p$: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to members of $P\cup P^{\prime}$, and cease participation in the epoch.
-Start Epoch by $q$: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
+By Proposition E.19, a leader block is finalized at all correct agents in each of these waves, ordering all non-empty ordinary blocks from earlier rounds.
+∎
+Appendix F The Amendment Setting
+We describe the general democratic operation.
+Unified proposing and voting.
+Following Bulteau et al.
+(Bulteau et al., 2021), there is no proposal phase.
+Whenever an agent decides to vote (afresh or as update), it issues a block with its timestamped and signed vote.
+A vote is persistent: if agent $i$ does not update its vote, the most recent one in the blocklace remains in effect.
+The prevailing constitution thus always reflects the current preferences of all agents without requiring repeated re-voting.
+As we show next, the decision spaces of $P$, $\sigma$, and $\Delta$ are all naturally embedded in corresponding metric spaces.
+Domain structure of the parameters. The three constitutional parameters have distinct domain structures, which determine the appropriate aggregation rule for each:
+  • $P$ (population): For each potential participant $p$, each agent votes yes or no—effectively, a separate referendum on each member. This is the simplest form of status-quo voting.
+  • $\sigma$ (supermajority threshold): Agents state a preferred value in $[0.5,1)$, a naturally single-peaked domain.
+Additionally, $\sigma$ is self-referential: it is the threshold used to amend the constitution, including $\sigma$ itself.
+The amendment rule must therefore be consistent with its own output.
+  • $\Delta$ (timeout): Agents state a preferred value in $\mathbb{R}^{+}$, again a single-peaked one-dimensional domain.
+Appendix G Constitutional Amendment Pseudocode
+Amendment Finalisation by $p$:: Upon issuing an ordinary block, if $p$’s amendment protocol has produced a valid constitutional amendment decision $d$ that is not final in $B$ then issue the block with $d$ instead of with payload.
+End Epoch by $p$:: Upon finalising in $B$ a constitutional block $b$ with constitutional amendment $d$, prevailing population $P$, and new population $P^{\prime}$, call $\tau(b)$, create a $(\textsc{coronate},d)$ block, send it to $P\cup P^{\prime}$, and cease participation in the epoch.
+Start Epoch by $q$:: Let $d$ be a valid decision to amend the prevailing constitution, with new population $P^{\prime}$.
 Upon receipt by $q\in P^{\prime}$ of a $\sigma$-supermajority among $P$ of $(\textsc{coronate},d)$ blocks, and provided $q$ previously issued such a block if $q\in P$, commence participation in the new epoch by $e\leftarrow d$ as the new epoch constitutional genesis block, $B\leftarrow B\cup\{d\}$, and retaining $D$ and payload.
-Respond to Belated nack by $q$: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
+Respond to Belated nack by $q$:: Upon receipt of a nack $q^{\prime}$-block for a $q$-block $b$ of a past epoch $e$ for which $q^{\prime}\in\Pi$ is a participant, judiciously send $[b]$ to $q^{\prime}$.
 Algorithm 2 Incorporating Constitutional Amendments in Constitutional Consensus.
 Code for any participant $p$ in the prevailing constitution and for any $q\in\Pi$.
-In the good case, a new constitutional amendment $d$ is issued by the leader of a wave and finalized within that same wave.
-In the worst case, it might take arbitrarily many waves to finalize $d$ prior to GST or due to faulty leaders.
-After GST, a constitutional amendment block issued by a correct leader among $P$ is bound to be finalized (and such a leader is bound to be selected eventually due to the fairness of $\textit{leader}_{P}$).
-Once finalized, correct members of $P$ will issue a coronate block and correct members of $P^{\prime}$ will receive it, allowing the current epoch to end and the next one to begin.
-This argument informally establishes the liveness of the transition among constitutions due to constitutional amendments.
-5.5 Safety and Liveness of Constitutional Amendments
-Theorem 5.4 (Constitutional Amendment Correctness).
+Epoch transition walkthrough.
+Algorithm 2 shows the additional logic supporting amendments in Algorithm 1.
+Every epoch commences with a constitutional genesis block that specifies its prevailing constitution.
+Amendement finalisation: If, during the epoch, members of the prevailing population produce a valid decision $d$ to amend the prevailing constitution, they endeavour to finalise this decision by producing only constitutional blocks with $d$ as payload.
+End Epoch: Once an agent finalises a constitutional block with $d$, it produces a coronate block with payload $(\textsc{coronate},d)$, sends it to both the old and new populations of $d$, and ceases participation in the epoch.
+Members of the old population that receive it and have yet to finalise $d$ can nack any missing dependent blocks.
+After ceasing to participate in the old epoch, an agent no longer adds blocks from that epoch to its blocklace.
+However, it must continue to respond to Belated nack requests from agents who are missing blocks in order to finalise the old epoch (due to Byzantine equivocation).
+An old epoch may be garbage-collected upon learning that all agents have coronated an ensuing epoch.
+Start Epoch: A member of the new population that receives a $\sigma$-supermajority of coronate blocks may commence the new epoch with $d$ as the constitutional genesis block.
+An agent that is a member of both the new and old populations commences the new epoch after both receiving a $\sigma$-supermajority of coronate blocks and sending one of its own.
+The algorithm’s output includes constitutional blocks with constitutional amendment decisions as payload.
+When $\tau$ orders a constitutional block with amendment $d$ for the first time, the protocol outputs the new constitution, suppressing subsequent copies.
+Appendix H Safety and Liveness of Constitutional Amendments
+Theorem H.1 (Constitutional Amendment Correctness).
 The Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) is correct.
 We need to prove that in every run of the protocol, there exists a sequence $T$ s.t.
-the two safety properties and three liveness properties of Definition 2.4 hold w.r.t.
+the two safety properties and two liveness properties of Definition 2.3 hold w.r.t.
 $T$.
@@
 Let $T$ be the concatenation of all these sequences, $T:=T^{1}\cdot T^{2}\cdots$. We next prove that the protocol satisfies safety and liveness properties w.r.t. $T$. The theorem will follow.
-Lemma 5.5 (Constitutional Amendment Safety).
+Lemma H.2 (Constitutional Amendment Safety).
 Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies safety w.r.t. $T$ defined above.
-Proof 5.6.
+Proof.
 Validity immediately follows from the fact that agents participate only in epochs of which they are members.
 Consistency follows from the fact that agents participate in epochs sequentially, namely, they stop running the protocol in epoch $i$ before starting epoch $j$.
-Lemma 5.7 (Constitutional Amendment Liveness).
+∎
+Lemma H.3 (Constitutional Amendment Liveness).
 Every run of the Constitutional Consensus Algorithm (Algorithm 1) with Constitutional Amendments (Algorithm 2) satisfies liveness w.r.t. $T$ defined above.
-Proof 5.8.
+Proof.
 Consider an epoch $e$ with prevailing constitution $(P,\sigma,\Delta)$, and the two liveness requirements.
-If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem 4.10.
+If $e$ is the last epoch then Epoch Liveness follows from the liveness of the single-epoch Algorithm 1, Theorem E.10.
 Else assume $e$ is not the last epoch, and therefore it must have ended with a coronation by a $\sigma$-supermajority among $P$, all observing a wave finalizing a constitutional block $b$ with payload $d$.
@@
 Else there is at least one correct agent $q\in P$ that coronated $d$ and sent to $p$ a coronate block $b^{\prime}$ observing a wave $W$ that finalizes $b$.
-Since $q$ and $p$ are correct, eventually a $\textsc{nack}(b^{\prime})$-block by $q$ will be received by $p$, who will respond (even after the epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $q$ any blocks in $W$ and their predecessors that $q$ is missing, upon receipt of which $p$ would issue a $(\textsc{coronate},d)$ block and output $\tau(b)$.
+Since the network is reliable, $p$ receives $b^{\prime}$ but may be missing some blocks in $b^{\prime}]$.
+Then $p$ sends a $\textsc{nack}(b^{\prime})$-block to $q$, who responds (even after epoch $e$ has ended, thanks to the Respond to Belated nack rule) by sending to $p$ the missing blocks in $W$ and their predecessors, upon receipt of which $p$ finalizes $b$, issues a $(\textsc{coronate},d)$ block, and outputs $\tau(b)$.
 Considering Constitution Amendment Liveness, once a constitutional amendment $d$ is known, all correct participants follow the high-throughput protocol and produce only blocks with payload $d$.
-Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem 4.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks.
-6 Constitutional Consensus is Grassroots
-We recall the formal definition of a grassroots protocol from reference [38], instantiate it informally (bypassing the formalities of the definition in [38]) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
-Definition 6.1 (Oblivious, Interactive, Grassroots).
+Eventually a correct leader will finalize its block, following the same argument as in the proof of Theorem E.10, upon which all correct participants will eventually issue $(\textsc{coronate},d)$ blocks.
+∎
+Appendix I Constitutional Consensus is Grassroots
+We recall the formal definition of a grassroots protocol from reference (Shapiro, 2026), instantiate it informally (bypassing the formalities of the definition in (Shapiro, 2026)) to the Constitutional Consensus protocol, and then argue that the protocol is grassroots.
+Definition I.1 (Oblivious, Interactive, Grassroots).
 A protocol $\mathcal{F}$ is:
-  1. 1.
+  1. (1)
     oblivious if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$, a run of the protocol over $P$ can proceed just the same in the presence of agents in $P^{\prime}$.
-  2. 2.
+  2. (2)
     interactive if for every $\emptyset\subset P\subset P^{\prime}\subseteq\Pi$ and every run of the protocol over $P$, the protocol can proceed from any configuration $c$ in $r$ to interact with members in $P^{\prime}\setminus P$, namely produce a run that is not possible only over $P$.
-  3. 3.
+  3. (3)
     grassroots if it is oblivious and interactive.
@@
 We thus conclude:
-Theorem 6.2.
+Theorem I.2.
 Constitutional Consensus is grassroots.
-Note that the two requirements cannot even be expressed without in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed.
+Note that the two requirements cannot even be expressed in the context of the basic (no constitutional amendments) Constitutional Consensus protocol, as the population $P$ is given and fixed.
 Also note that Constitutional Consensus being interactive critically-depends on the ability of the prevailing population $P$ to decide to add new members.
-7 Conclusions
-The Constitutional Consensus protocol and its extensions are novel in several respects, all devised in order to serve the needs of future grassroots platforms that operate digital social contracts.
-Initially, we expect low-throughput applications to prevail, as small digital communities engaged in constitutional democratic conduct use the platform to grow and evolve [39].
-Even small community banks and cooperatives [37] may still be, at least initially, on the low-throughput side.
-As communities federate to form ever-larger communities, and local community banks federate to form regional and national banks, higher-throughput execution would follow.
-Until then, simulating the various scenarios would be productive future work.
+Appendix J Related Work
+Constitutional Consensus combines three properties: (i) democratic one-person-one-vote governance, (ii) decentralised BFT consensus, and (iii) self-governing constitutional amendment within the protocol.
+We survey prior work across five areas, corresponding to the five aspects of the protocol described in the introduction, and find that every existing system lacks at least one of these properties.
+J.1. Smart Contracts and Digital Social Contracts
+Deployed smart contract bytecode on Ethereum is immutable; amendments are achieved through proxy schemes that separate mutable state from swappable logic, controlled by admin keys, multisigs, or token-weighted governance (De Filippi et al., 2021).
+Tezos (Goodman, 2014) embeds self-amendment at the protocol level, and the amendment mechanism is itself subject to amendment.
+In both cases, governance is plutocratic and amendments apply to code, not to a broader protocol constitution.
+A digital social contract (Cardelli et al., 2020) differs architecturally: it is a voluntary agreement among people, executed on the participants’ own devices, and enforced by the participants themselves.
+Constitutional Consensus is designed for such contracts.
+Governance is democratic (one person, one vote) rather than plutocratic (one token, one vote), and amendments apply to the protocol constitution—including the participant set, supermajority threshold, and timing—not merely to code.
+These distinctions are summarised in Table 1.
+Table 1. Comparing Smart Contracts and Digital Social Contracts.
+Smart Contract (De Filippi et al., 2021) Digital Social Contract (Cardelli et al., 2020)
+Among: Accounts People
+Executed by: Third-party validators The people’s smartphones
+Cost: Gas None
+Consensus: Always required Only when required by the contract
+Governance: Plutocratic (token-weighted voting) Democratic (one person, one vote)
+Amended via: Proxy schemes and redeployment; bytecode is immutable Constitutional amendments
+Amendment authority: Token-weighted governance Participants themselves via supermajority
+Applications: Decentralized Finance (DeFi) (Schueffel, 2021; Jensen et al., 2021),
+Decentralized Autonomous Organizations (DAOs) (Ethereum, ndao; Ding et al., 2023),
+Non-Fungible Tokens (NFTs) (Wang et al., 2021)
+Social graph and social networking (Shapiro, 2023b),
+grassroots currencies (Shapiro, 2024; Lewis-Pye et al., 2023),
+democratic communities and their federation (Shapiro and Talmon, 2025)
+J.2. DAOs and Blockchain Governance
+Decentralised Autonomous Organisations (Hassan and De Filippi, 2021) are the prevailing method for participant-governed digital organisations.
+Several blockchains can modify their own protocol rules through on-chain stakeholder voting: Tezos (Goodman, 2014) through a five-period procedure requiring an 80% supermajority; Polkadot (Wood, 2016) through forkless runtime upgrades via on-chain referenda; Cardano (Corduan et al., 2023) through a tripartite structure of Constitutional Committee, Delegated Representatives, and Stake Pool Operators; Cosmos (Kwon and Buchman, 2016) through its governance module; and the Internet Computer (Camenisch et al., 2022) through its Network Nervous System.
+All are universally plutocratic: [^5] Cardano is utilizing one-person-one-vote but only for its Constitutional Committee, which has essentially only veto power.
+voting power is proportional to staked tokens.
+This is not accidental; stake-weighting provides built-in Sybil resistance, whereas democratic voting requires an external mechanism to establish unique personhood.
+DAO frameworks built on existing blockchains are similarly plutocratic.
+Aragon (Cuende and Izquierdo, 2017) defaults to ERC-20 token-weighted voting.
+Compound Governor (Leshner, 2020) and OpenZeppelin Governor (OpenZeppelin, 2021) are built around ERC20Votes, tying governance power to token holdings.
+Snapshot (Marino and Snapshot Labs, 2020) offers configurable voting strategies including an equal-weight option, but is purely off-chain and admin-controlled.
+Optimism’s Citizens’ House (Optimism Foundation, 2022) is a notable exception: it implements one-person-one-vote governance for retroactive public goods funding.
+However, the Citizens’ House relies on Ethereum’s proof-of-stake consensus, and the Optimism Foundation retains significant authority over proposal execution.
+Schneider (Schneider, 2022) argues that the economic logic embedded in blockchain systems systematically constrains democratic possibilities.
+Siddarth et al.
+(Siddarth et al., 2020) identify the gap directly: “the lack of a robust notion of personhood … has led to the development of plutocracies.” Talmon (Talmon, 2023) examines DAO governance from a computational social choice perspective.
+Buterin (Buterin, 2021) critiques coin voting as creating plutocratic governance vulnerable to vote buying and proposes proof-of-personhood alternatives.
+No DAO or self-amending blockchain achieves democratic governance with its own BFT consensus and self-governing constitutional amendment.
+Consortium blockchain platforms such as Hyperledger Fabric (Androulaki et al., 2018) and R3 Corda (Brown et al., 2016) implement organisational governance rather than democratic governance.
+In Fabric, configuration changes—including membership, block parameters, and governance policies—require signatures from a majority of organisation administrators, collected off-chain and submitted as configuration update transactions.
+Corda centralises governance further: a network operator controls certificate issuance and network parameters.
+While Fabric’s hierarchical policy system bears structural resemblance to constitutional amendment (each policy specifies who can modify it), decision-making authority rests with designated administrators, not with individual participants exercising equal voting rights.
+J.3. Reconfiguration of Consensus Protocols
+Constitutional amendments generalise reconfiguration in two respects:
+  1. (1)
+    Scope.
+Reconfiguration refers to changing the set of active agents (servers, processes), whereas a constitution describes the entire consensus protocol, and in principle any aspect of it may be amended.
+  2. (2)
+    Authority.
+Even in self-governing systems like Raft (Ongaro and Ousterhout, 2014), reconfiguration is a mechanism available to whoever holds the leader role.
+Constitutional amendments, by contrast, are decided by the participants themselves through a process specified by the constitution.
+In Constitutional Consensus, constitutional amendment preserves the protocol instance: the transaction history (the blocklace) carries over from one epoch to the next under the new constitution.
+Starting a new instance, by contrast, would discard all prior state.
+In this work, the amendable component of the constitution consists of $(P,\sigma,\Delta)$ (Definition 2.1), parameters which affect the safety, liveness, and performance of almost any consensus protocol for eventual synchrony.
+[…]
+The reconfiguration literature has focused on how membership changes execute safely.
+Vertical Paxos (Lamport et al., 2009) introduced an auxiliary configuration master that determines the acceptor set for each configuration.
+DynaStore (Aguilera et al., 2011) allows any client to propose reconfigurations but treats them as externally triggered.
+Spiegelman et al.
+(Spiegelman et al., 2017) explicitly model an administrator invoking reconfiguration operations.
+SMART (Lorch et al., 2006) handles crash-fault reconfiguration via an external coordinator.
+Raft (Ongaro and Ousterhout, 2014) treats configuration changes as special log entries committed through joint consensus of both old and new configurations.
+In the BFT setting, BFT-SMaRt (Bessani et al., 2014) provides the first practical BFT library with reconfiguration support, but relies on an external trusted third party for membership decisions.
+Duan and Zhang (Duan and Zhang, 2022) showed that prior BFT reconfiguration approaches can compromise liveness, and introduced Dyno, the first formally analysed dynamic BFT protocol.
+Proof-of-Stake protocols achieve a form of self-reconfiguration, where validator sets change based on staking transactions (Komatovic et al., 2025), but governance is plutocratic.
+ComChain (Vizier and Gramoli, 2020) and the subsequent SMaRtChain (Bessani et al., 2020) achieve BFT reconfiguration without a trusted third party: current participants run Byzantine consensus to decide on membership changes and employ a forgetting protocol that decouples permanent keys from per-view consensus keys to prevent removed nodes from creating forks.
+However, neither system provides a governance mechanism: the criteria for accepting new members are application-specific (e.g., proof-of-work, certification by an authority).
+Moreover, both are limited to membership changes and do not support amending protocol parameters or governance rules.
+Across all of these, decision authority is either left unspecified, delegated to an external entity, determined by application-specific criteria, or determined plutocratically.
+The mechanism of reconfiguration is well-studied; the question of who decides—through democratic governance—is not.
+J.4. BFT Consensus Protocols
+PBFT (Castro and Liskov, 1999) introduced the first practical algorithm for BFT state machine replication, tolerating up to $\lfloor(n-1)/3\rfloor$ Byzantine faults.
+HotStuff (Yin et al., 2019) achieved linear communication complexity with responsiveness and became the basis for the Diem blockchain.
+Tendermint (Buchman, 2016) brought BFT consensus to blockchains and underlies the Cosmos network.
+A second line of work replaces the leader-based approach with DAG-based protocols.
+DAG-Rider (Keidar et al., 2021) is the first asynchronous BFT Atomic Broadcast protocol achieving optimal resilience, amortised communication complexity, and time complexity simultaneously.
+Narwhal and Tusk (Danezis et al., 2021) separate reliable transaction dissemination from ordering, enabling high throughput.
+Bullshark (Giridharan et al., 2022) adds a practical low-latency synchronous fast path.
+Jolteon and Ditto (Gelashvili et al., 2022) combine leader-based consensus with an asynchronous DAG fallback.
+Mysticeti (Babel et al., 2023) achieves the theoretical lower bound of three message delays for commit latency.
+Raikwar et al.
+(Raikwar et al., 2024) provide a systematisation of knowledge on DAG-based consensus.
+Constitutional Consensus builds on two protocols from this line of work: Cordial Miners (Keidar et al., 2023), a DAG-based BFT Atomic Broadcast protocol, and Morpheus (Lewis-Pye and Shapiro, 2025), which adapts between quiescent and high-throughput modes.
+Both are used as building blocks; Constitutional Consensus adds constitutional amendment and democratic governance above them.
+None of the BFT protocols surveyed here addresses governance or self-amendment: they provide the consensus mechanism but leave the question of who participates and how rules change to external processes.
+J.5. Grassroots Protocols and Democratic Identity
+Democratic digital governance platforms achieve one-person-one-vote participation but rely on centralised infrastructure.
+Decidim (Barandiaran et al., 2024), used by over 500 institutions, runs as a client-server web application.
+Loomio (Loomio Cooperative Limited, 2024) provides equal-weight deliberation and voting, but all data resides on centralised servers.
+CONSUL Democracy (Royo et al., 2020; Consul Democracy Foundation, 2025), serving approximately 250 governments, authenticates users via national ID but operates as a centralised web application.
+LiquidFeedback (Behrens et al., 2018) published a blockchain prototype with a swarm-behaviour-inspired consensus algorithm where node influence derives from the number of accredited persons trusting each node, and the roadmap envisions voting on application code.
+However, LiquidFeedback implements liquid democracy with transitive delegation rather than one-person-one-vote; its consensus is not characterised as BFT; and the self-amendment capability remained a roadmap item.
+Proof-of-personhood (PoP) systems solve the Sybil resistance prerequisite for democratic governance.
+Ford (Ford, 2020) argues that digital personhood—inalienable participation rights built on proof-of-personhood mechanisms—is a necessary foundation for digital democracy.
+PoPCoin (Borge et al., 2017) uses pseudonym parties to enable equal-weight participation in ByzCoin consensus, achieving democratic identity with BFT consensus but not addressing self-governing amendment.
+Humanode (Kavazi et al., 2023) is the closest existing system to combining all three properties: it runs BFT consensus (BABE/GRANDPA on Substrate) with equal per-node weight, using biometric identity verification.
+However, its self-governance mechanism (Vortex DAO) operates as an off-chain simulation as of early 2026, with full authority transfer planned approximately four years after mainnet launch; Vortex is designed as a DAO overlay on the consensus protocol, not as constitutional amendment embedded within it.
+Idena (Idena Team, 2018) achieves equal-weight consensus via flip-test identity verification, but its consensus is not formally characterised as BFT, and protocol upgrades are controlled by the development team.
+Encointer (Brenzikofer, 2019) provides proof of personhood through physical key-signing ceremonies but operates as a Polkadot parachain, inheriting its consensus.
+Because Constitutional Consensus requires no external authority, global ledger, or staked capital, multiple independent instances can form, operate, and grow concurrently—the grassroots property (Shapiro, 2023a, 2026).
+This contrasts with Proof-of-Stake protocols like Ethereum 2.0 (Buterin et al., 2020), whose security depends on the staked token having real-world value, creating a winner-take-all dynamic that precludes independent instances.
+Democratic platforms achieve one-person-one-vote but are centralised; PoP systems solve identity but remain building blocks.
+To the best of our knowledge, no existing system combines democratic one-person-one-vote governance, decentralised BFT consensus, and self-governing constitutional amendment into a single protocol.